### v2

# Existence of 5 minimal tori in 3-spheres of positive Ricci curvature

Abstract.
In 1989, B. White conjectured that every Riemannian 3-sphere contains at least 5 embedded minimal tori. We confirm this conjecture for 3-spheres of positive Ricci curvature. While our proof uses min-max theory, the underlying heuristics are largely inspired by mean curvature flow.
1. Introduction
While enumerative geometry is traditionally a branch of algebraic geometry, there are, in fact, enumerative problems in differential geometry as well.
In 1905, Poincaré conjectured the existence of 333 simple closed geodesic in every smooth, Riemannian 2-sphere [Poi05]. The number 333 is optimal, as is shown in the example of an ellipsoid by Morse [Mor96]. In 1917, G. D. Birkhoff [Bir17] initially confirmed the existence of an (immersed) closed geodesic using a min-max argument, which can be further refined to show that the geodesic is simple. A universally accepted solution to Poincaré’s conjecture was later given by M. Grayson in 1989 through curve shortening flow [Gra89], building upon the earlier work of Lyusternik–Schnirelmann [LS29].
In one dimension higher, S.-T. Yau [Yau82] conjectured that there exist at least 444 embedded minimal 222-spheres in every Riemannian 333-sphere. Simon-Smith [Smi83] showed that there is always at least one embedded minimal 222-sphere using min-max theory. More recently, Wang-Zhou [WZ24] confirmed Yau’s conjecture for the case of metrics with positive Ricci curvature and bumpy metrics (see also the work of Haslhofer-Ketover [HK19]). Furthermore, B. White [Whi89, Whi91] conjectured that every Riemannian 333-sphere contains at least 555 embedded minimal tori. In fact, he proved that every 333-sphere of positive Ricci curvature contains at least one embedded minimal torus using a degree theory he developed alongside Hamilton’s Ricci flow.
In this paper, we confirm B. White’s conjecture for 3-spheres with positive Ricci curvature using min-max theory:
Theorem A.
Every smooth Riemannian 333-sphere of positive Ricci curvature contains at least 555 embedded minimal tori.
Let us briefly explain where the number 555 comes from.
The well-known Willmore conjecture [Wil65, Wil71], which was proved by Marques-Neves [MN14], implies that in the unit 3-sphere $\mathbb{S}^{3}\subset\mathbb{R}^{4}$, the Clifford torus
$T^{2}=\left\{x\in\mathbb{R}^{4}\mathrel{\mathop{\ordinarycolon}}x^{2}_{1}+x^{2}_{2}=x^{2}_{3}+x^{2}_{4}=\frac{1}{2}\right\}\,,$
up to rigid motions in $\mathbb{S}^{3}$, is the unique embedded minimal surface with the second smallest area, after the equator. Moreover, the resolution of Lawson’s conjecture on minimal tori [Law70] by S. Brendle [Bre13] indicates that the Clifford torus is the only embedded minimal torus in the $\mathbb{S}^{3}$. Consequently, the space of embedded minimal tori in $\mathbb{S}^{3}$ is exactly the space $\mathcal{C}$ of Clifford tori, which is homeomorphic to $\mathbb{RP}^{2}\times\mathbb{RP}^{2}$.
Let us also denote $\mathbb{S}^{3}$ by $(S^{3},\bar{g})$, where $\bar{g}$ is the Riemannian metric induced from $\mathbb{R}^{4}$. If we perturb the metric $\bar{g}$, the area functional defined on $\mathcal{C}$ changes. However, the Lyusternik-Schnirelmann number of ${\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$ is 555, meaning every smooth real function on ${\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$ has at least 555 critical points. This is due to the theorem that the Lyusternik-Schnirelmann number of a space is always strictly greater than its cup length in $\mathbb{Z}_{2}$-coefficients (the maximum number of first cohomology classes whose cup product is non-zero), and the topological fact that the cup length of ${\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$ is 444. Based on these facts, B. White proved that every 333-sphere equipped with a metric sufficiently close to the round metric contains at least 555 embedded minimal tori.
It is worth noting that the topological fact that the cup length of ${\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$ is 444 will also play a crucial role in our proof of Theorem A.
For more works on the construction of geodesics or minimal surfaces with controlled topological types, we refer to the works  [Str84, GJ86, Zho16, HK23, Ko23, Ko23a]. Recently, X. Li and Z. Wang also have an independent work on the existence of minimal tori [LW24].
1.1. Heuristics for the proof
Let us describe the heuristics behind the proof of Theorem A. It is worth noting that while the heuristics rely on mean curvature flow, we do not actually use mean curvature flow in practice. Instead, we primarily use Simon-Smith min-max theory, incorporating some concepts from Almgren-Pitts min-max theory.
Let $(S^{3},g_{0})$ be a Riemannian 333-sphere of positive Ricci curvature as given in Theorem A. In their proof of the Willmore conjecture [MN14], F. C. Marques and A. Neves defined a 555-parameter family of flat cycles, known as the canonical family, each associated to a closed embedded surface in the unit 333-sphere. Building upon their construction, we define in $(S^{3},g_{0})$ a 9-parameter family $\Psi$ of surfaces of genus 111 or 0, possibly with singularities; see details in §1.2.1.
Now, we apply mean curvature flow to each member of $\Psi$. Recall that given a smooth, embedded surface $\Sigma\subset M$, the mean curvature flow starting from $\Sigma$ is a family of surfaces $\{\Sigma_{t}\}_{t\in[0,T)}$ in $M$, with $\Sigma_{0}=\Sigma$, satisfying the equation
$\partial_{t}x=\mathbf{H}(x),$
where $x$ is the position vector and $\mathbf{H}$ is the mean curvature vector. Moreover, mean curvature flow can be viewed as the gradient flow of the area functional.
Following the heuristics from gradient flow in the finite-dimensional setting, we assume all such mean curvature flows exist uniquely for all time until the surfaces vanish. Although this assumption is likely false in reality due to the possibility of fattening (i.e. non-unique solutions of mean curvature flow [IW24]), we will ignore this complication in our heuristic argument. Thus, as $t\to+\infty$, we expect the family $\Psi$ to evolve into some family $\Xi$ that is “optimal”, in the sense that $\Xi$ contains minimal tori or minimal spheres, as well as eternal or ancient mean curvature flows connecting these minimal surfaces to one another or to the empty set. These mean curvature flows act as gradient flow lines connecting critical points (see Figure 1). Note that we assume all minimal tori or minimal spheres appearing in $\Xi$ have multiplicity one: This assumption is heuristically justified, as it is conjectured that under the condition of positive Ricci curvature, long time limits of mean curvature flow have multiplicity one [CS24, Conjecture 1.3].
Figure 1. The figure on the left shows the original family $\Psi$. The figure on the right shows $\Xi$, with the minimal surfaces $S_{1},S_{2},S_{3}$ detected, and ancient mean curvature flow lines originating from them. The red part is $T(S_{1})$, the green part is $T(S_{2})$, and the blue part is $T(S_{3})$.
For each minimal surface $S$ detected by $\Xi$, there are eternal or ancient mean curvature flows originating from $S$. Let $T(S)$ denote the collection of all time slices in such mean curvature flows. Note that $T(S)$ is a subset of $\Xi$, and can be viewed as the “unstable manifold” for $S$ (see Figure 1). Moreover, $\Xi=\bigcup_{S}T(S)$ where $S$ ranges over all minimal surfaces (of genus 0 or 111) detected by $\Xi$. If $S$ is a minimal sphere, then each element of $T(S)$ should have genus 0, possibly with singularities, since mean curvature flow does not increase genus [BK24].
To prove Theorem A, suppose for the sake of contradiction that there are only 444 minimal tori $T_{1},T_{2},T_{3},T_{4}$ detected by $\Xi$. The case with fewer than 444 minimal tori can be handled similarly. Then, letting $D_{0}\mathrel{\mathop{\ordinarycolon}}=\bigcup_{S}T(S)$ where $S$ ranges over the minimal spheres detected by $\Xi$, and $D_{i}\mathrel{\mathop{\ordinarycolon}}=T(T_{i})$ for $i=1,\cdots,4$, we have
(1.1) $\Xi=D_{0}\cup D_{1}\cup D_{2}\cup D_{3}\cup D_{4}\,.$
Next, let us describe a crucial topological fact about $\Psi$. In fact, ensuring this topological fact holds is the main reason we constructed $\Psi$ using the Marques-Neves canonical family. Recall that the cycle space $\mathcal{Z}_{2}(S^{3};\mathbb{Z}_{2})$ is weakly homotopy equivalent to ${\mathbb{RP}}^{\infty}$, and thus, the cohomology ring $H^{*}(\mathcal{Z}_{2}(S^{3};\mathbb{Z}_{2});\mathbb{Z}_{2})$ can be written as $\mathbb{Z}_{2}[\bar{\lambda}]$, where $\bar{\lambda}$ denotes the unique non-trivial element in $H^{1}(\mathcal{Z}_{2}(S^{3};\mathbb{Z}_{2});\mathbb{Z}_{2})$. Since each member of $\Psi$ can be viewed as a mod 222 cycle, we can view $\Psi$ as a subset of $\mathcal{Z}_{2}(S^{3};\mathbb{Z}_{2})$. Let $\lambda$ be the element in $H^{1}(\Psi;\mathbb{Z}_{2})$ induced by $\bar{\lambda}$. The crucial topological fact is that, there exist $\alpha,\beta\in H^{1}(\Psi;\mathbb{Z}_{2})$ such that
(1.2) $\lambda^{5}\cup\alpha^{2}\cup\beta^{2}\neq 0\,.$
By construction, $\Xi$ and $\Psi$ are homotopic to each other as sweepouts, so we can view $\lambda,\alpha$ and $\beta$ as elements of $H^{1}(\Xi;\mathbb{Z}_{2})$ as well. Hence, by an elementary Lyusternik–Schnirelmann argument, (1.1) and (1.2) together immediately imply that one of the following must hold:
  1. (1)
    The element $\lambda^{5}|_{D_{0}}$ in $H^{5}(D_{0};\mathbb{Z}_{2})$ induced by $\lambda^{5}$ is non-zero.
  2. (2)
    The element $\alpha|_{D_{1}}$ in $H^{1}(D_{1};\mathbb{Z}_{2})$ induced by $\alpha$ is non-zero.
  3. (3)
    The element $\alpha|_{D_{2}}$ in $H^{1}(D_{2};\mathbb{Z}_{2})$ induced by $\alpha$ is non-zero.
  4. (4)
    The element $\beta|_{D_{3}}$ in $H^{1}(D_{3};\mathbb{Z}_{2})$ induced by $\beta$ is non-zero.
  5. (5)
    The element $\beta|_{D_{4}}$ in $H^{1}(D_{4};\mathbb{Z}_{2})$ induced by $\beta$ is non-zero.
And to derive a contradiction, it suffices to show that statements (1) to (5) are all impossible.
To show that (1) is impossible, first note that by construction, $D_{0}$ consists entirely of genus 0 surfaces, possibly with singularities. Now, let us consider $D_{0}$ as a family in the unit round 333-sphere $\mathbb{S}^{3}$, instead of the original 333-sphere $(S^{3},g_{0})$. Note the following two facts:
  • If we apply the Simon-Smith min-max theorem to $D_{0}$ in $\mathbb{S}^{3}$, then by Wang-Zhou’s multiplicity one theorem [WZ24], we detect the multiplicity-one equatorial 222-sphere, which has area $4\pi$.
  • On the other hand, if (1) holds, meaning $\lambda^{5}|_{D_{0}}\neq 0$, then by definition, $D_{0}$ is a 555-sweepout in the sense of the Almgren-Pitts min-max theory. Since the Simon-Smith width must be greater than or equal to the Almgren-Pitts width, the Simon-Smith width of the family $D_{0}$ (viewed in $\mathbb{S}^{3}$) is at least the 555-width of $\mathbb{S}^{3}$, which is $2\pi^{2}$ by C. Nurser [Nur16].
The above two facts are contradictory, so (1) is impossible.
To show that (2) is impossible, note that by backtracking the mean curvature flow lines, $D_{1}$ can be deformation retracted to the single minimal torus $\{T^{1}\}$. Thus, $H^{1}(D_{1};\mathbb{Z}_{2})$ is zero, making (2) is impossible. Similarly, (3) through (5) are also impossible. This concludes the proof of Theorem A.
1.2. Details behind the heuristics.
The above heuristic argument contains several details and inaccuracy that need to be addressed.
1.2.1. The 999-parameter family $\Psi$
Let us now explain the construction of the family $\Psi$. Let $\mathcal{C}$ (resp. $\widetilde{\mathcal{C}}$) denote the set of unoriented (resp. oriented) Clifford tori in the unit 3-sphere $\mathbb{S}^{3}$. Note that $\mathcal{C}\cong{\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$, and $\widetilde{\mathcal{C}}$ is a double cover of this space. For each $\Sigma\in\widetilde{\mathcal{C}}$, based on Marques-Neves’ canonical family, C. Nurser in [Nur16] constructed a 5-sweepout
$\Phi^{\Sigma}_{5}\mathrel{\mathop{\ordinarycolon}}{\mathbb{RP}}^{5}\to\mathcal{Z}_{2}(\mathbb{S}^{3};\mathbb{Z}_{2})\,.$
Here is his construction: The set of conformal diffeomorphisms of $\mathbb{S}^{3}$ can be parametrized by the open unit 4-ball $\mathbb{B}^{4}$, which we denote by $\{F_{v}\}_{v\in{\mathbb{B}^{4}}}$. For each image $F_{v}(\Sigma)$, which is oriented, we consider the level surfaces of the signed distance function to $F_{v}(\Sigma)$. This gives us a 555-parameter family of surfaces, possibly with singularities and possibly empty, parametrized by ${\mathbb{B}^{4}}\times[-\pi,\pi]$, since the set of possible signed distances from $F_{v}(\Sigma)$ lies in the range $[-\pi,\pi]$.
Marques-Neves [MN14, §5] showed that by ingeniously reparametrizing this family near the boundary of its parameter space, one can continuously extend the family to the boundary, obtaining a $\overline{{\mathbb{B}^{4}}}\times[-\pi,\pi]$-family. Moreover, any two “antipodal” points on the boundary of this parameter space can be identified, as they represent the same surface (but with opposite orientation), thereby yielding an ${\mathbb{RP}}^{5}$-family $\Phi^{\Sigma}_{5}$.
Repeating this for every oriented Clifford torus $\Sigma$, we obtain an ${\mathbb{RP}}^{5}\times\widetilde{\mathcal{C}}$-family. The subfamilies $\Psi^{\Sigma}_{5}$ and $\Psi^{-\Sigma}_{5}$ have the same images for each $\Sigma\in\widetilde{\mathcal{C}}$, where $-\Sigma$ denotes $\Sigma$ with an opposite orientation. By identifying all such pairs of subfamilies $\Psi^{\Sigma}_{5}$ and $\Psi^{-\Sigma}_{5}$, we obtain a family $\Psi$ with parameter space $Y$, which is an ${\mathbb{RP}}^{5}$-bundle over $\mathcal{C}\cong{\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$. This family consists only of genus 0 and genus 111 surfaces, possibly with singularities, and is suitable for running the Simon-Smith min-max process.
1.2.2. Overcoming the issue of fattening
In the heuristic argument of §1.1, we proposed using mean curvature flow to obtain an “optimal family” $\Xi$ from $\Psi$. In practice, the authors were unable to implement this strategy due to the phenomenon of fattening. Indeed, if any member of $\Psi$ fattens along the mean curvature flow, then the flow no longer generates a canonical deformation that is continuous across the entire family $\Psi$ of initial conditions. As a result, it is unclear how one might obtain the desired family $\Xi$ using this approach.
Instead of using mean curvature flow, we employ a scheme of repeatedly applying the Simon-Smith min-max to $\Psi$ in order to obtain $\Xi$. Furthermore, we will define a process called pinch-off process that resembles mean curvature flow, so that each member $\Psi(x)$ will deform under this pinch-off process to some member $\Xi(x)$. Crucially, this process is genus-non-increasing in time. More precisely, a pinch-off process involves three types of behavior: (1) isotopy, (2) neck-pinch surgery, (3) shrinking some components in points: See Figure 2. The “genus-non-increasing” nature of pinch-off process (discussed in §9) should be compared with the analogous results on mean curvature flow by B. White [Whi95]. At the end, although the family $\Xi$ we obtain may not be as well-behaved as to be considered “optimal” in the sense described in §1.1, it is sufficient for our purposes.
Figure 2. This is an example of pinch-off process. It has one neck-pinch surgery, and one connected component shrunk to a point.
One might ask why we need the pinch-off process. Indeed, solely by the usual Simon-Smith min-max theorem, one already could homotope the initial family $\Psi$ to some other family that is, a certain sense, optimal. The problem is the following. When this optimal family detects a minimal sphere, let us consider the “cap” region of this optimal family that is near the minimal sphere. Every member of this cap would only be varifold-close to the minimal sphere, and thus could actually have genus 111 but with a short non-trivial loop. As we will see in §1.2.3, this would be problematic (and deviates from the heuristic picture in §1.1 too). Hence, we need to use the pinch-off process, to eliminate the short non-trivial loops, and ensure that the cap consists of only genus 0 surfaces.
We remark that, this “interpolation” process of modifying the whole high area cap such that every member becomes genus 0 is a non-trivial step. Namely, we need to perform the pinch-off process continuously for every member. We will briefly explain the strategy in §1.2.6.
1.2.3. Obtaining $\Xi$ through repeated applications of the min-max theorem
First, we assume that the 3-sphere $(S^{3},g_{0})$ given has only finitely many minimal tori. Then, we perturb the metric $g_{0}$ such that it has only finitely many minimal spheres, all of which are non-degenerate and have linearly independent areas over $\mathbb{Z}$, with no linear combination over $\mathbb{Z}$ of their areas equal to the area of any minimal torus; see §6.
In the first stage of min-max process, we apply the Simon-Smith min-max construction to $\Psi$, homotoping it obtain a new pulled-tight “optimal” family $\widetilde{\Psi}^{1}$ that detects finitely many min-max minimal surfaces of genus 0 or 111. For simplicity, we will assume the new parameter space $\mathrm{dmn}(\widetilde{\Psi}^{1})$ is also $Y$.
In the new perturbed metric, using D. Ketover’s strengthened genus bound [Ket19] and Wang-Zhou’s multiplicity one theorem [WZ24], which builds on Sarnataro-Stryker’s work [SS23], it is not hard to see that there are two possible cases about the detected minimal surfaces: (1) each min-max minimal surface is a multiplicity one minimal torus, or (2) only one min-max minimal surface is detected, and it is a multiplicity-one sphere. Note that the Frankel property is needed, which is ensured by the assumption of positive Ricci curvature. For simplicity, in case (1), we assume that there is only one minimal torus.
In both cases, we consider a subset $C^{1}\subset\mathrm{dmn}(\widetilde{\Psi}^{1})$ such that the restriction $\widetilde{\Psi}^{1}|_{C^{1}}$ is close to the minimal surface $V^{1}$ detected. The family $\widetilde{\Psi}^{1}|_{C^{1}}$ is like a “cap” of large area, such that its complement $\Psi^{1}\mathrel{\mathop{\ordinarycolon}}=\widetilde{\Psi}^{1}|_{\overline{Y\backslash C^{1}}}$ has the property that the maximal area of $\Psi^{1}$ is strictly less than the width of $\Psi$, $\operatorname{area}(V^{1})$. Let $Y^{1}$ denote the new domain $\mathrm{dmn}(\Psi^{1})=\overline{Y\backslash C^{1}}$.
For case (2) only, we will prove certain interpolation results that allow us, without loss of generality, to assume that $\widetilde{\Psi}^{1}|_{C^{1}}$ consists entirely of genus 0 surfaces, possibly with singularities: This is where the pinch-off process described in §1.2.2 comes into place.
Now, we proceed to the second stage of the min-max process, this time applied to $\Psi^{1}$, regardless of whether case (1) or (2) occurred previously. The width at this stage will be strictly smaller than the previous width, leading to some new minimal surface. We repeat this process iteratively.
Namely, at the $k$-th stage of the min-max process, we start with a family $\Psi^{k-1}$ with parameter space $Y^{k-1}$. Applying the Simon-Smith min-max theorem to $\Psi^{k-1}$ would produce a family $H^{k}$ with parameter space $[0,1]\times Y^{k-1}$ , resembling a homotopy, such that $H^{k}(0,\cdot)=\Psi^{k-1}$, and the new family $\widetilde{\Psi}^{k}\mathrel{\mathop{\ordinarycolon}}=H^{k}(1,\cdot)$ (with parameter space $Y^{k-1}$) is an “optimal” family detecting a multiplicity-one minimal surface $V^{k}$ of genus 0 or 111.
[…]
Since we have perturbed the metric $g_{0}$ so that all minimal spheres are non-degenerate, there must be only finitely many of them, by Choi-Schoen’s compactness result [CS85]. Consequently, since the width of $\Psi^{k}$ decreases strictly with each iteration, the repetitive min-max process must terminate. Namely, there exists a $K>0$ such that at the $K$-th stage of min-max process, the width of $\Psi^{K-1}$ is zero. At this point, there exists a family $H^{K}$ (which acts like a homotopy) with parameter space $[0,1]\times Y^{K-1}$, where $H^{K}(0,\cdot)=\Psi^{K-1}$, and $\widetilde{\Psi}^{K}\mathrel{\mathop{\ordinarycolon}}=H^{K}(1,\cdot)$ has a maximal area arbitrarily close to zero.
[…]
$\widetilde{\Psi}^{1}|_{C^{1}}$
$H^{2}|_{[0,1]\times\partial Y^{1}}$
$\widetilde{\Psi}^{2}|_{C^{2}}$
$H^{3}|_{[0,1]\times\partial Y^{2}}$
($2K-4$)
$H^{K-1}|_{[0,1]\times\partial Y^{K-2}}$
($2K-3$)
$\widetilde{\Psi}^{K-1}|_{C^{K-1}}$
($2K-2$)
$H^{K}|_{[0,1]\times\partial Y^{K-1}}$
($2K-1$)
$\Psi^{K}$
Figure 3. Constructing $\Xi$.
Immediately from the definition of these $2K-1$ families, there exists a natural way to glue their boundaries together, as illustrated in Figure 3. Through this process, we obtain a new 999-parameter Simon-Smith family $\Xi$ that can be deformed to the original family $\Psi$; see Proposition 3.11.
1.2.4. Decomposing $\Xi$ into “unstable manifolds”
In the heuristic argument in §1.1, we decomposed $\Xi$ into a union of “unstable manifolds” $T(S)$ associated with each minimal surface $S$ detected in $\Xi$. There are two issues with this approach.
First, the family $\Xi$ we obtained in §1.2.3, when viewed as a subset of the cycle space $\mathcal{Z}_{2}(S^{3};\mathbb{Z}_{2})$, may have complicated topology and thus be difficult to work with from an algebraic topology perspective. Thus, in practice we actually work with its parameter space $\mathrm{dmn}(\Xi)$, and decompose $\mathrm{dmn}(\Xi)$ instead.
Second, our family $\Xi$ no longer consists of mean curvature flow lines. Instead of defining $T(S)$, we are going to define the trace $T(C)$ of each cap $C=C^{1},C^{2},\cdots,C^{K-1}$, and each trace $T(C)$ will be a subset of $\mathrm{dmn}(\Xi)$.
Let us fix a cap $C=C^{k}$, obtained at the $k$-th stage of the min-max process. Consider the following sets, which can all be viewed as subsets of $\mathrm{dmn}(\Xi)$ (see Figure 4):
  • $B_{k}$:$\;\mathrel{\mathop{\ordinarycolon}}=C$,
  • ${\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}B_{k+1}}\mathrel{\mathop{\ordinarycolon}}=[0,1]\times\partial Y^{k}=\mathrm{dmn}(H^{k+1})$,
  • ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}B_{k+2}}\mathrel{\mathop{\ordinarycolon}}=[0,1]\times\left((\{1\}\times\partial Y^{k})\cap(\{0\}\times\partial Y^{k+1})\right)\subset\mathrm{dmn}(H^{k+2})$. Here, the subsets $\{1\}\times\partial Y^{k},\{0\}\times\partial Y^{k+1}\subset\mathrm{dmn}(\Xi)$ can both also be viewed as subsets of $\mathrm{dmn}(\widetilde{\Psi}^{k+1})$
  • ${B_{k+3}}\mathrel{\mathop{\ordinarycolon}}=[0,1]\times\left((\{1\}\times\partial Y^{k+1})\cap(\{0\}\times\partial Y^{k+2})\right)\subset\mathrm{dmn}(H^{k+3})$.
Finally, we define the trace of $C$ by
$T(C)\mathrel{\mathop{\ordinarycolon}}=B_{k}\cup B_{k+1}\cup\cdots\cup B_{K}\subset\mathrm{dmn}(\Xi)\,.$
A crucial property of $T(C)$ is that, by definition, $T(C)$ admits a strong deformation retraction back onto the cap $C$.
Figure 4. The trace $T(C)$ in $\mathrm{dmn}(\Xi)$.
In fact, if $C=C^{k}$ is a cap associated with a genus 0 minimal surface, then by construction $\Xi|_{T(C)}$ consists entirely of genus 0 surfaces, possibly with singularities, because $\widetilde{\Psi}^{k}|_{C^{k}}$ does and the “homotopies” given by $H^{k+1},H^{k+2},...$ do not increase genus (this should be compared with the situation in §1.1, where mean curvature flow does not increase genus).
Now, let $N$ be the number of minimal tori we detected throughout the first $K-1$ stages of the min-max process. To prove Theorem A, we only need to show that $N\geq 5$.
Let us see why $N=4$ is impossible; the cases for $N=0,...,3$ are similar. Given a cap $C=C^{k}$, we say that it has genus 0 if the minimal surface $V^{k}$ associated has genus 0, and genus 111 if $V^{k}$ has genus 111. For example, when $N=4$, we have four genus 111 caps, and finitely many genus 0 caps. We can express $\mathrm{dmn}(\Xi)$ as the union of the following five subsets:
  • Let $D_{0}$ be the union of $\mathrm{dmn}(\Psi^{K})$ and $\bigcup_{C}T(C)$, where $C$ ranges over all the genus 0 caps.
  • For each genus 111 cap $C$, we consider its trace $T(C)$. There are in total 444 such traces, which we will denote by $D_{1},\cdots,D_{4}$.
It follows directly from the definition of $\Xi$ that
$\mathrm{dmn}(\Xi)=D_{0}\cup D_{1}\cup D_{2}\cup D_{3}\cup D_{4}\,.$
1.2.5. Topological arguments
With this setup, we can proceed with the remaining steps of the heuristic argument in §1.1. However, instead of considering the cohomology rings of $\Psi$ and $\Xi$, we should focus on those of $Y$ and $\mathrm{dmn}(\Xi)$. For example, we should instead define $\lambda$ as an element of $H^{1}(Y;\mathbb{Z}_{2})$, by $\lambda\mathrel{\mathop{\ordinarycolon}}=\Psi^{*}(\bar{\lambda})$, where $\Psi$ is viewed as a map into $\mathcal{Z}_{2}(S^{3};\mathbb{Z}_{2})$. The crucial topological fact is that, there exist $\alpha,\beta\in H^{1}(Y;\mathbb{Z}_{2})$ such that
(1.3) $\lambda^{5}\cup\alpha^{2}\cup\beta^{2}\neq 0\,.$
Moreover, since $Y$ and $\mathrm{dmn}(\Xi)$ are actually homotopy equivalent, $\lambda,\alpha$ and $\beta$ can be viewed as elements of $H^{1}(\mathrm{dmn}(\Xi);\mathbb{Z}_{2})$ as well.
Then, as in §1.1, we apply a Lyusternik-Schnirelmann argument to derive a contradiction. The proof that $\lambda^{5}|_{D_{0}}=0$ proceeds similarly to the earlier argument. However, proving $\alpha|_{D_{i}}=0$ for $i=1,...,4$ requires more essential modifications. For instance, unlike in §1.1, $D_{1}$ may not be contractible (let us focus on $i=1$, as the cases of $i=2,3,4$ are the same). Thus, we would instead prove $\alpha|_{D_{1}}=0$ by showing that the map
$H_{1}(D_{1};\mathbb{Z}_{2})\to H_{1}(\mathrm{dmn}(\Xi);\mathbb{Z}_{2})$
induced by the inclusion $D_{1}\hookrightarrow\mathrm{dmn}(\Xi)$ is trivial.
[…]
Let $Z_{0}\subset Y$ be the set of parameters corresponding to (unoriented) Clifford tori under $\Psi$. From the definition of $\Psi$ and the fact that all members of $\Psi|_{c_{0}}$ have genus one, it is not hard to show that one can deform $c_{0}$ into $Z_{0}$. Thus, it suffices to show that $c_{0}$ is homologically trivial in $Z_{0}$. Note that $Z_{0}\cong{\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$, so $H_{1}(Z_{0};\mathbb{Z}_{2})$ consists four elements: (0,0),(0,1),(1,0),(1,1)(0,0),(0,1),(1,0),(1,1)( 0 , 0 ) , ( 0 , 1 ) , ( 1 , 0 ) , ( 1 , 1 ).
First, using the fact that $\Xi|_{c}$ is close to a single minimal torus, $[c_{0}]$ cannot be (0,1)(0,1)( 0 , 1 ) or (1,0)(1,0)( 1 , 0 ), as both correspond to 111-sweepouts under $\Psi$. As for (1,1)(1,1)( 1 , 1 ), one can check that it gives rise to, under $\Psi$, a loop of oriented Clifford tori, $\{\Sigma(\theta)\}_{\theta\in[0,2\pi]}$, such that the followings hold:
  • We can continuously pick an interior region $\mathrm{in}(\Sigma(\theta))$, bounded by $\Sigma(\theta)$, for each $\theta$.
  • While the interior regions $\mathrm{in}(\Sigma(0))$ and $\mathrm{in}(\Sigma(2\pi))$ coincide, the isomorphism map from $\pi_{1}(\mathrm{in}(\Sigma(0)))$ to $\pi_{1}(\mathrm{in}(\Sigma(2\pi)))$ (both groups are $\mathbb{Z}$) induced by the motion $\{\Sigma(\theta)\}_{\theta\in[0,2\pi]}$ is not the identity but 1-1- 1.
We would show that this is impossible, due to the fact that $\Xi|_{c}$ is close to a single minimal torus, and that $\Psi|_{c_{0}}$ was obtained from $\Xi|_{c}$ through a homotopy which is a family consisting of genus one surfaces possibly with singularities.
Hence, $[c_{0}]$ must be the trivial element (0,0)(0,0)( 0 , 0 ), as desired. This means $\alpha|_{D_{1}}=0$. Similarly, $\alpha|_{D_{2}}$, $\beta|_{D_{3}}$, $\beta|_{D_{4}}$ are all trivial. This leads to a contradiction, as shown in §1.1. This completes the proof of Theorem A.
1.2.6. Interpolation
Lastly, let us briefly explain the interpolation process mentioned at the end of §1.2.2.
Suppose we have a “cap” region $C$ consisting of genus 111 surfaces, each being $\mathbf{F}$-close to a minimal sphere ($\mathbf{F}$ is the distance for currents). It is expected that each such surface contains a short non-trivial loop. If such short non-trivial loops could be selected in a continuous manner, one might perform a continuous family of surgeries on the cap to produce a new cap consisting of genus 0 surfaces. However, to the best of the authors’ knowledge, there is no canonical method for locating such a continuous family of short loops.
Instead, we replace the neck-pinch surgery by a two-step “pinch-off process”. First, for each surface $\Sigma$ in $C$, we identify a 333-ball $B(p_{\Sigma},r_{\Sigma})\subset S^{3}$ that contains a short non-trivial loop of $\Sigma$. We then perform a sequence of neck-pinch surgeries on $\Sigma$ along each loop component of $\Sigma\cap\partial B(p_{\Sigma},r_{\Sigma})$, so that afterwards $\Sigma$ avoids the sphere $\partial B(p_{\Sigma},r_{\Sigma})$. In the second step, we shrink everything inside $B(p_{\Sigma},r_{\Sigma})$ to a point.
The advantage of this alternative method is that it avoids the need to choose the 333-balls continuously. Instead, we fix an open covering $\{O_{i}\}_{i}$ of $C$ and select a suitable 333-balls $B_{i}$ for each $O_{i}$ such that $\partial B_{i}\cap\partial B_{j}=\emptyset$ if $O_{i}\cap O_{j}\neq\emptyset$. If such a selection exists, one can construct a continuous family of two-step pinch-off processes in $C$ using a standard cut-off argument. Fortunately, this is indeed feasible, and a similar construction was already carried out by Pitts [Pit81]; see Section 8 for details.
The actual construction is more involved than the outline given above, due to the following reasons:
  1. (1)
    Proving the existence of a short non-trivial loop on each surface requires some careful argument.
  2. (2)
    It is crucial to control the area during interpolation in order to ensure at each repetition of the min-max procedure, the width drops.
  3. (3)
    The “surfaces” in the cap $C$ may not be smooth, so we need multiple balls to handle each one appropriately.
1.3. Open problems
First of all, the full conjecture of B. White is still open. However, in light of the work of [WZ24], we are more optimistic about the existence of at least five embedded minimal tori for a generic metric, though this number might not be sharp for generic metrics, as discussed below.
Recall that the Morse number of a manifold is the minimum number of critical points a Morse function can have. Since $\mathbb{RP}^{2}\times\mathbb{RP}^{2}$ has Morse number 999, B. White [Whi91] conjectured the existence of at least 999 embedded minimal tori in a 333-sphere equipped with a generic metric.
Inspired by the works of [MN16, Li23a, Li23], and as a further inquiry based on the current work, we conjecture that in every 333-sphere with positive Ricci curvature, there exist at least 5 embedded minimal tori of Morse index at most 999.
It is also natural to raise the question of counting minimal surfaces with higher genus. For instance, let $\mathcal{M}_{2}$ denote the space of embedded, unoriented, genus 222 Lawson surfaces in the unit 333-sphere. This space is a closed 666-dimensional manifold. In the spirit of [Whi91], we conjecture that in every Riemannian 333-sphere, the number of closed, embedded, genus 222 minimal surfaces is at least one plus the cup length of $\mathcal{M}_{2}$. This conjecture is, of course, based on the fact that the Lyusternik–Schnirelmann number of a manifold is bounded below by one plus the cup length. We also conjecture that in every Riemannian 333-sphere equipped with a generic metric, the number of closed, embedded, genus 222 minimal surfaces is at least the Morse number of $\mathcal{M}_{2}$.
Finally, we pose the following questions as potential steps towards resolving the fattening issue described in §1.2.2 from a purely mean curvature flow perspective. Given a smooth, embedded, closed surface $\Sigma$ in a 3-manifold, let $\mathfrak{M}$ denote the union of all time slices of all mean curvature flows with initial data $\Sigma$, in whichever appropriate weak sense. Is the set $\mathfrak{M}$ connected? More ambitiously, is $\mathfrak{M}$ contractible?
If $\mathfrak{M}$ is contractible, it may be possible to run mean curvature flow to the whole family $\Psi$ and surpass the time of fattening; see also the recent work of J. Bernstein, L. Chen, and L. Wang regarding the set of expanders flowing out of a cone [BW22, BCW24].
1.4. Organization
This paper is divided in two main parts. In part I, we first present the preliminary results in §2, which primarily concern about various min-max theories, and then prove the main theorem in §3. In part II, we prove various theorems and propositions used in part I.
Acknowledgment
We are deeply grateful to Zhihan Wang for his generous help, which has been instrumental in the completion of this project. We also wish to thank André Neves for his constant support and guidance. The second author would also like to thank Jonanthan Zung for answering questions on geometric topology, especially those related to Example 2.6.
This material is based upon work supported by the National Science Foundation under Grant No. DMS-1928930, while the first author was in residence at the Simons Laufer Mathematical Sciences Institute (formerly MSRI) in Berkeley, California, during the Fall semester of 2024. The second author was partially supported by the AMS-Simons travel grant.
Part I Main arguments
2. Preliminaries
Throughout this paper, all the ambient Riemannian manifolds $(M,g)$ are assumed to be smooth and closed. We will use $M$ to denote the manifold when the metric $g$ is clear from the context.
2.1. Notations
  • $\mathbf{I}_{n}(M;\mathbb{Z}_{2})$: the set of integral $n$-dimensional currents in $M$ with $\mathbb{Z}_{2}$-coefficients.
  • $\mathcal{Z}_{n}(M;\mathbb{Z}_{2})\subset\mathbf{I}_{n}(M;\mathbb{Z}_{2})$: the subset consisting of elements $T$ such that $T=\partial Q$ for some $Q\in\mathbf{I}_{n+1}(M;\mathbb{Z}_{2})$. Such $T$ are also referred to as (modulo 222) flat $n$-cycles.
  • $\mathcal{Z}_{n}(M;\nu;\mathbb{Z}_{2})$ with $\nu=\mathcal{F},\mathbf{F},\mathbf{M}$: the set $\mathcal{Z}_{n}(M;\mathbb{Z}_{2})$ equipped with the three topologies given corresponding respectively to the flat norm $\mathcal{F}$, the $\mathbf{F}$-metric, and the mass norm $\mathbf{M}$ (see [Pit81] and the survey paper [MN20]). For the flat norm, there are two definitions that, by the isoperimetric inequality, would induce the same topology:
    $\mathcal{F}(T)\mathrel{\mathop{\ordinarycolon}}=\inf\{\mathbf{M}(P)+\mathbf{M}(Q)\mathrel{\mathop{\ordinarycolon}}T=P+\partial Q\}\,,$
    and
    $\mathcal{F}(T)\mathrel{\mathop{\ordinarycolon}}=\inf\{\mathbf{M}(Q)\mathrel{\mathop{\ordinarycolon}}T=\partial Q\}\,.$
    In this paper, we will use the second definition.
  • $\mathcal{V}_{n}(M)$: the closure, in the varifold weak topology, of the space of $n$-dimensional rectifiable varifolds in $M$. There is a $\mathbf{F}$ metric on $\mathcal{V}_{n}(M)$, which also induces the varifold weak topology when restricted to any subset of $\mathcal{V}_{n}(M)$ whose elements have volume bounded above by a constant.
  • $\mathinner{\!\left\lVert V\right\rVert}$: the Radon measure induced on $M$ by $V\in\mathcal{V}_{n}(M)$.
  • $|T|$: the varifold in $\mathcal{V}_{n}(M)$ induced by a current $T\in\mathcal{Z}_{n}(M;\mathbb{Z}_{2})$, or a countable $n$-rectifiable set $T$. In the same spirit, given a map $\Phi$ into $\mathcal{Z}_{n}(M;\mathbb{Z}_{2})$, the associated map into $\mathcal{V}_{n}(M)$ is denoted by $|\Phi|$.
  • $\operatorname{spt}(\cdot)$: the support of a current or a measure.
[…]
  • $[\mathcal{W}]\mathrel{\mathop{\ordinarycolon}}=\{[\Sigma_{i}]\}\subset\mathcal{Z}_{n}(M;\mathbb{Z}_{2})$ for a set $\mathcal{W}$ of varifolds $\{V_{i}\}\subset\mathcal{V}_{n}(M)$, each associated with a countably $n$-rectifiable set $\Sigma_{i}$.
  • $\mathbf{B}^{\nu}_{\varepsilon}(\cdot)$: the open $\varepsilon$-neighborhood of an element or a subset of the space $\mathcal{Z}_{n}(M;\nu;\mathbb{Z}_{2})$.
  • $\mathbf{B}^{\mathbf{F}}_{\varepsilon}(\cdot)$: the open $\varepsilon$-neighborhood of an element or a subset of $\mathcal{V}_{n}(M)$ under the $\mathbf{F}$-metric.
  • $\mathcal{C}(M)$: the space of Caccioppoli sets in $M$, equipped with the metric induced by the Lebesgue measure of the symmetric difference.
  • $\partial^{*}\Omega$: the reduced boundary of $\Omega\in\mathcal{C}(M)$.
  • $\nu_{\Omega}$: the inward pointing normal of $\Omega\in\mathcal{C}(M)$.
  • $\Gamma^{\infty}(M)$: the set of smooth Riemannian metrics on $M$.
  • $B_{r}(p)$: the open $r$-neighborhood of a point $p$.
  • $\mathfrak{g}(S)$: the genus of a surface $S$.
  • $\mathrm{in}(S)$ ($\mathrm{out}(S)$): the inside (outside), open region of an oriented hypersurface $S$, provided that $S$ separates $M$.
For an $m$-dimensional cube $I^{m}=\mathbb{R}^{m}\cap\{x\mathrel{\mathop{\ordinarycolon}}0\leq x_{i}\leq 1,i=1,2,\cdots,m\}$, we can give it cubical complex structures as follows.
  • $I(1,j)$: the cubical complex on $I\mathrel{\mathop{\ordinarycolon}}=[0,1]$ whose 111-cells and 0-cells are respectively
    $[0,1/3^{j}],[1/3^{j},2/3^{j}],\dots,[1-1/3^{j},1]\quad\textrm{ and }\quad[0],[1/3^{j}],[2/3^{j}],\dots,[1]\,.$
  • $I(m,j)$: the cubical complex structure
    $I(m,j)=I(1,j)\otimes\dots\otimes I(1,j)\quad(m\textrm{ times})$
    on $I^{m}$. $\alpha=\alpha_{1}\otimes\cdots\otimes\alpha_{m}$ is a $q$-cell of $I(m,j)$ if and only if each $\alpha_{i}$ is a cell in $I(1,j)$ and there are exactly $q$ 111-cells. A cell $\beta$ is a face of a cell $\alpha$ if and only if $\beta\subset\alpha$ as sets.
[…]
Given a cubical subcomplex $X$ of some $I(m,j)$, for $j^{\prime}>j$, one can refine $X$ to a cubical subcomplex
$X(j^{\prime})\mathrel{\mathop{\ordinarycolon}}=\{\sigma\in I(m,j^{\prime})\mathrel{\mathop{\ordinarycolon}}\sigma\cap|X|\neq\emptyset\}$
of $I(m,j^{\prime})$. For the sake of convenience, we will denote the refined cubical subcomplex by $X$ unless there is ambiguity.
2.2. Simon-Smith min-max theory
In his Ph.D. thesis [Smi83], written under the supervision of Simon, F. Smith combined the Almgren-Pitts theory with results from Almgren-Simon [AS79] and Meeks-Simon-Yau [MSY82], proving the existence of a minimal 2-sphere in any Riemannian 3-sphere. This theory is now known as Simon-Smith min-max theory. Later, Pitts and Rubinstein [PR86] announced, without publishing proofs, an extension of the work to more general min-max constructions in 3-dimensional manifolds, dealing with multiparameter families, surfaces of higher genus, and bounds on the Morse index. This program was eventually completed by De Lellis-Pellandini [DP10] and Ketover [Ket19]. Most recently, Wang-Zhou [WZ24] proved a multiplicity one theorem for the Simon-Smith min-max theory.
In most of the aforementioned literature, the focus is on the space of smooth surfaces, occasionally augmented with certain degenerating sets of measure zero. This setup allows the use of smooth topology, which helps simplify arguments related to regularity and genus bounds. However, in this paper, such a space is not sufficiently broad, as we need to interpolate between tori and spheres, requiring allowance for singularities both on and off the surface; see §8. We will refer to this type of singular surface as a punctate surface to distinguish it from existing notions in the literature.
In this subsection, let $(M,g)$ be a closed Riemannian 3-manifold.
Definition 2.1 (Punctate surfaces).
A closed subset $\Sigma\subset M$ is called a punctate surface provided that:
  1. (1)
    The 2-dimensional Hausdorff measure $\mathcal{H}^{2}(\Sigma)\in(0,\infty)$.
  2. (2)
    There exists a finite set $P\subset\Sigma$ such that $\Sigma\setminus P$ is an orientable, smooth, embedded surface. In this case, we call $P$ a punctate set of $\Sigma$.
  3. (3)
    Let $\Sigma_{\mathrm{iso}}$ be the set of isolated points of $\Sigma$. Then the complement of $\Sigma\backslash\Sigma_{\mathrm{iso}}$ is a disjoint union of two open regions of $M$, each having $\Sigma\backslash\Sigma_{\mathrm{iso}}$ as its topological boundary.
We denote the space of punctate surfaces in $M$ by $\mathcal{S}(M)$.
For each $\mathfrak{g}_{0}\in\mathbb{N}$ (the set of non-negative integers), $\Sigma\in\mathcal{S}(M)$ is called a genus $\mathfrak{g}_{0}$ punctate surface provided that there exists a punctate set $P$ and a sequence $r_{i}\to 0$ such that
$\mathfrak{g}(\Sigma)\mathrel{\mathop{\ordinarycolon}}=\lim_{r\to 0}\mathfrak{g}(\Sigma\setminus B_{r_{i}}(P))=\mathfrak{g}_{0}\,.$
Additionally, as in the aforementioned literature, we augment the space of punctate surfaces with degenerating sets of measure zero, and denote
$\mathcal{S}^{*}(M)\mathrel{\mathop{\ordinarycolon}}=\mathcal{S}(M)\cup\{\Sigma\subset M\mathrel{\mathop{\ordinarycolon}}\Sigma\textrm{ is a closed subset},\mathcal{H}^{2}(\Sigma)=0\}\,.$
Figure 5. A punctate surface of genus 111 with a punctate set (red dots).
Remark 2.2.
A punctate surface is simply the union of a surface with finitely many singularities and a finite set of points outside the surface (see Figure 5). Therefore, the smallest possible punctate set of $\Sigma$ is the union of the singular points of $\Sigma\setminus\Sigma_{\text{iso}}$ and $\Sigma_{\text{iso}}$. As for why we concern about isolated points of a punctate surface, and why we allow the punctate set to contain some regular points of a punctate surface, see Remark 2.7.
Remark 2.3.
For any $\Sigma\in\mathcal{S}(M)$ with a punctate set $P$, by Sard’s theorem, there exists a set of full measure $E\subset(0,\infty)$ such that for every $r\in E$, $\Sigma\setminus B_{r}(P)$ is a smooth surface with boundary. Note that the boundary is a finite union of disjoint loops.
A smooth surface with boundary can be obtained from a closed surface by removing finitely many disjoint disks. The genus of this surface is the genus of the original closed surface. Also note that when a surface is disconnected, its genus is defined as the sum of the genus of each of its connected components.
Moreover, for $r_{1}>r_{2}>0$ such that both $\Sigma\setminus B_{r_{i}}(P)$ are smooth surfaces with boundary, by [CM12, Chapter 1 §2.1 Lemma 1.5], we have the inequality on their genus,
$\mathfrak{g}(\Sigma\setminus B_{r_{1}}(P))\leq\mathfrak{g}(\Sigma\setminus B_{r_{2}}(P))\,.$
Hence, the limit
${}\lim_{E\ni r\to 0}\mathfrak{g}(\Sigma\setminus B_{r}(P))\in\mathbb{N}\cup\{\infty\}$
always exists.
In particular, the definition of $\mathfrak{g}(\Sigma)$ is independent of the choice of the sequence $r_{i}$. In addition, one can verify that the definition is independent of the choice of the punctate set $P$.
Definition 2.4 (Simon-Smith family).
Let $X$ be a cubical subcomplex of some $I(m,j)$. A map $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}^{*}(M)$ is called a Simon-Smith family, provided that:
  1. (1)
    The composition map $x\mapsto\mathcal{H}^{2}\circ\Phi(x)$ is continuous.
  2. (2)
    For any $x_{0}\in X$ and any open set $U\supset\Phi(x_{0})$, there exists a neighborhood $O\subset X$ of $x_{0}$ such that for any $x\in O$, $\Phi(x)\subset U$.
  3. (3)
    For each $x_{0}\in X$ with $\Phi(x_{0})\in\mathcal{S}(M)$, we can choose a punctate set $P_{\Phi}(x_{0})\subset\Phi(x_{0})$ such that
    $N_{P}(\Phi)\mathrel{\mathop{\ordinarycolon}}=\sup_{x\mathrel{\mathop{\ordinarycolon}}\Phi(x)\in\mathcal{S}(M)}|P_{\Phi}(x)|<\infty\,.$
  4. (4)
    For the punctate set $P_{\Phi}(x_{0})$ chosen in (3), for any $x_{0}\in X$ with $\Phi(x_{0})\in\mathcal{S}(M)$ and on any open set $U\subset\subset M\backslash P_{\Phi}(x_{0})$, $\Phi(x)\to\Phi(x_{0})$ smoothly whenever $x\to x_{0}$.
In this case, we call $X$ the parameter space of $\Phi$.
In addition, for a non-negative integer $\mathfrak{g}_{0}$, we call $\Phi$ a Simon-Smith family of genus $\leq\mathfrak{g}_{0}$, if $\Phi$ also satisfies:
  1. (5)
    For each $x\in X$ with $\Phi(x)\in\mathcal{S}(M)$, $\Phi(x)$ has genus at most $\mathfrak{g}_{0}$.
Remark 2.5.
Condition (2) implies that the family $\Phi$ is in some sense closed. More precisely, the subset
$\{(x,p)\in(X,M)\mathrel{\mathop{\ordinarycolon}}p\in\Phi(x)\}$
is closed in $X\times M$. This closedness property will be really convenient in later parts (in §9 and the proof of Proposition 10.4).
In the literature, there exist various notions of generalized family of surfaces in the Simon-Smith min-max setting. Our definition of Simon-Smith family can be viewed as a generalization of those in the following sense: Our continuity conditions (2), (3) and (4) are essentially the same as those in [DR18, Definition 2.2], but we also allow general measure-zero sets in the family. Our genus bound condition (5) generalizes the definitions in [DP10, CFS22, Fra21] without assuming a dense set of smooth surfaces in the family.
Example 2.6.
At the end of §1.2.2 we mentioned a “pinch-off process,” for the purpose of interpolating from a family of tori to a family of genus 0 surfaces; see Theorem 3.6. This is, in fact, why we need to consider surfaces with singularities, more precisely, punctate surfaces. Let us illustrate the necessity of this with the following example.
[…]
  1. (1)
    For each $p\in\mathbb{S}^{2}$, $H(0,p)=\Sigma_{p}$ and $H(1,p)=\mathbb{S}^{2}$.
  2. (2)
    $H|_{[0,1)}$ is a continuous family of tori in the smooth topology.
  3. (3)
    For any sequence $(t_{i},p_{i})\to(1,p)$, the tori $H(t_{i},p_{i})$ converge smoothly to $\mathbb{S}^{2}$ except at finitely many points in $\mathbb{S}^{2}$.
Figure 6. A torus close to a round sphere.
Indeed, suppose by contradiction such map $H$ exists. Then for each $t\in[0,1)$, consider a solid torus bundle $E_{t}$ over $\mathbb{S}^{2}$, where the fiber at each $p\in\mathbb{S}^{2}$ is the solid torus bounded by $H(t,p)$. Note that $E_{0}$ is homotopy equivalent to the unit tangent bundle to $\mathbb{S}^{2}$, and thus, has a nontrivial Euler class. However, for $t$ sufficiently close to 111, every fiber of $E_{t}$ contains the origin, implying that $E_{t}$ has a section, and thus the Euler class of $E_{t}$ is 0. This leads to a contradiction.
On the other hand, if surfaces with singularities are allowed in the interpolation, then the above task of interpolating the family $\{\Sigma_{p}\}_{p\in\mathbb{S}^{2}}$ to $\mathbb{S}^{2}$ can clearly be done, by pinching all thin necks at the origin.
Remark 2.7.
In the definition of punctate surface, we also care about isolated points of a punctate surface, and allow the punctate set to contain some regular points of a punctate surface for the following reasons:
  1. (1)
[…]
  2. (2)
    In the construction of the family $\Psi$ in §5, we allow a family to contain a sequence of smooth tori converging to a smooth sphere, smoothly except at a regular point of the sphere. Therefore, we need to consider punctate sets that contain regular points of the surfaces.
Proposition 2.8 (Simon-Smith families are sweepouts).
  1. (1)
    Every element $\Sigma$ in $\mathcal{S}^{*}(M)$ is associated with a unique 222-cycle $[\Sigma]$ in $\mathcal{Z}_{2}(M;\mathbb{Z}_{2})$. In particular, if $\mathcal{H}^{2}(\Sigma)=0$, then $[\Sigma]=0$.
  2. (2)
    Moreover, for a Simon-Smith family $\Phi$, the induced map
    $[\Phi]\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{Z}_{2}(M;\mathbf{F};\mathbb{Z}_{2})$
    is continuous with respect to the $\mathbf{F}$-metric.
Remark 2.9.
(2) implies that our Simon-Smith family induces a sweepout in the Almgren-Pitts min-max theory.
Proof of Proposition 2.8.
[…]
$[\Sigma]\mathrel{\mathop{\ordinarycolon}}=[\Sigma\setminus P]\in\mathcal{Z}_{2}(M;\mathbb{Z}_{2})\,.$
Note that the definition is independent of the choice of punctate sets $P$. For $\mathcal{H}^{2}(\Sigma)=0$, we can simply define
$[\Sigma]\mathrel{\mathop{\ordinarycolon}}=0\in\mathcal{Z}_{2}(M;\mathbb{Z}_{2})\,.$
For (2), for any $x\in X$ and any sequence $\{y_{i}\}\subset X$ with $\lim_{i}y_{i}=x$, it suffices to prove that $[\Phi(y_{i})]\to[\Phi(x)]$ with respect to the $\mathbf{F}$-metric. There are two cases: (1) $\mathcal{H}^{2}(\Phi(x))=0$, or (2) $\mathcal{H}^{2}(\Phi(x))>0$.
Case 1: $\mathcal{H}^{2}(\Phi(x))=0$
In this case, by Definition 2.4 (1),
$\mathbf{M}([\Phi(y_{i})])=\mathcal{H}^{2}(\Phi(y_{i}))\to\mathcal{H}^{2}(\Phi(x))=0\,,$
and thus, $[\Phi(y_{i})]\to 0=[\Phi(x)]$ with respect to the $\mathbf{F}$-metric.
Case 2: $\mathcal{H}^{2}(\Phi(x))>0$
In this case, we can assume that every $\Phi(y_{i})\in\mathcal{S}(M)$ and let $P\mathrel{\mathop{\ordinarycolon}}=P_{\Phi}(x)$ defined in Definition 2.4 (3). For any $\varepsilon>0$, by the facts that $\mathcal{H}^{2}(\Phi(x))<\infty$, that $\Phi(x)$ is closed, and by Sard’s theorem, there exists $r>0$ such that
(2.1) $\mathcal{H}^{2}(\Phi(x)\cap\overline{B}_{2r}(P))<\frac{\varepsilon}{4}\,,$
(2.2) $\mathcal{H}^{3}(\overline{B}_{2r}(P))<\varepsilon/4\,,$
and $\partial B_{2r}(P)$ intersects transversally with $\Phi(x)$ and each $\Phi(y_{i})$, where each intersection is a finite union of loops.
By Definition 2.4 (4), $\Phi(y_{i})\to\Phi(x)$ smoothly outside $\overline{B}_{r}(P)$. In particular, we have:
  • (2.3) $\limsup_{i}\mathcal{F}([\Phi(y_{i})],[\Phi(x)])\leq\mathcal{H}^{3}(\overline{B}_{2r}(P))\leq\frac{\varepsilon}{4}\,.$
  • Outside $\overline{B}_{2r}(P)$, we have
    (2.4) $\lim_{i}\mathbf{F}(|\Phi(y_{i})\setminus\overline{B}_{2r}(P)|,|\Phi(x)\setminus\overline{B}_{2r}(P)|)=0\,.$
    In particular, $\lim_{i}\mathcal{H}^{2}(\Phi(y_{i})\setminus\overline{B}_{2r}(P))=\mathcal{H}^{2}(\Phi(x)\setminus\overline{B}_{2r}(P))$.
  • Inside $\overline{B}_{2r}(P)$,
    (2.5) $\quad\limsup_{i}\mathcal{H}^{2}(\Phi(y_{i})\cap\overline{B}_{2r}(P))$
    $=\limsup_{i}\left(\mathcal{H}^{2}(\Phi(y_{i}))-\mathcal{H}^{2}(\Phi(y_{i})\setminus\overline{B}_{2r}(P))\right)$
    $=\lim_{i}\mathcal{H}^{2}(\Phi(y_{i}))-\lim_{i}\mathcal{H}^{2}(\Phi(y_{i})\setminus\overline{B}_{2r}(P))$
    $=\mathcal{H}^{2}(\Phi(x))-\mathcal{H}^{2}(\Phi(x)\setminus\overline{B}_{2r}(P))$
    $=\mathcal{H}^{2}(\Phi(x)\cap\overline{B}_{2r}(P))<\frac{\varepsilon}{4}\,.$
Note that by the triangle inequality, we have
$\quad\mathbf{F}(|\Phi(y_{i})|,|\Phi(x)|)$
$=\mathbf{F}(|\Phi(y_{i})|,|\Phi(x)|)+\mathcal{F}(|\Phi(y_{i})|,|\Phi(x)|)$
$\leq\mathbf{F}(|\Phi(y_{i})|,|\Phi(y_{i})\setminus\overline{B}_{2r}(P)|)+\mathbf{F}(|\Phi(x)\setminus\overline{B}_{2r}(P)|,|\Phi(x)|)$
$\quad+\mathbf{F}(|\Phi(y_{i})\setminus\overline{B}_{2r}(P)|,|\Phi(x)\setminus\overline{B}_{2r}(P)|)$
$=\mathbf{F}(|\Phi(y_{i})\cap\overline{B}_{2r}(P)|,0)+\mathbf{F}(0,|\Phi(x)\cap\overline{B}_{2r}(P)|)$
$\quad+\mathbf{F}(|\Phi(y_{i})\setminus\overline{B}_{2r}(P)|,|\Phi(x)\setminus\overline{B}_{2r}(P)|)$
$=\mathcal{H}^{2}(\Phi(y_{i})\cap\overline{B}_{2r}(P))+\mathcal{H}^{2}(\Phi(x)\cap\overline{B}_{2r}(P))$
$\quad+\mathbf{F}(|\Phi(y_{i})\setminus\overline{B}_{2r}(P)|,|\Phi(x)\setminus\overline{B}_{2r}(P)|)\,.$
Applying (2.1), (2.3), (2.4) and (2.5), for any $\varepsilon>0$, we obtain
$\limsup_{i}\mathbf{F}([\Phi(y_{i})],[\Phi(x)])$ $\leq\limsup_{i}\mathcal{F}([\Phi(y_{i})],[\Phi(x)])$
$\qquad+\limsup_{i}\mathbf{F}(|\Phi(y_{i})|,|\Phi(x)|)$
$\leq\varepsilon\,,$
and thus, $\lim_{i}\mathbf{F}([\Phi(y_{i})],[\Phi(x)])=0$. ∎
Definition 2.10 (Homotopy class).
Two Simon-Smith families $\Phi_{0}$ and $\Phi_{1}$, which are parametrized by the same parameter space $X$, are said to be homotopic to each other if there exists a continuous map
$\varphi\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\operatorname{Diff}^{\infty}(M)$
such that:
  1. (1)
    $\varphi(0,x)=\operatorname{Id}$ for all $x\in X$.
  2. (2)
    $\varphi(1,x)(\Phi_{0}(x))=\Phi_{1}(x)$ for all $x\in X$.
Note that $\textrm{Diff}^{\infty}(M)$ denotes the diffeomorphism group of $M$ equipped with the $C^{\infty}$ topology.
The set of all families homotopic to a Simon-Smith family $\Phi$ is called the homotopy class associated to $\Phi$, and is denoted by $\Lambda(\Phi)$.
Remark 2.11.
The family
$H\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}^{*}(M),\quad(t,x)\mapsto\varphi(t,x)(\Phi(x))$
is also a Simon-Smith family. In particular, for every $t\in[0,1]$,
$\Phi_{t}\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}^{*}(M),\quad x\mapsto\varphi(t,x)(\Phi(x))$
is a Simon-Smith family in $\Lambda(\Phi_{0})$.
Furthermore, for any choice of punctate sets
$P_{\Phi}\mathrel{\mathop{\ordinarycolon}}\{x\in X\mathrel{\mathop{\ordinarycolon}}\Phi(x)\in\mathcal{S}(M)\}\to M$
for $\Phi$ satisfying the conditions (3) and (4) in Definition 2.4, any isotopy $\varphi$ would induce a choice of $P_{\Phi^{\prime}}(x)\mathrel{\mathop{\ordinarycolon}}=\varphi(1,x)\circ P_{\Phi}(x)$ for every $\Phi^{\prime}\mathrel{\mathop{\ordinarycolon}}=\varphi(1,\cdot)\circ\Phi\in\Lambda(\Phi)$. Consequently, we may assume that $N_{P}<\infty$ is a constant within the homotopy class $\Lambda(\Phi)$.
In addition, if $\Phi$ is of genus $\leq\mathfrak{g}_{0}$, then so is every $\Phi^{\prime}\in\Lambda(\Phi)$.
Definition 2.12 (Simon-Smith width).
Given a homotopy class $\Lambda$, its width is defined by
$\mathbf{L}(\Lambda)\mathrel{\mathop{\ordinarycolon}}=\inf_{\Phi\in\Lambda}\sup_{x\in X}\mathcal{H}^{2}(\Phi(x))\,.$
Definition 2.13 (Minimizing sequence and min-max sequence).
A sequence $\{\Phi_{i}\}$ in $\Lambda$ is said to be minimizing if
$\lim_{i\to\infty}\sup_{x\in X}\mathcal{H}^{2}(\Phi_{i}(x))=\mathbf{L}(\Lambda)\,.$
If $\{\Phi_{i}\}$ is a minimizing sequence in $\Lambda$ and $\{x_{i}\}\subset X$ satisfies
$\lim_{i\to\infty}\mathcal{H}^{2}(\Phi_{i}(x_{i}))=\mathbf{L}(\Lambda)\,,$
then $\{\Phi_{i}(x_{i})\}$ is called a min-max sequence. For a minimizing sequence $\{\Phi_{i}\}$, we define its critical set $\mathbf{C}(\{\Phi_{i}\})$ to be the set of all subsequential varifold-limit of its min-max sequences:
$\mathbf{C}(\{\Phi_{i}\})\mathrel{\mathop{\ordinarycolon}}=\{V=\lim_{j}|\Phi_{i_{j}}(x_{j})|\mathrel{\mathop{\ordinarycolon}}x_{j}\in X,\|V\|(M)=\mathbf{L}(\Lambda)\}\,.$
And $\{\Phi_{i}\}$ is called pulled-tight if every varifold in $\mathbf{C}(\{\Phi_{i}\})$ is stationary.
Definition 2.14.
In a closed Riemannian 333-manifold $(M,g)$, for $L>0$, let $\mathcal{W}_{L}(M,g)$ be the set of all varifolds $W\in\mathcal{V}_{2}(M)$, with $\|W\|(M)=L$, of the form
$W=m_{1}|\Gamma_{1}|+\cdots+m_{l}|\Gamma_{l}|\,,$
[…]
For a nonnegative integer $\mathfrak{g}_{0}\geq 0$, let $\mathcal{W}_{L,\leq\mathfrak{g}_{0}}(M,g)\subset\mathcal{W}_{L}(M,g)$ consist of all the varifolds satisfying the genus bound
(2.6) $\sum_{j\in I_{O}}m_{j}\mathfrak{g}(\Gamma_{j})+\frac{1}{2}\sum_{j\in I_{N}}m_{j}(\mathfrak{g}(\Gamma_{j})-1)\leq\mathfrak{g}_{0}\,,$
where $\Gamma_{j}$ is orientable if $j\in I_{O}$ and non-orientable if $j\in I_{N}$.
For simplicity, when the ambient manifold $(M,g)$ is clear from context, we may use the notations $\mathcal{W}_{L}$ and $\mathcal{W}_{L,\leq\mathfrak{g}_{0}}$ without further specification.
Remark 2.15.
Note that the genus of a non-orientable surface $\Sigma$ is the number of cross-caps needed to be attached to a two-sphere in order to obtain a surface homeomorphic to $\Sigma$.
Theorem 2.16 (Simon-Smith min-max theorem).
Given an orientable closed Riemannian 333-manifold $(M,g)$, a cubical subcomplex $X$ of some $I(m,k)$, a Simon-Smith family $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}^{*}(M)$ of genus $\leq\mathfrak{g}_{0}$, and a positive real number $r>0$, if $L\mathrel{\mathop{\ordinarycolon}}=\mathbf{L}(\Lambda(\Phi))>0$, then there exists a minimizing sequence $\{\Phi_{i}\}$ in $\Lambda(\Phi)$ such that:
  1. (1)
    The sequence $\{\Phi_{i}\}$ is pulled-tight and moreover,
    $\mathbf{C}(\{\Phi_{i}\})\cap\mathcal{W}_{L,\leq\mathfrak{g}_{0}}\neq\emptyset\,.$
  2. (2)
    There exists a $W\in\mathcal{W}_{L,\leq\mathfrak{g}_{0}}$ such that every connected component $\Gamma_{j}$ in $W$ satisfies:
    1. (a)
      If $\Gamma_{j}$ is unstable and two-sided, then $m_{j}=1$.
    2. (b)
      If $\Gamma_{j}$ is one-sided, then its connected double cover is stable.
  3. (3)
    Furthermore, there exists $\eta>0$ such that for all sufficiently large $i$,
    $\mathcal{H}^{2}(\Phi_{i}(x))\geq L-\eta\implies|\Phi_{i}(x)|\in\mathbf{B}^{\mathbf{F}}_{r}(\mathcal{W}_{L,\leq\mathfrak{g}_{0}})\,.$
This min-max theorem essentially builds upon the foundational results in [Smi83, DP10, Ket19, MN21, WZ24]. We postpone the detailed proof of this theorem to §4.
2.3. Relative Simon-Smith min-max theory
[…]
Definition 2.17 (Relative homotopy class).
Two Simon-Smith families $\Phi_{0}$ and $\Phi_{1}$ parametrized by the same parameter space $X$ with $\Phi_{0}|_{Z}=\Phi_{1}|_{Z}$ are said to be homotopic relative to $\Phi_{0}|_{Z}$ to each other if there exists a continuous map
$\varphi\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\operatorname{Diff}^{\infty}(M)$
such that:
  1. (1)
    $\varphi(0,x)=\operatorname{Id}$ for all $x\in X$.
  2. (2)
    $\varphi(t,z)=\operatorname{Id}$ for all $t\in[0,1]$ and $z\in Z$.
  3. (3)
    $\varphi(1,x)(\Phi_{0}(x))=\Phi_{1}(x)$ for all $x\in X$.
[…]
Definition 2.18 (Relative Simon-Smith min-max width).
Given a $(X,Z)$-relative homotopy class $\Lambda_{Z}$, its width is defined by
$\mathbf{L}(\Lambda_{Z})\mathrel{\mathop{\ordinarycolon}}=\inf_{\Phi\in\Lambda_{Z}}\sup_{x\in X}\mathcal{H}^{2}(\Phi(x))\,.$
Definition 2.19 (Minimizing sequence and min-max sequence for Relative Simon-Smith min-max).
A sequence $\{\Phi_{i}\}$ in $\Lambda_{Z}$ is said to be minimizing if
$\lim_{i\to\infty}\sup_{x\in X}\mathcal{H}^{2}(\Phi_{i}(x))=\mathbf{L}(\Lambda_{Z})\,.$
If $\{\Phi_{i}\}$ is a minimizing sequence and $\{x_{i}\}\subset X$ satisfies
$\lim_{i\to\infty}\mathcal{H}^{2}(\Phi_{i}(x_{i}))=\mathbf{L}(\Lambda_{Z})\,,$
then $\{\Phi_{i}(x_{i})\}$ is called a min-max sequence. For a minimizing sequence $\{\Phi_{i}\}$, we define its critical set $\mathbf{C}(\{\Phi_{i}\})$ to be the set of all subsequential varifold-limit of its min-max sequences:
$\mathbf{C}(\{\Phi_{i}\})\mathrel{\mathop{\ordinarycolon}}=\{V=\lim_{j}|\Phi_{i_{j}}(x_{j})|\mathrel{\mathop{\ordinarycolon}}x_{j}\in X,\|V\|(M)=\mathbf{L}(\Lambda_{Z})\}\,.$
And $\{\Phi_{i}\}$ is called pulled-tight if every varifold in $\mathbf{C}(\{\Phi_{i}\})$ is stationary.
Lemma 2.20.
For a Simon-Smith family $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}^{*}(M)$ and a subcomplex $Z\subset X$, we have
$\mathbf{L}(\Lambda(\Phi))\leq\mathbf{L}(\Lambda_{Z}(\Phi))\,.$
Proof.
This follows immediately from the fact that $\Lambda_{Z}(\Phi)\subset\Lambda(\Phi)$. ∎
Theorem 2.21 (Relative Simon-Smith min-max theorem).
Given an orientable closed Riemannian 333-manifold $(M,g)$, a pair of cubical subcomplexes $Z\subset X$ of some $I(m,k)$, a Simon-Smith family $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}^{*}(M)$ of genus $\leq\mathfrak{g}_{0}$, and a positive real number $r>0$, if
$L=\mathbf{L}(\Lambda_{Z}(\Phi))>\sup_{z\in Z}\mathcal{H}^{2}(\Phi(z))\,,$
then there exists a minimizing sequence $\{\Phi_{i}\}$ in $\Lambda_{Z}(\Phi)$ such that conclusions (1), (2) and (3) in Theorem 2.16 still hold.
This relative min-max theorem can be proved in much the same way as 2.16, carefully following the steps of that proof. A proof is also provided in §4.
2.4. Almgren-Pitts min-max theory
In 1962, after F.J. Almgren completed his thesis [Alm62], H. Federer suggested that he develop a Morse theory for minimal surfaces. In [Alm65], Almgren initiated a program to develop a geometric version of the calculus of variations in the large, aimed at proving the existence of critical points for the area functional. This program is based on what is nowadays called the Almgren isomorphism theorem [Alm62, Alm65] (see also [LMN18, §2.5]). In particular, when equipped with the flat topology, the modulo 222 cycle space $\mathcal{Z}_{n}(M;\mathbb{Z}_{2})$ is weakly homotopy equivalent to $\mathbb{R}\mathbb{P}^{\infty}$. The rich topological structure of $\mathcal{Z}_{n}(M;\mathbb{Z}_{2})$ makes Morse theory possible.
In the following, we denote the cohomology ring of $\mathcal{Z}_{n}(M;\mathbb{Z}_{2})$ by $\mathbb{Z}_{2}[\bar{\lambda}]$, where $\bar{\lambda}\in H^{1}(\mathcal{Z}_{n}(M;\mathbb{Z}_{2}),\mathbb{Z}_{2})$ is the generator. Let $\mathcal{P}_{p}$ be the set of all $\mathbf{F}$-continuous maps $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{Z}_{2}(M;\mathbf{F};\mathbb{Z}_{2})$, where $X$ is a finite simplicial complex, such that $\Phi^{*}(\bar{\lambda}^{p})\neq 0$. Elements of $\mathcal{P}_{p}$ are called $p$-sweepouts. Note that every finite cubical complex is homeomorphic to a finite simplicial complex and vice versa (see [BP02, §4]). So in the above notion of $p$-sweepouts, we may as well require $X$ to be a finite cubical complex.
Definition 2.22 (Almgren-Pitts $p$-width).
[…]
$\omega_{p}(M,g)\mathrel{\mathop{\ordinarycolon}}=\inf_{\Phi\in\mathcal{P}_{p}}\sup_{x\in\mathrm{dmn}(\Phi)}\mathbf{M}(\Phi(x))\,.$
Remark 2.23.
There is an equivalent definition of $p$-widths from [MN21, Remark 5.7]: First, an $\mathcal{F}$-continuous map $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{Z}_{n}(M;\mathbb{Z}_{2})$ is said to have no concentration of mass if
$\lim_{r\to 0}\sup_{x\in X,q\in M}\mathinner{\!\left\lVert\Phi(x)\right\rVert}(B_{r}(q))=0\,.$
Then, when defining the $p$-width, instead of using the collection $\mathcal{P}_{p}$, we use the collection of all $\mathcal{F}$-continuous maps $\Phi$ with no concentration of mass such that $\Phi^{*}(\bar{\lambda}^{p})\neq 0$.
The interpolation results therein imply that the two definitions yield the same $p$-widths.
As in Simon-Smith min-max theory, we can define the notion of minimizing sequences, critical sets, and min-max sequences for the $p$-widths.
In this paper, the only result we require from the Almgren-Pitts min-max theory is that the 555-width of the unit 333-sphere $(S^{3},\bar{g})$ is greater than $4\pi$. In fact, C. Nurser [Nur16] proved that
$\omega_{5}(S^{3},\bar{g})=2\pi^{2}\,.$
3. Proof of Theorem A
In this section, we will prove our main theorem, Theorem A: Given a 333-sphere $(S^{3},g_{0})$ of positive Ricci curvature, it contains at least 555 embedded minimal tori. The proof consists of five key steps:
  1. (1)
    Define a 9-parameter Simon-Smith family $\Psi$ of genus 1$\leq$ 1≤ 1, building upon the 555-parameter canonical family discovered by Marques-Neves in their proof of the Willmore conjecture [MN14].
  2. (2)
    Repeatedly apply the Simon-Smith min-max theory to $\Psi$. Namely, each time we detect a minimal surface, which must be either a multiplicity one minimal sphere or torus by Wang-Zhou’s multiplicity one theorem [WZ24], we remove the region in the family near the minimal surface (which is like a cap), resulting in a new family of maximal area strictly less than the area of the minimal surface. We repeat the min-max process on this new family.
  3. (3)
    Assemble pieces we obtained in step 2 to get a 999-parameter family $\Xi$, which is, in a certain sense, homotopic to $\Psi$.
  4. (4)
    Let $N$ be the number of embedded minimal tori obtained via min-max in Step 2. Then we show that the parameter space $\mathrm{dmn}(\Xi)$ of $\Xi$ can be decomposed into a union of subsets $D_{0},D_{1},\cdots,D_{N}$, such that the restriction $\Xi|_{D_{0}}$ is not a 5-sweepout, while for each $i=1,\cdots,N$, the image of $H_{1}(D_{i};\mathbb{Z}_{2})$ mapped into $H_{1}(\mathrm{dmn}(\Xi);\mathbb{Z}_{2})$ is trivial.
  5. (5)
    Show that $N\geq 5$ by analyzing the cohomology ring of $\mathrm{dmn}(\Xi)$ and applying a Lyusternik-Schnirelmann argument.
3.1. A 9-parameter family
Let us first define a 999-parameter family $\Psi$ of surfaces in the unit 333-sphere. It is worth noting that F. C. Marques previously considered a similar family to prove that the 888-width of the unit 333-sphere is $2\pi^{2}$ [Mar23].
For simplicity, throughout §3.1, we denote the unit 333-sphere by $\mathbb{S}^{3}$, and the unit open 444-ball in $\mathbb{R}^{4}$ by $\mathbb{B}^{4}$. We fix an unoriented Clifford torus $\Sigma_{0}$ in $\mathbb{S}^{3}$.
Consider for each $v\in\mathbb{B}^{4}$ the conformal diffeomorphism $F_{v}\mathrel{\mathop{\ordinarycolon}}\mathbb{S}^{3}\to\mathbb{S}^{3}$ given by
(3.1) $F_{v}(x)=\frac{1-|v|^{2}}{|x-v|^{2}}(x-v)-v\,,$
[…]
Next, we choose an orientation for $\Sigma_{0}$. This gives an inward normal direction for $\Sigma_{0}$, and consequently, for each $F_{v}(\Sigma_{0})$. With this orientation, we can define the signed distance $d_{v}\mathrel{\mathop{\ordinarycolon}}\mathbb{S}^{3}\to\mathbb{R}$ to $F_{v}(\Sigma_{0})$, which is positive outside $F_{v}(\Sigma_{0})$ and negative inside.
Then we define the following continuous family of 222-cycles: For each pair $(v,t)\in\mathbb{B}^{4}\times[-\pi,\pi]$, let
(3.2) $\Sigma(v,t)\mathrel{\mathop{\ordinarycolon}}=\partial\{x\in\mathbb{S}^{3}\mathrel{\mathop{\ordinarycolon}}d_{v}(x)<t\}\in\mathcal{Z}_{2}(\mathbb{S}^{3};\mathbb{Z}_{2})\,.$
[…]
$\Phi_{5}\mathrel{\mathop{\ordinarycolon}}{\mathbb{RP}}^{5}\to\mathcal{Z}_{2}(\mathbb{S}^{3};\mathbb{Z}_{2})\,,$
which he showed to be a 5-sweepout by proving $\Phi_{5}^{*}(\bar{\lambda}^{5})\neq 0$ and $\Phi_{5}$ has no concentration of mass ($\bar{\lambda}$ is the non-trivial element of $H^{1}(\mathcal{Z}_{2}(\mathbb{S}^{3};\mathbb{Z}_{2});\mathbb{Z}_{2})$). The details of this procedure was carried out carefully in  [Nur16, §3.4.2], but for the sake of completeness, we will also provide the details in §5.
Now, for each oriented Clifford torus $\Sigma$, one can repeat the above procedure, with $\Sigma$ in place of $\Sigma_{0}$, to obtain a map
$\Phi^{\Sigma}_{5}\mathrel{\mathop{\ordinarycolon}}{\mathbb{RP}}^{5}\to\mathcal{Z}_{2}(\mathbb{S}^{3};\mathbb{Z}_{2})\,.$
Thus, if we let $\widetilde{\mathcal{C}}$ be the space of oriented Clifford tori, we immediately obtain a map
(3.3) ${\mathbb{RP}}^{5}\times\widetilde{\mathcal{C}}\to\mathcal{Z}_{2}(\mathbb{S}^{3};\mathbb{Z}_{2})\,.$
However, the topology of the parameter space of this family would not satisfy our need. We would want to identify Clifford tori of opposite orientations. Thus, we wish to identify the subfamily $\Phi^{\Sigma}_{5}$ and $\Phi^{-\Sigma}_{5}$ of the family (3.3), where $-\Sigma$ is $\Sigma$ with an opposite orientation. We claim that this can be done. More precisely, let us consider instead the space $\mathcal{C}$ of unoriented Clifford tori, which is topologically an ${\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$ [Nur16, §3.4.3]. We will prove in §5 the following theorem.
Theorem 3.1.
There exists an ${\mathbb{RP}}^{5}$-bundle $Y$ over ${\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$, and a Simon-Smith family of genus 1$\leq$ 1≤ 1
$\Psi\mathrel{\mathop{\ordinarycolon}}Y\to\mathcal{S}^{*}(S^{3})\,,$
satisfying the following property: If $\Sigma$ is an oriented Clifford torus and $b\in{\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$ corresponds to $\{\Sigma,-\Sigma\}$, then there exists a homeomorphism $f\mathrel{\mathop{\ordinarycolon}}Y_{b}\to{\mathbb{RP}}^{5}$ such that $\Phi^{\Sigma}_{5}\circ f=[\Psi|_{Y_{b}}]$.
Note that in the above, $Y_{b}$ denotes the fiber at $b$, $\Psi|_{Y_{b}}$ the restriction of $\Psi$ on $Y_{b}$, and $[\Psi|_{Y_{b}}]$ the induced map into $\mathcal{Z}_{2}(\mathbb{S}^{3};\mathbb{Z}_{2})$. Also note that since we are in the Simon-Smith setting, $\Psi$ maps into the set $\mathcal{S}^{*}(S^{3})$ of punctate surfaces instead of the cycle space $\mathcal{Z}_{2}(\mathbb{S}^{3};\mathbb{Z}_{2})$.
Remark 3.2.
In [Nur16, §3.4.3] C. Nurser introduced a 999-parameter family which is parametrized by ${\mathbb{RP}}^{5}\times{\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$. However, that family is actually not well-defined. The issue is that to define an ${\mathbb{RP}}^{5}$-family as described above, one needs to specify an orientation for the Clifford torus. There is not a continuous way to assign an orientation to each unoriented Clifford torus in ${\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$.
3.2. Repetitive min-max
Let $(S^{3},g_{0})$ be the 333-sphere given in our main theorem, Theorem A. To prove the theorem, we can assume $(S^{3},g_{0})$ has only finitely many embedded minimal tori; otherwise, we would already have at least five embedded minimal tori and the proof would be complete.
In §3.1, we obtained a 999-parameter Simon-Smith family $\Psi$ of genus 1$\leq$ 1≤ 1 on the unit 333-sphere. But this family can also be viewed as a Simon-Smith family in $S^{3}$ with any metric.
We will soon run the min-max process repeatedly starting from $\Psi$. To ensure that the process terminates in finitely many steps and that all the minimal surfaces detected have multiplicity one, we need to perturb $g_{0}$ to a new metric $g^{\prime}$ with the following properties:
Proposition 3.3.
For a Riemannian metric $g$ of positive Ricci curvature on $S^{3}$, if there are finitely many embedded minimal tori in $(S^{3},g)$, then there exists another metric $g^{\prime}$ of positive Ricci curvature such that:
  1. (1)
    $(S^{3},g)$ and $(S^{3},g^{\prime})$ have the same number of embedded minimal tori.
  2. (2)
    $(S^{3},g^{\prime})$ has finitely many embedded minimal spheres, each of which is non-degenerate.
  3. (3)
    For any embedded minimal sphere $S$ in $(S^{3},g^{\prime})$ and any positive integers $m$, the value $m\cdot\operatorname{area}_{g^{\prime}}(S)$ can never equal the $g^{\prime}$-area of an embedded minimal torus in $(S^{3},g^{\prime})$.
  4. (4)
    For a pair of distinct embedded minimal spheres $S_{1},S_{2}$ in $(S^{3},g^{\prime})$ and a pair of positive integers $m_{1},m_{2}$, we have
    $m_{1}\cdot\operatorname{area}_{g^{\prime}}(S_{1})\neq m_{2}\cdot\operatorname{area}_{g^{\prime}}(S_{2})\,.$
The proof of the proposition above will be postponed to §6. In fact, from the proof, the metric $g^{\prime}$ can be made arbitrarily close to $g$ in $C^{\infty}$, and all the $g^{\prime}$-minimal tori coincide with the $g$-minimal tori. However, we will not need these facts.
Now, let us apply the above proposition to $(S^{3},g_{0})$ to obtain a modified metric $g^{\prime}$ satisfying the above properties. To prove Theorem A, it suffices to show that $(S^{3},g^{\prime})$ has at least five embedded minimal tori. Therefore, in what follows, we will focus on $(S^{3},g^{\prime})$. We will need the following useful lemma.
Lemma 3.4.
In $(S^{3},g^{\prime})$, for any $L>0$, suppose that $\mathcal{W}_{L,\leq 1}$ (defined using (2.6)) has a varifold associated with a multiplicity-one minimal surface $\Sigma$.
  1. (1)
    If $\Sigma$ is a minimal torus, then every varifold in $\mathcal{W}_{L,\leq 1}$ is associated with a multiplicity-one minimal torus.
  2. (2)
    If $\Sigma$ is a multiplicity-one minimal sphere, then
    $\mathcal{W}_{L,\leq 1}=\{|\Sigma|\}\,.$
In particular, $\mathcal{W}_{L,\leq 1}$ consists of finitely many varifolds, each associated with a multiplicity-one minimal surface.
Proof.
Since $(S^{3},g^{\prime})$ has positive Ricci curvature, the Frankel property holds, i.e., every pair of connected embedded minimal surfaces intersects with each other. Consequently, every varifold in $\mathcal{W}_{L,\leq 1}$ is associated with a surface, which is either a minimal torus with multiplicities, or a minimal sphere with multiplicities.
For (1), by (3) of Proposition 3.3, every varifold in $\mathcal{W}_{L,\leq 1}$ is associated with a multiplicity-one minimal torus.
For (2), by (3) of Proposition 3.3 again, every varifold in $\mathcal{W}_{L,\leq 1}$ is associated with a minimal sphere with multiplicities. By (4) of Proposition 3.3, we have $\mathcal{W}_{L,\leq 1}=\{[\Sigma]\}$. ∎
We now apply the min-max process repeatedly, starting with our canonical family $\Psi$, treated as a family in $(S^{3},g^{\prime})$. We choose some $d_{0}>0$ to be less than the distance $\mathbf{F}(\Sigma,\Sigma^{\prime})$ for any two minimal surfaces $\Sigma,\Sigma^{\prime}$ of genus 0 or 1 in $(S^{3},g^{\prime})$ ($\mathbf{F}$ denotes the distance for currents). Note, this is possible as we assumed $(S^{3},g^{\prime})$ has only finitely many minimal tori and spheres. (Though, in §10, we will need to choose an even smaller $d_{0}$, depending on $g^{\prime}$ only; see the paragraph following the proof of Proposition 10.1.)
3.2.1. First stage.
In $(S^{3},g^{\prime})$, applying Theorem 2.16 to $\Psi$ with $r=d_{0}$, we obtain a varifold $V^{1}$ in $\mathcal{W}^{1}\mathrel{\mathop{\ordinarycolon}}=\mathcal{W}_{\mathbf{L}(\Lambda(\Psi)),\leq 1}$, where $V^{1}$ is induced by a multiplicity-one minimal surface of genus 0 or 111. By Lemma 3.4, we have the following dichotomy:
  • Case 1
    $\mathcal{W}^{1}$ consists finitely many finitely many varifolds, each associated to a multiplicity-one minimal torus.
  • Case 2
    $\mathcal{W}^{1}=\{V^{1}\}$ where $\operatorname{spt}\|V^{1}\|$ has genus 0.
Let us investigate case 1 first.
Case 1.
From Theorem 2.16 (3) it follows immediately that there exists a $\delta>0$ and some $\Psi^{\prime}\in\Lambda(\Psi)$ such that
$\mathbf{M}(\Psi^{\prime}(x))\geq\mathbf{L}(\Lambda(\Psi))-\delta\;\;\Rightarrow\;\;|\Psi^{\prime}(x)|\in\mathbf{B}^{\mathbf{F}}_{d_{0}}(\mathcal{W}^{1})\,.$
However, this is not enough for us to prove Theorem 3.16. A stronger result is required.
Theorem 3.5.
Let $(M,g)$ be an orientable closed Riemannian 333-manifold, $\Lambda$ be a homotopy class of a Simon-Smith families of genus $\leq\mathfrak{g}_{0}$ such that $L=\mathbf{L}(\Lambda)>0$. Suppose that $\mathcal{W}_{L,\leq\mathfrak{g}_{0}}$ consists of finitely many varifolds, each associated with a connected multiplicity-one minimal surface. Then for any $r>0$, there exists $\eta>0$ and $\Phi\in\Lambda$ such that
$\mathbf{M}(\Phi(x))\geq L-\eta\implies[\Phi(x)]\in\mathbf{B}^{\mathbf{F}}_{r}([\mathcal{W}_{L,\leq\mathfrak{g}_{0}}])\,.$
The notation $[\mathcal{W}_{L,\leq\mathfrak{g}_{0}}]$ used here is as introduced in §2, and $\mathbf{B}^{\mathbf{F}}_{r}([\mathcal{W}_{L,\leq\mathfrak{g}_{0}}])$ denotes the $r$-neighborhood of $[\mathcal{W}_{L,\leq\mathfrak{g}_{0}}]$ in $\mathcal{Z}_{2}(M;\mathbf{F};\mathbb{Z}_{2})$. We will prove Theorem 3.5 in §7.
Applying the above theorem with $(M,g)=(S^{3},g^{\prime})$, $\Lambda=\Lambda(\Psi)$, $\mathfrak{g}_{0}=1$ and $r=d_{0}$, we obtain $\eta>0$ and $\Phi\in\Lambda(\Psi)$ satisfying the aforementioned property. For later reference, we denote $\delta_{1}\mathrel{\mathop{\ordinarycolon}}=\eta$ and $\widetilde{\Psi}^{1}\mathrel{\mathop{\ordinarycolon}}=\Phi$.
By refining the cubical complex structure of the parameter space $Y$, we can obtain a 999-dimensional subcomplex $C^{1}$ of $Y$, such that
(3.4) $\mathbf{M}(\widetilde{\Psi}^{1}(x))\geq\mathbf{L}(\Lambda(\Psi))-\delta_{1}/2\implies x\in C^{1},$
and
(3.5) $x\in C^{1}\implies[\widetilde{\Psi}^{1}(x)]\in\mathbf{B}^{\mathbf{F}}_{d_{0}}([\mathcal{W}^{1}]).$
Roughly speaking, $C^{1}$ is the “cap” where $\widetilde{\Psi}^{1}$ has large area and is close to some embedded minimal tori.
Since $(S^{3},g^{\prime})$ has only finitely many embedded minimal tori, by the definition of $d_{0}$, and by further refining $Y$, we can assume that $C^{1}$ can be decomposed into a disjoint union of 999-dimensional subcomplexes $C^{1}_{1},\cdots,C^{1}_{n_{1}}$ such that: For each $j=1,\cdots,n_{1}$ there is a distinct embedded minimal torus $T^{1}_{j}\in\mathcal{S}^{*}(S^{3})$, with $|T^{1}_{j}|\in\mathcal{W}^{1}$, such that the image
$[\widetilde{\Psi}^{1}](C^{1}_{j})\subset\mathbf{B}^{\mathbf{F}}_{d_{0}}([T^{1}_{j}])\,.$
Finally, we define
$\Psi^{1}\mathrel{\mathop{\ordinarycolon}}=\widetilde{\Psi}^{1}|_{\overline{Y\backslash C^{1}}}\,.$
By (3.4), $\Psi^{1}$ satisfies that
(3.6) $\sup\mathcal{H}^{2}\circ\Psi^{1}<\mathbf{L}(\Lambda(\Psi)),$
Later when we apply the min-max process for the second time, we will start with $\Psi^{1}$, and the area upper bound guarantees new min-max minimal surfaces will be detected. For notational convenience later, let us denote $\widetilde{Y}^{1}\mathrel{\mathop{\ordinarycolon}}=\mathrm{dmn}(\widetilde{\Psi}^{1})=Y$ and $Y^{1}=\overline{Y\backslash C^{1}}=\mathrm{dmn}(\Psi^{1})$.
Case 2.
In this case, the goal is once again to construct a map $\Psi^{1}$ that satisfies the bound (3.6). Additionally, we want $\Psi^{1}$, when restricted to the boundary of the removed cap, to form a Simon-Smith family of genus 0. To achieve this, we will need to apply an important interpolation theorem. But first, let us introduce some concepts.
Figure 7. The figure on the left shows a map $f\mathrel{\mathop{\ordinarycolon}}X^{\prime}\to X$. The figure on the right shows the mapping cylinder $W\mathrel{\mathop{\ordinarycolon}}=M_{f}$, where the blue lines describe the deformation retraction of $W$ onto $\partial_{0}W$ induced by the map $F_{W}$.
Given two topological spaces $X$ and $X^{\prime}$, and a continuous map $f\mathrel{\mathop{\ordinarycolon}}X^{\prime}\to X$, the mapping cylinder of $f$ is defined by
$M_{f}\mathrel{\mathop{\ordinarycolon}}=(([0,1]\times X^{\prime})\sqcup X)/\sim\,,$
where $(0,x)\sim f(x)$ for all $x\in X^{\prime}$.
In the following, we will primarily be interested in the case where $X^{\prime}$ and $X$ are both finite cubical complexes, and $f\mathrel{\mathop{\ordinarycolon}}X^{\prime}\to X$ is a surjective homotopy equivalence that is a cubical map, i.e., every cell of $X^{\prime}$ is mapped to a cell of $X$. In this case, given the mapping cylinder $W\mathrel{\mathop{\ordinarycolon}}=M_{f}$, which can be viewed as a simplicial complex, we define the following subsets of $W$:
$\partial_{0}W\mathrel{\mathop{\ordinarycolon}}=(\{0\}\times X^{\prime})/\sim\,,\quad\partial_{1}W\mathrel{\mathop{\ordinarycolon}}=(\{1\}\times X^{\prime})/\sim\,.$
Note that $\partial_{0}W\cong X$ and $\partial_{1}W\cong X^{\prime}$. Heuristically, $W$ is like a cobordism between $\partial_{0}W$ and $\partial_{1}W$. Also, we will sometimes by abuse of notation just denote $\partial_{0}W$ by $X$, and $\partial_{1}W$ by $X^{\prime}$. Now, consider the map
$[0,1]\times[0,1]\times X^{\prime}\to[0,1]\times X^{\prime}$
defined by sending each $(t,s,x^{\prime})$ to $(ts,x^{\prime})$. This induces a map
(3.7) $F_{W}\mathrel{\mathop{\ordinarycolon}}[0,1]\times W\to W$
(see Figure 7). The key property of $F_{W}$ is that the map $(t,w)\mapsto F_{W}(1-t,w)$, $(t,w)\in[0,1]\times W$, is a strong deformation retraction of $W$ onto $\partial_{0}W$: This means for all $t\in[0,1],w\in W,x\in\partial_{0}W$ we have
$F_{W}(1,w)=w\,,\quad F_{W}(0,w)\in\partial_{0}W\,,\quad F_{W}(t,x)=x\,.$
By [BP02, §4], $W$ can also be viewed as a cubical complex.
Now, we can state the interpolation theorem.
Theorem 3.6.
In a closed 333-dimensional Riemannian manifold $(M,g)$, let $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}^{*}(M)$ be a Simon-Smith family of genus 1$\leq$ 1≤ 1, where $X$ is a cubical subcomplex in $I^{m}$. Suppose that $L=\mathbf{L}(\Lambda(\Phi))>0$, and the set $\mathcal{W}_{L,\leq 1}$ consists of exactly one varifold associated with a multiplicity-one nondegenerate minimal sphere. Then after refining $X$, there exists a cubical subcomplex $X^{\prime}$ in $I^{m+1}$, a homotopy equivalence $f\mathrel{\mathop{\ordinarycolon}}X^{\prime}\to X\,,$ which is a surjective cubical map, and a Simon-Smith family $\Phi^{\prime}\mathrel{\mathop{\ordinarycolon}}X^{\prime}\to\mathcal{S}^{*}(M)$ of genus 1$\leq$ 1≤ 1 with the following properties.
  1. (1)
    There exists an $\eta>0$ such that for $x\in X^{\prime}$,
    $\mathcal{H}^{2}(\Phi^{\prime}(x))\geq L-\eta\quad\Longrightarrow\quad\mathfrak{g}(\Phi^{\prime}(x))=0\,.$
  2. (2)
    Let $W$ be the mapping cylinder $M_{f}=(([0,1]\times X^{\prime})\sqcup X)/\sim$ of $f$, where $(0,x)\sim f(x)$ for all $x\in X^{\prime}$. There exists a Simon-Smith family $H\mathrel{\mathop{\ordinarycolon}}W\to\mathcal{S}^{*}(M)$ of genus 1$\leq$ 1≤ 1 such that:
    1. (a)
      $H|_{\partial_{0}W}=\Phi$ and $H|_{\partial_{1}W}=\Phi^{\prime}$.
    2. (b)
      For all $x\in X^{\prime}$, $t\mapsto H(t,x)$ is a pinch-off process (see Definition 3.7).
Heuristically, if a torus is sufficiently to a sphere, one can find a loop within a small ball. Under mean curvature flow, it is expected that the flow would undergo a neck pinch along the loop, yielding a genus 0 punctate surface. However, as explained in §1.2.2, we cannot employ mean curvature flow to reduce the genus. Instead, we pinch off the small ball containing the loop. This process involves a finite number of neck-pinch surgeries, so that the surface becomes disjoint from the small ball, and then shrinking the components inside the ball into a point.
Note that the following definition is purely topological.
Definition 3.7.
[…]
There exist finitely many spacetime points $(t_{1},p_{1}),\cdots,(t_{n},p_{n})$, where $t_{i}\in[a,b]$ and $p_{i}\in\Phi(t_{i})$, such that:
  1. (1)
    For each $(t,p)$, where $t\in[a,b]$ and $p\in\Phi(t)$, different from any $(t_{i},p_{i})$, there exists an open time interval $J\subset[a,b]$ around $t$, a one-parameter group of diffeomorphisms $\{\varphi_{t^{\prime}}\}_{t^{\prime}\in J}\subset\operatorname{Diff}(M)$ and a open ball $U\subset S^{3}$ around $p$ such that for all $t^{\prime}\in J$,
    $\varphi_{t^{\prime}}(\Phi(t))\cap U=\Phi(t^{\prime})\cap U\,.$
  2. (2)
    For each $i=1,\cdots,n$, there exists an open time interval $J\subset[a,b]$ around $t_{i}$ and a ball $U\subset S^{3}$ around $p_{i}$ such that the family
    $\{\Phi(t^{\prime})\cap U\}_{t^{\prime}\in J}$
    is of one of the following types:
    1. (a)
      A surgery process: Pinching a (topological) cylinder at the point $p_{i}$, at time $t_{i}$, to obtain a (topological) double cone, and then either splitting it into two smooth discs or leaving it intact.
    2. (b)
      A shrinking process: For each $t^{\prime}\in J$, $\Phi(t^{\prime})$ does not intersect $\partial U$, and $\{\Phi(t^{\prime})\cap U\}_{t^{\prime}\in J\cap[a,t_{i})}$ is induced by a one-parameter group of diffeomorphisms in $U$, and $\Phi(t^{\prime})\cap U=p_{i}$ for all $t^{\prime}\in J\cap[t_{i},b]$.
Remark 3.8.
Let us compare pinch-off process and mean curvature flow. Mean curvature flow could introduce more complicated singularities than pinch-off process. Namely, under mean curvature flow, besides the above two types of singularities, there could also be singularities modeled by non-compact self-shrinkers of genus greater than 0.
Furthermore, as we will prove in §9, a pinch-off process simplifies the topology of the initial condition in certain sense. This can be compared to B. White’s result that mean curvature flow simplifies the topology of the surfaces [Whi95].
We will prove the interpolation theorem, Theorem 3.6, in §8. Here is a brief outline: First, we show that if a torus is close to a sphere, in the varifold sense, it must contain a non-trivial loop in a small ball. Heuristically, we can then pinch this loop to obtain a punctate surface with genus 0. The main challenge lies in the fact that the choice of loops may not be continuous in a Simon-Smith family, causing the pinching process to fail in producing a desired deformation map $H$. To address this, we will extend Pitts’s combinatorial arguments [Pit81, Theorem 4.10], and construct a map $H$ which consists of a union of pinch-off processes.
[…]
Applying Theorem 3.6 to $\Psi$, we obtain:
  • a cubical complex $\widetilde{Y}^{1}$,
  • a surjective, cubical, homotopy equivalence $f^{1}\mathrel{\mathop{\ordinarycolon}}\widetilde{Y}^{1}\to Y$,
  • a mapping cylinder $W^{1}\mathrel{\mathop{\ordinarycolon}}=M_{f^{1}}$ (which can be viewed as a simplicial complex) with
    $Y=\partial_{0}W^{1},\quad\widetilde{Y}^{1}=\partial_{1}W^{1}\,,$
    and the associated map $F^{1}\mathrel{\mathop{\ordinarycolon}}=F_{W^{1}}\mathrel{\mathop{\ordinarycolon}}[0,1]\times W^{1}\to W^{1}$ (see (3.7)),
  • a Simon-Smith family of genus 1$\leq$ 1≤ 1
    $\widetilde{\Psi}^{1}\mathrel{\mathop{\ordinarycolon}}\widetilde{Y}^{1}\to\mathcal{S}^{*}(S^{3})\,,$
  • a Simon-Smith family of genus 1$\leq$ 1≤ 1
    $H^{1}\mathrel{\mathop{\ordinarycolon}}W^{1}\to\mathcal{S}^{*}(S^{3})\,,$
    with $H^{1}|_{\partial_{0}W^{1}}=\Psi$, $H^{1}|_{\partial_{1}W^{1}}=\widetilde{\Psi}^{1}$,
  • a constant $\delta_{1}>0$ (in place of $\eta$),
satisfying the properties listed in the theorem.
Then by refining $\widetilde{Y}^{1}$, we can obtain a cubical subcomplex $C^{1}$ of $\widetilde{Y}^{1}$ such that
$\mathcal{H}^{2}(\widetilde{\Psi}^{1}(x))\geq\mathbf{L}(\Lambda(\Psi))-\delta_{1}/2\implies x\in C^{1}\,,$
and $\widetilde{\Psi}^{1}|_{C^{1}}$ is a Simon-Smith family of genus 0. Finally, we let
$\Psi^{1}\mathrel{\mathop{\ordinarycolon}}=\widetilde{\Psi}^{1}|_{\overline{\widetilde{Y}^{1}\backslash C^{1}}}\,.$
As in Case 1, a key property of $\Psi^{1}$ is that
$\sup\mathcal{H}^{2}\circ\Psi^{1}<\mathbf{L}(\Lambda(\Psi)).$
Let us denote $Y^{1}\mathrel{\mathop{\ordinarycolon}}=\mathrm{dmn}(\Psi^{1})$.
In summary, in both cases, we obtain a Simon-Smith family $\Psi^{1}$ of genus 1$\leq$ 1≤ 1 with parameter space $\overline{\widetilde{Y}^{1}\backslash C^{1}}$, and with $\sup\mathcal{H}^{2}\circ\Psi^{1}$ strictly less than the width $\mathbf{L}(\Lambda(\Psi))$. Thus, if we apply the Simon-Smith min-max construction again to $\Psi^{1}$, all min-max minimal surfaces will have smaller area than those obtained previously. We will continue repeating the min-max process.
3.2.2. $k$-th stage
In this subsubsection, we explain in detail what is obtained in the $k$-th stage of min-max for $k\geq 1$. Before the $k$-th stage, we have a Simon-Smith family of genus 1$\leq$ 1≤ 1
$\Psi^{k-1}\mathrel{\mathop{\ordinarycolon}}Y^{k-1}\to\mathcal{S}^{*}(S^{3})\,,$
where $Y^{k-1}$ is some cubical complex. If $k=1$, we take $\Psi^{0}\mathrel{\mathop{\ordinarycolon}}=\Psi$ and $Y^{0}\mathrel{\mathop{\ordinarycolon}}=Y$.
Applying the min-max theorem, Theorem 2.16, to $\Psi^{k-1}$ with $r=d_{0}$, we obtain a varifold $V^{k}$ in $\mathcal{W}^{k}\mathrel{\mathop{\ordinarycolon}}=\mathcal{W}_{\mathbf{L}(\Lambda(\Psi^{k-1})),\leq 1}$, where $V^{k}$ is induced by a multiplicity one, smooth, embedded, minimal surface of genus 0 or 111. Again, there are two cases:
  • Case 1.
    each element of $\mathcal{W}^{k}$ is associated with a multiplicity one torus;
  • Case 2.
    $V^{k}$ is associated with a smooth, embedded, minimal sphere of multiplicity one and $\mathcal{W}^{k}=\{V^{k}\}$.
Remark 3.9.
In the following, the superscript k indicates that the object concerned is constructed from running the min-max process for the $k$-th stage.
Case 1.
Arguing as before, we obtain the following objects:
  • a Simon-Smith family of genus 1$\leq$ 1≤ 1
    $\widetilde{\Psi}^{k}\mathrel{\mathop{\ordinarycolon}}Y^{k-1}\to\mathcal{S}^{*}(S^{3})$
    in $\Lambda(\Psi^{k-1})$,
  • a “cap” $C^{k}$, which is a subcomplex of $Y^{k-1}$ and can be decomposed into a disjoint union of cubical subcomplexes $C^{k}_{1},\cdots,C^{k}_{n_{k}}$,
  • the restriction
    $\Psi^{k}\mathrel{\mathop{\ordinarycolon}}=\widetilde{\Psi}^{k}|_{\overline{Y^{k-1}\backslash C^{k}}}\,,$
  • a new domain $Y^{k}\mathrel{\mathop{\ordinarycolon}}=\mathrm{dmn}(\Psi^{k})=\overline{Y^{k-1}\backslash C^{k}}$,
satisfying the following properties:
  • For each $j=1,\cdots,n_{k}$, there is a distinct embedded minimal torus $T^{k}_{j}\in\mathcal{S}^{*}(S^{3})$, with $|T^{k}_{j}|\in\mathcal{W}^{k}$, such that the image
    $[\widetilde{\Psi}^{k}](C^{k-1}_{j})\subset\mathbf{B}^{\mathbf{F}}_{d_{0}}([T^{k}_{j}])\,.$
  • $\sup_{Y^{k}}\mathcal{H}^{2}\circ\Psi^{k}<\mathbf{L}(\Lambda(\Psi^{k-1}))$.
Since $\widetilde{\Psi}^{k}$ and $\Psi^{k-1}$ are homotopic in the sense of Simon-Smith min-max theory, there exists a homotopy
$H^{k}\mathrel{\mathop{\ordinarycolon}}[0,1]\times Y^{k-1}\to\mathcal{S}^{*}(S^{3})\,,$
such that
$H(0,\cdot)=\Psi^{k-1},\quad H^{k}(1,\cdot)=\widetilde{\Psi}^{k}\,,$
and for each $(t,x)$, $H(t,x)$ is obtained from $\Psi^{k-1}(x)$ via an isotopy of $S^{3}$, according to Definition 2.10.
For consistency in both cases, we introduce the following notation, which mirrors the setup in Case 2. Define
$W^{k}\mathrel{\mathop{\ordinarycolon}}=[0,1]\times Y^{k-1},\quad\partial_{0}W^{k}=\{0\}\times Y^{k-1},\quad\partial_{1}W^{k}=\{1\}\times Y^{k-1}\,,$
and a map
$F^{k}\mathrel{\mathop{\ordinarycolon}}[0,1]\times W^{k}\to W^{k}\,,\quad(s,(t,x))\mapsto(st,x)\,.$
Note that the map $(s,w)\mapsto F^{k}(1-s,w)$, with $(s,w)\in[0,1]\times W^{k}$, is a strong deformation retraction of $W^{k}$ onto $\{0\}\times Y^{k-1}$. For convenience, we will identify $Y^{k-1}$ with the subset $\{0\}\times Y^{k-1}$ in $W^{k}$, and denote $\{1\}\times Y^{k-1}$ by $\widetilde{Y}^{k}$.
Remark 3.10.
For example, in Case 1, if we have a subset $C\subset Y^{k-1}$, then $(F^{k}(0,\cdot))^{-1}(C)$ is the subset $[0,1]\times C$ of $W^{k}$.
Case 2.
In this case, by applying Theorem 3.6 to $\Psi^{k-1}$ with $r=d_{0}$, we obtain:
  • a cubical complex $\widetilde{Y}^{k}$,
  • a surjective, cubical, homotopy equivalence $f^{k}\mathrel{\mathop{\ordinarycolon}}\widetilde{Y}^{k}\to Y^{k-1}$,
  • a mapping cylinder $W^{k}\mathrel{\mathop{\ordinarycolon}}=M_{f^{k}}$ (viewed as a simplicial complex) with
    $Y^{k-1}=\partial_{0}W^{k}\,,\quad\widetilde{Y}^{k}=\partial_{1}W^{k}\,,$
    and the associated map $F^{k}\mathrel{\mathop{\ordinarycolon}}=F_{W^{k}}\mathrel{\mathop{\ordinarycolon}}[0,1]\times W^{k}\to W^{k}$ (see (3.7)),
  • a Simon-Smith family of genus 1$\leq$ 1≤ 1
    $\widetilde{\Psi}^{k}\mathrel{\mathop{\ordinarycolon}}\widetilde{Y}^{k}\to\mathcal{S}^{*}(S^{3})\,,$
  • a Simon-Smith family of genus 1$\leq$ 1≤ 1
    $H^{k}\mathrel{\mathop{\ordinarycolon}}W^{k}\to\mathcal{S}^{*}(S^{3}),$
    with $H^{k}|_{\partial_{0}W^{k}}=\Psi^{k-1}$, $H^{k}|_{\partial_{1}W^{k}}=\widetilde{\Psi}^{k}$,
  • a “cap” $C^{k}$, which is a cubical subcomplex of $\widetilde{Y}^{k}$,
  • the restriction
    $\Psi^{k}\mathrel{\mathop{\ordinarycolon}}=\widetilde{\Psi}^{k}|_{\overline{Y^{k-1}\backslash C^{k}}}\,,$
  • a new domain $Y^{k}\mathrel{\mathop{\ordinarycolon}}=\mathrm{dmn}(\Psi^{k})=\overline{Y^{k-1}\backslash C^{k}}$,
satisfying the following:
  • The map $(t,w)\mapsto F^{k}(1-t,w)$, where $(t,w)\in[0,1]\times W^{k}$, is a strong deformation retraction of $W^{k}$ onto $Y^{k-1}$.
  • For each $x\in\widetilde{Y}^{k}$, the family $t\mapsto H^{k}(F^{k}(t,x))$, $t\in[0,1]$, is a pinch-off process.
  • $\widetilde{\Psi}^{k}|_{C^{k}}$ is a Simon-Smith family of genus 0.
  • $\sup\mathcal{H}^{2}\circ\Psi^{k}<\mathbf{L}(\Lambda(\Psi^{k-1}))$.
As before, in both cases, the mass bound
(3.8) $\sup\mathcal{H}^{2}\circ\Psi^{k}<\mathbf{L}(\Lambda(\Psi^{k-1}))$
holds. Thus, the width $\mathbf{L}(\Lambda(\Psi^{k}))$ is strictly decreasing with respect to $k$. Combined with the assumption that $(S^{3},g^{\prime})$ admits only finitely many embedded minimal spheres and tori, our repetitive min-max process must terminate in finitely many steps. Specifically, for some $K\in\mathbb{N}^{+}$, when we apply the min-max process for the $K$-th time to the family $\Psi^{K-1}$, the width $\mathbf{L}(\Lambda(\Psi^{K-1}))$ reaches zero.
3.2.3. Last stage
We have that the width $\mathbf{L}(\Lambda(\Psi^{K-1}))$ is zero. Thus, there exists some Simon-Smith family $\Psi^{K}\in\Lambda(\Psi^{K-1})$ and a homotopy in the sense of Simon-Smith min-max,
$H^{K}\mathrel{\mathop{\ordinarycolon}}[0,1]\times Y^{K-1}\to\mathcal{S}^{*}(S^{3}),$
such that:
  • $H^{K}(0,\cdot)=\Psi^{K-1}$, $H^{K}(1,\cdot)=\Psi^{K}$.
  • For each $(t,x)$, $H^{K}(t,x)$ is obtained from $\Psi^{K-1}(x)$ via some diffeomorphism of $S^{3}$ according to Definition 2.10.
  • $\sup_{Y^{K-1}}\mathcal{H}^{2}_{\bar{g}}\circ\Psi^{K}<1$, where $\bar{g}$ denotes the standard round metric on $S^{3}$.
Crucially, note that in the last bullet point, the Hausdorff measure is taken with respect to the unit 3-sphere $(S^{3},\bar{g})$.
We denote
$W^{K}\mathrel{\mathop{\ordinarycolon}}=[0,1]\times Y^{K-1}\,,\quad\partial_{0}W^{K}=\{0\}\times Y^{K-1}\,,\quad\partial_{1}W^{K}=\{1\}\times Y^{K-1}\,,$
and define a map
$F^{K}\mathrel{\mathop{\ordinarycolon}}[0,1]\times W^{K}\to W^{K}\,,\quad(s,(t,x))\mapsto(st,x)\,.$
Note that the map $(s,w)\mapsto F^{K}(1-s,w)$, with $(s,w)\in[0,1]\times W^{K}$, is a strong deformation retraction of $W^{K}$ onto $\{0\}\times Y^{K-1}$. As before, we identify $Y^{K-1}$ with the subset $\{0\}\times Y^{K-1}$ in $W^{K}$.
3.3. A new family $\Xi$
We have already run the min-max process $K$ times, generating multiple families from the original family $\Psi$, each with a different parameter space. In this subsection, we will use these families to reconstruct a new Simon-Smith family $\Xi$, whose parameter space is homotopy equivalent to the original parameter space $Y$ (which is an ${\mathbb{RP}}^{5}$-bundle over ${\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$). This new family remains a Simon-Smith family of genus 1$\leq$ 1≤ 1.
Consider the following list of $2K-1$ Simon-Smith families of genus 1$\leq$ 1≤ 1. For visualization purposes, readers may find it helpful to refer to Figure  8. When interpreting the notations below, we should keep in mind that we always identify $\partial_{0}W^{k}$ with $Y^{k-1}=\mathrm{dmn}(\Psi^{k-1})$ (and also thus their respective subsets), and we identify $\partial_{1}W^{k}$ with $\mathrm{dmn}(\widetilde{\Psi}^{k})$.
$\widetilde{\Psi}^{1}|_{C^{1}}$.
$H^{2}|_{(F^{2}(1,\cdot))^{-1}(A_{1})}$, where $A_{1}\mathrel{\mathop{\ordinarycolon}}=C^{1}\cap Y^{1}\;(\subset\partial_{0}W^{2}\subset W^{2})$. Note that $A_{1}$ is simply $\partial C^{1}$ if $\widetilde{Y}^{1}$ is a topological manifold and $C^{1}$ is a $\dim(\widetilde{Y}^{1})$-chain in it. Also, $H^{2}|_{(F^{2}(1,\cdot))^{-1}(A_{1})}$ is the same as $H^{2}|_{[0,1]\times A_{1}}$ if Case 1 occurred at the second stage of the min-max process; see Remark 3.10.
$\widetilde{\Psi}^{2}|_{C^{2}}$.
$H^{3}|_{(F^{3}(1,\cdot))^{-1}(A_{2})}$, where $A_{2}\subset\partial_{0}W^{3}$ is defined as the union of the following two subsets of $\partial_{0}W^{3}\;(\subset W^{3})$:
  • $(f^{2})^{-1}(A_{1})\cap Y^{2}\;(\subset Y^{2}\cong\partial_{0}W^{3})$.
  • $C^{2}\cap Y^{2}\;(\subset\partial_{0}W_{3})$.
In other words,
$A_{2}\mathrel{\mathop{\ordinarycolon}}=\left((f^{2})^{-1}(A_{1})\cap Y^{2}\right)\cup\left(C^{2}\cap Y^{2}\right)\,.$
(In Figure  8, we see a blue bridge, and a boundary part of it (on the right). This boundary part consists of two curves: The “top curve” corresponds to the set described in the second bullet point, while the “side curve” (which coincides with part of the left boundary of the red bridge) corresponds to the set in the first bullet point.) Note that $H^{3}|_{(F^{3}(1,\cdot))^{-1}(A_{2})}$ coincides with $H^{3}|_{[0,1]\times A_{2}}$ if Case 1 occurred at the third stage of min-max.
$\widetilde{\Psi}^{3}|_{C^{3}}$.
$H^{4}|_{(F^{4}(1,\cdot))^{-1}(A_{3})}$, where $A_{3}\subset\partial_{0}W^{4}$ is defined by
$A_{3}\mathrel{\mathop{\ordinarycolon}}=\left((f^{3})^{-1}(A_{2})\cap Y^{3}\right)\cup\left(C^{3}\cap Y^{3}\right)\,.$
($2K-4$)
($2K-3$)
$\widetilde{\Psi}^{K-1}|_{C^{K-1}}$.
($2K-2$)
$H^{K}|_{(F^{K}(1,\cdot))^{-1}(A_{K-1})}$, where $A_{K-1}\subset\partial_{0}W^{K}$ is defined by
$A_{K-1}\mathrel{\mathop{\ordinarycolon}}=\left((f^{K-1})^{-1}(A_{K-2})\cap Y^{K-1}\right)\cup\left(C^{K-1}\cap Y^{K-1}\right)\,.$
($2K-1$)
$\Psi^{K}$.
Figure 8. Constructing $\Xi$.
By the definition of these $2K-1$ families, there is a natural way to glue them together. Namely, for each $k=1,\cdots,K-2$, the following pairs of families share a common subfamily:
[…]
Let the newly obtained family be called $\Xi$. It is also a Simon-Smith family of genus 1$\leq$ 1≤ 1.
Proposition 3.11.
There exist:
  • a simplicial complex $\widetilde{W}$, containing $Y$ and $\mathrm{dmn}(\Xi)$ as subcomplexes,
  • a map $\widetilde{F}\mathrel{\mathop{\ordinarycolon}}[0,1]\times\widetilde{W}\to\widetilde{W}$ such that the map $(t,w)\mapsto\widetilde{F}(1-t,w)$, with $(t,w)\in[0,1]\times\widetilde{W}$, is a strong deformation retraction of $\widetilde{W}$ onto $Y$, and $\widetilde{F}(0,\cdot)|_{\mathrm{dmn}(\Xi)}$ is a surjective homotopy equivalence from $\mathrm{dmn}(\Xi)$ onto $Y$,
  • a Simon-Smith family of genus 1$\leq$ 1≤ 1,
    $\widetilde{\Xi}\mathrel{\mathop{\ordinarycolon}}\widetilde{W}\to\mathcal{S}^{*}(S^{3}),$
such that:
  1. (1)
    $\widetilde{\Xi}|_{Y}=\Psi,\;\;\widetilde{\Xi}|_{\mathrm{dmn}(\Xi)}=\Xi.$
  2. (2)
    For each $x\in\mathrm{dmn}(\Xi)$, the family $t\mapsto\widetilde{\Xi}(\widetilde{F}(t,x))$, with $t\in[0,1]$, is a pinch-off process.
Thus, heuristically, the map $(t,w)\mapsto\widetilde{\Xi}(\widetilde{F}(1-t,w))$, where $(t,w)\in[0,1]\times\mathrm{dmn}(\Xi)$, can be viewed as a generalized homotopy that deforms $\Xi$ back to $\Psi$.
Proof.
By examining the construction of $\Xi$, there is a natural, canonical way to define the objects mentioned in the proposition. Specifically, we introduce a sequence of “$\mathbf{F}$-homotopies” that, in a certain sense, homotope $\Xi$ back to $\Psi$. Before proceeding, we introduce the following notations:
  • The $\mathbf{F}$-metric in $\mathcal{Z}_{2}(S^{3};\mathbb{Z}_{2})$ does not induce a metric on the space $\mathcal{S}^{*}(S^{3})$. However, the $\mathbf{F}$-metric would induce a pseudometric, and therefore a topology, on $\mathcal{S}^{*}(S^{3})$. With this in mind, the list in the next paragraph would describe a sequence of homotopies, called $\mathbf{F}$-homotopies under this topology.
  • We use “$+$” to define the addition of chains with $\mathbb{Z}_{2}$-coefficients in $\mathcal{S}^{*}(S^{3})$, where overlapping portions cancel out each other (see, for example, Figure 8).
  • We use the numbering $(1),(2),\cdots,(2K-1)$ for the subfamilies of $\Xi$ as stated at the beginning of §3.3.
Let us now consider the following sequence of $\mathbf{F}$-homotopies. For visualization purposes, readers may refer to Figure  8.
  • The family $\Xi$ can be written as the sum of three subfamilies:
    $\Xi=(2K-1)+(2K-2)+\textrm{ remaining part}.$
    We homotope the subfamily $(2K-1)+(2K-2)$ back to $\Psi^{K-1}$ (using the family $H^{K}$ and the strong deformation retraction of $W^{K}$ onto $Y^{K-1}$ induced by $F^{K}$), while fixing the remaining part. Call this new family $\Xi^{K-1}$.
  • The family $\Xi^{K-1}$ can be written as the sum of three subfamilies:
    $\Xi^{K-1}=\widetilde{\Psi}^{K-1}+(2K-4)+\textrm{ remaining part.}$.
    We homotope the subfamily $\widetilde{\Psi}^{K-1}+(2K-4)$ to $\Psi^{K-2}$ (using the family $H^{K-1}$ and the strong deformation retraction of $W^{K-1}$ onto $Y^{K-2}$ induced by $F^{K-1}$), while fixing the remaining part. Call this new family $\Xi^{K-2}$.
  • The family $\Xi^{2}$ can be written as the sum of three subfamilies:
    $\Xi^{2}=\widetilde{\Psi}^{2}+(2)+\textrm{ remaining part}.$
    We homotope the subfamily $\widetilde{\Psi}^{2}+(2)$ to $\Psi^{1}$ (using the family $H^{2}$ and the strong deformation retraction of $W^{2}$ onto $Y^{1}$ induced by $F^{2}$), while fixing the remaining part. Call this new family $\Xi^{1}$, which is actually $\widetilde{\Psi}^{1}$.
  • Finally, we homotope $\Xi^{1}=\widetilde{\Psi}^{1}$ to $\Psi$ using the family $H^{1}$ and the strong deformation retraction of $W^{1}$ onto $Y^{0}=Y$ induced by $F^{1}$.
From the above chain of $\mathbf{F}$-homotopies, one can easily construct $\widetilde{W}$, $\widetilde{F}$, and $\widetilde{\Xi}$ satisfying the desired conditions. ∎
3.4. Decomposing $\Xi$
Recall that for each $k=1,\cdots,K-1$, during the $k$-th min-max process, one of two cases occurred: (1) Some multiplicity one tori $T^{k}_{1},\cdots,T^{k}_{n_{k}}$ were detected, and (2) a multiplicity one sphere $V^{k}$ was detected. For each $k$, if Case 1 occurred, then $C^{k}\subset\mathrm{dmn}(\widetilde{\Psi}^{k})$ can be decomposed into a union of $C^{k}_{1},\cdots,C^{k}_{n_{k}}$, and we will call each $C^{k}_{j}$ a genus 111 cap. Recall that $[\widetilde{\Psi}^{k}|_{C^{k}_{j}}]$ is $\mathbf{F}$-close to $[T^{k}_{j}]$. If instead Case 2 occurred, we call $C^{k}$ a genus 0 cap. Recall that $\widetilde{\Psi}^{k}|_{C^{k}}$ is a Simon-Smith family of genus 0.
If $C\subset\mathrm{dmn}(\widetilde{\Psi}^{k})$ is any such cap obtained at the $k$-th stage of min-max, we define the trace of $C$, which is a subset of $\mathrm{dmn}(\Xi)$, as follows. We first define the following sets (see Figure 9):
  • $B_{k}$:$\;\mathrel{\mathop{\ordinarycolon}}=C$, a subcomplex of $\mathrm{dmn}(\widetilde{\Psi}^{k})=\widetilde{Y}^{k}$.
  • ${\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}B_{k+1}}\mathrel{\mathop{\ordinarycolon}}=(F^{k+1}(1,\cdot))^{-1}(B_{k}\cap Y^{k})\subset W^{k+1}$. Here $B_{k}\cap Y^{k}$ can also be viewed as a subset of $\mathrm{dmn}(\Psi^{k})=\partial_{0}W^{k+1}$.
  • ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}B_{k+2}}\;\mathrel{\mathop{\ordinarycolon}}=(F^{k+2}(1,\cdot))^{-1}(B_{k+1}\cap Y^{k+1})\subset W^{k+2}$.
Note that each of $B_{k},B_{k+1},\cdots,B_{K}$ is regarded as a subset of $\mathrm{dmn}(\Xi)$. Then the trace of $C$ is defined as their union
$T(C)\mathrel{\mathop{\ordinarycolon}}=B_{k}\cup B_{k+1}\cup\cdots\cup B_{K}\subset\mathrm{dmn}(\Xi)\,.$
Figure 9. The trace $T(C)$ in $\mathrm{dmn}(\Xi)$.
Now, let $N$ be the number of all the genus 1 caps we obtained throughout the first $K-1$ stages of min-max. To prove Theorem A, we just need to show that $N\geq 5$.
We decompose $\mathrm{dmn}(\Xi)$ as the union of the following $N+1$ subcomplexes:
  • Let $D_{0}$ be the union of $\mathrm{dmn}(\Psi^{K})$ and $\bigcup_{C}T(C)$, where $C$ ranges over the genus 0 caps.
  • For each genus 111 cap $C$, we consider its trace $T(C)$. There are in total $N$ such traces, which are denoted by $D_{1},\cdots,D_{N}$.
It follows directly from the definition of $\Xi$ that
$\mathrm{dmn}(\Xi)=D_{0}\cup D_{1}\cup\cdots\cup D_{N}\,.$
Proposition 3.12.
The decompositions above satisfy the following two properties:
  1. (1)
    $\Xi|_{\bigcup_{C}T(C)}$, where $C$ ranges over the genus 0 caps, is a Simon-Smith family of genus 0.
  2. (2)
    For each genus 0 or genus 111 cap $C$, there exists a strong deformation retraction of the trace $T(C)$ back to $C$ in $\mathrm{dmn}(\Xi)$.
Proof.
For (1), if $C$ is a genus 0 cap obtained at the $k$-stage of min-max, then $\widetilde{\Psi}^{k}|_{C^{k}}$ is a Simon-Smith family of genus 0. Then it follows from the definition of the homotopies $H^{k}$, namely Theorem 3.6 (2) (2)(b), that $\Xi|_{T(C)}$ is also a Simon-Smith family of genus 0. This proves (1). Item (2) follows directly from the definition of $T(C)$, using the maps $F^{k+1},\cdots,F^{K}$. ∎
3.5. Topological arguments
In this subsection, we will show that $N\geq 5$.
First, we define three elements $\lambda,\alpha,\beta$ in $H^{1}(Y;\mathbb{Z}_{2})$:
  • Let $\bar{\lambda}$ be the non-trivial element of $H^{1}(\mathcal{Z}_{2}(S^{3};\mathbb{Z}_{2});\mathbb{Z}_{2})$. Define $\lambda$ as the pullback of $\bar{\lambda}$ under $[\Psi]$.
  • Let $a={\mathbb{RP}}^{1}\times{\mathbb{RP}}^{2}\subset{\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$, and $A\subset Y$ be the ${\mathbb{RP}}^{5}$-subbundle over $a$. Define $\alpha$ to be the Poincaré dual $PD(A)$ of $A$.
  • Let $b={\mathbb{RP}}^{2}\times{\mathbb{RP}}^{1}\subset{\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$, and $B\subset Y$ be the ${\mathbb{RP}}^{5}$-subbundle over $b$. Define $\beta$ to be the Poincaré dual $PD(B)$ of $B$.
Here we assumed ${\mathbb{RP}}^{2}$ is built from attaching a 222-disk to the non-trivial loop ${\mathbb{RP}}^{1}\subset{\mathbb{RP}}^{2}$.
The following topological fact about $Y$ is crucial and is the main reason we used the family $\Psi$ in the first place.
Theorem 3.13.
In the cohomology ring $H^{*}(Y;\mathbb{Z}_{2})$,
$\lambda^{5}\cup\alpha^{2}\cup\beta^{2}\neq 0\,.$
We will prove this theorem in §5.
Now, to prove $N\geq 5$, let us suppose by contradiction that $N\leq 4$.
By Proposition 3.11, $\mathrm{dmn}(\Xi)$ and $Y$ are homotopy equivalent, so, by abuse of notation, we will also treat view $\lambda,\alpha,\beta$ as elements of $H^{1}(\mathrm{dmn}(\Xi);\mathbb{Z}_{2})$ as well. For each $j=0,1,\cdots,N$, let $i_{j}$ be the inclusion map $D_{j}\hookrightarrow\mathrm{dmn}(\Xi)$. Moreover, given a cohomology class $\gamma$ of $\mathrm{dmn}(\Xi)$, we denote its pullback under
$(i_{j})^{*}\mathrel{\mathop{\ordinarycolon}}H^{*}(\mathrm{dmn}(\Xi);\mathbb{Z}_{2})\to H^{*}(D_{j};\mathbb{Z}_{2})$
by $\gamma|_{D_{j}}$. Then by Theorem 3.13 and a Lyusternik-Schnirelmann argument, we obtain:
Lemma 3.14.
  • If $N=4$, then one of $\lambda^{5}|_{D_{0}},\alpha|_{D_{1}},\alpha|_{D_{2}},\beta|_{D_{3}},\beta|_{D_{4}}$ is non-zero.
  • If $N=3$, then one of $\lambda^{5}|_{D_{0}},\alpha|_{D_{1}},\alpha|_{D_{2}},\beta^{2}|_{D_{3}}$ is non-zero.
  • If $N=2$, then one of $\lambda^{5}|_{D_{0}},\alpha|_{D_{1}},(\alpha\cup\beta^{2})|_{D_{2}}$ is non-zero.
  • If $N=1$, then one of $\lambda^{5}|_{D_{0}},(\alpha^{2}\cup\beta^{2})|_{D_{1}}$ is non-zero.
  • If $N=0$, i.e. $\mathrm{dmn}(\Xi)=D_{0}$, then $\lambda^{5}|_{D_{0}}$ is non-zero.
By Proposition 3.11, regarding the induced maps $[\Psi]$ and $[\Xi]$ into $\mathcal{Z}_{2}(S^{3};\mathbb{Z}_{2})$, we have
$[\Xi]^{*}(\bar{\lambda})=[\Psi]^{*}(\bar{\lambda})=\lambda\,.$
Hence,
$\lambda^{5}|_{D_{0}}=i_{0}^{*}([\Xi]^{*}(\bar{\lambda})^{5})=([\Xi\circ i_{0}]^{*}(\bar{\lambda}))^{5}=([\Xi|_{D_{0}}]^{*}(\bar{\lambda}))^{5}\,,$
which is non-zero if and only if $[\Xi|_{D_{0}}]$ is a 555-sweepout. Consequently, it suffices to prove the following two facts:
Proposition 3.15.
$\Xi|_{D_{0}}$ is not a 555-sweepout.
Theorem 3.16.
Suppose $1\leq N\leq 4$. Then for each $j=1,\cdots,N$, the map
$(i_{j})_{*}\mathrel{\mathop{\ordinarycolon}}H_{1}(D_{j};\mathbb{Z}_{2})\to H_{1}(\mathrm{dmn}(\Xi);\mathbb{Z}_{2})$
is the zero map.
Indeed, Proposition 3.15 implies $\lambda^{5}|_{D_{0}}=0$, and Theorem 3.16, via the universal coefficients theorem, implies the map
$(i_{j})^{*}\mathrel{\mathop{\ordinarycolon}}H^{1}(\mathrm{dmn}(\Xi);\mathbb{Z}_{2})\to H^{1}(D_{j};\mathbb{Z}_{2})$
is trivial for each $j=1,\cdots,N$. This implies that all the cohomology classes in Lemma 3.14 must be zero, contradicting the assumption that $N\leq 4$. Thus, it remains to prove Proposition 3.15 and Theorem 3.16.
Proof of Proposition 3.15.
In this proof, we will treat $\Xi|_{D_{0}}$ as a Simon-Smith family of genus 1$\leq$ 1≤ 1 in the unit 3-sphere $(S^{3},\bar{g})$ instead of $(S^{3},g^{\prime})$. Recall that $D_{0}$ is the union of $E\mathrel{\mathop{\ordinarycolon}}=\bigcup_{C}T(C)$, where $C$ runs over the genus 0 caps, and $\mathrm{dmn}(\Psi^{K})$.
By Proposition 3.12, $\Xi|_{E}$ is a Simon-Smith family of genus 0, while
$\sup\mathcal{H}^{2}_{\bar{g}}\circ\Xi|_{\mathrm{dmn}(\Psi^{K})}=\sup\mathcal{H}^{2}_{\bar{g}}\circ\Psi^{K}<1\;\;(<4\pi)$
by definition. Thus, if we apply the relative Simon-Smith min-max, Theorem 2.21, on the unit round 3-sphere to $\Xi|_{D_{0}}$, relative to $\mathrm{dmn}(\Psi^{K})\subset D_{0}$, we will detect the multiplicity one equatorial 2-sphere. This immediately implies that the width $\mathbf{L}(\Lambda(\Xi|_{D_{0}}))$, in the non-relative Simon-Smith min-max setting, is at most $4\pi$.
Note the following fact:
Claim 3.17.
In a Riemannian 333-manifold $M$, if $\Phi$ is a Simon-Smith family such that $[\Phi]$ is a $p$-sweepout, then the Simon-Smith min-max width $\mathbf{L}(\Lambda(\Phi))$ satisfies
$\mathbf{L}(\Lambda(\Phi))\geq\omega_{p}(M)\,.$
Proof.
This is true because it follows from definition that a homotopy in the sense of Simon-Smith min-max is continuous in the flat topology. ∎
Thus, whenever $[\Xi|_{D_{0}}]$ is a $p$-sweepout,
$4\pi=\mathbf{L}(\Lambda(\Xi|_{D_{0}}))\geq\omega_{p}(S^{3},\bar{g})\,.$
But by C. Nurser [Nur16], the first four widths of the unit 333-sphere are $4\pi$, while the fifth is $2\pi^{2}$. Thus, $[\Xi|_{D_{0}}]$ is not a 555-sweepout. This finishes the proof of Proposition 3.15
We postpone the detailed proof of Theorem 3.16 to §10. For now, let us outlined the idea here. Let $1\leq j\leq N$. By definition, $D_{j}$ is the trace $T(C)$ of some genus 1 cap $C\subset\mathrm{dmn}(\Xi)$ obtained at the $k$-th stage of min-max, for some $1\leq k\leq K-1$. By Proposition 3.12, there exists a strong deformation retraction of $D_{j}$ onto $C$, so it suffices to show that for the inclusion map $i\mathrel{\mathop{\ordinarycolon}}C\hookrightarrow\mathrm{dmn}(\Xi)$, the induced map $i_{*}$ between the first homology groups of $C$ and $\mathrm{dmn}(\Xi)$ is zero. By definition, the entire image of $[\Xi|_{C}]$ is $d_{0}$-close to a single minimal torus in the $\mathbf{F}$-metric for currents. We will show in §10 that this means $i_{*}$ indeed has to be the zero map, thereby proving Theorem 3.16.
Finally, as previously explained, Proposition 3.15 and Theorem 3.16 together contradict Lemma 3.14, thereby proving that the assumption $N\leq 4$ is false. Consequently, at least 555 multiplicity one, embedded, minimal tori are detected through the repetitive min-max process. This completes the proof of Theorem A.
Part II Technical ingredients
4. Min-max results I: The Simon-Smith min-max theorem
The goal of this section is to prove Theorem 2.16 and subsequently Theorem 2.21.
4.1. Deformations in annuli
One of the innovative ingredients in the Almgren-Pitts min-max theory is the notion of almost-minimizing varifolds in annuli, which was established using combinatorial arguments for deformations in annuli. In this subsection, we adapt these tools to our setting.
For an open subset $U\subset M$, we use $\mathfrak{Is}(U)$ to denote the set of isotopies $\{\varphi(t)\}_{t\in[0,1]}$ in $U$.
Definition 4.1 ($(\varepsilon,\delta)$-deformation).
Given $\varepsilon,\delta>0$, an open set $U\subset M$, and a punctate surface $\Sigma\in\mathcal{S}(M)$, we call an isotopy $\psi\in\mathfrak{Is}(U)$ an $(\varepsilon,\delta)$-deformation of $\Sigma$ in $U$ provided that:
  1. (1)
    $\mathcal{H}^{2}(\psi(t,\Sigma))\leq\mathcal{H}^{2}(\Sigma)+\delta$ for all $t\in[0,1]$.
  2. (2)
    $\mathcal{H}^{2}(\psi(1,\Sigma))\leq\mathcal{H}^{2}(\Sigma)-\varepsilon$.
We define $\mathfrak{a}(U;\varepsilon,\delta)$ to be the set of all punctate surfaces that do not admit $(\varepsilon,\delta)$-deformations in $U$.
Definition 4.2 (Admissible annuli).
Given $K\in\mathbb{N}$ and $p\in M$, a collection of annuli centered in $p$
$A(p;s_{1},r_{1}),\cdots,A(p;s_{K},r_{K}),$
is called $K$-admissible if $2r_{i+1}<s_{i}$ for all $i=1,\cdots,K-1$.
For $R\in(0,\infty]$, if $\sup_{i}r_{i}\leq R$, we will say that these annuli are of outer radius at most $R$.
The following proposition is an adaptation of Pitts’ combinatorial result [Pit81, Theorem 4.10].
Proposition 4.3.
Given $X$ a finite cubical subcomplex of some cube $I(m,k)$, $D\subset X$ a compact subset (not necessarily a subcomplex) of $X$, $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}^{*}(M)$ a Simon-Smith family, $R\in(0,\infty]$ and a pair positive numbers $(\varepsilon,\delta)$, we set $K=K(m)\mathrel{\mathop{\ordinarycolon}}=3^{m3^{m}}$. Suppose that for every $x\in D$, there exists a $K$-admissible collection of annuli of outer radius at most $R$, denoted by $\{A_{x,i}\}^{K}_{i=1}$, so that
$\Phi(x)\notin\bigcup^{K}_{i=1}\mathfrak{a}(A_{x,i};2\varepsilon,\delta/2)\,.$
Then there exists $\Phi^{*}\in\Lambda(\Phi)$ in the homotopy class of $\Phi$ such that:
  1. (1)
    For every $x\in D$,
    $\mathcal{H}^{2}(\Phi^{*}(x))<\mathcal{H}^{2}(\Phi(x))-\varepsilon+(3^{m}-1)\delta\,.$
  2. (2)
    For every $x\in X\setminus D$,
    $\mathcal{H}^{2}(\Phi^{*}(x))<\mathcal{H}^{2}(\Phi(x))+3^{m}\delta\,.$
  3. (3)
    For every $x\in X$, there exist $t$ many points $p_{1},\cdots,p_{t}\in M$ such that $t\leq K(m)$ and
    $\Phi^{*}(x)\setminus(\overline{B}_{2R}(p_{1})\cup\cdots\cup\overline{B}_{2R}(p_{t}))=\Phi(x)\setminus(\overline{B}_{2R}(p_{1})\cup\cdots\cup\overline{B}_{2R}(p_{t}))\,.$
  4. (4)
    Moreover, for another compact set $D^{\prime}\subset X$ with $D^{\prime}\cap D=\emptyset$, we can choose $\Phi^{*}$ such that $\Phi^{*}(x)=\Phi(x)$ holds for all $x\in D^{\prime}$.
Proof.
First, by our assumption, for each $x\in D$ and $i\in\{1,2,\cdots,K\}$, there exists an isotopy $\psi_{x,i}\in\mathfrak{Is}(A_{x,i})$ such that:
  • $\mathcal{H}^{2}(\psi_{x,i}(t,\Phi(x)))\leq\mathcal{H}^{2}(\Phi(x))+\delta/2$ for all $t\in[0,1]$.
  • $\mathcal{H}^{2}(\psi_{x,i}(1,\Phi(x)))\leq\mathcal{H}^{2}(\Phi(x))-2\varepsilon$.
By Proposition 2.8, $[\Phi]$ is continuous in the $\mathbf{F}$-metric, and thus, for each $x\in D$, there exists an open neighborhood $O_{x}\subset X$ of $x$ such that for every $y\in O_{x}$ and every $i\in\{1,2,\cdots,K\}$, we have:
  • $\mathcal{H}^{2}(\psi_{x,i}(t,\Phi(y)))\leq\mathcal{H}^{2}(\Phi(y))+\delta$ for all $t\in[0,1]$.
  • $\mathcal{H}^{2}(\psi_{x,i}(1,\Phi(y)))\leq\mathcal{H}^{2}(\Phi(y))-\varepsilon$.
Next, since the set $D$ is compact, we can refine $X$ so that it is a cubical subcomplex of $I(m,k^{\prime})$ and every cell $\sigma$ of $X$ with $\sigma\cap D\neq\emptyset$ is contained within some $O_{x_{\sigma}}$ where $x_{\sigma}\in D$. We denote by $X_{b}$ the smallest cubical subcomplex of $X$ containing
$\{\sigma\in X\mathrel{\mathop{\ordinarycolon}}\sigma\cap D\neq\emptyset\}\,.$
Moreover, to ensure that conclusion (4) holds, we also require that $X^{\prime}_{b}$ the smallest cubical subcomplex of $X$ containing
$\{\sigma\in X\mathrel{\mathop{\ordinarycolon}}\sigma\cap D^{\prime}\neq\emptyset\}\,,$
satisfies the condition that $X_{b}\cap X^{\prime}_{b}=\emptyset$.
It follows from Proposition 4.9 of [Pit81] that for every cell $\sigma$ of $X_{b}$ we can choose an annulus $A_{\sigma}\in\{A_{x_{\sigma},1},\cdots,A_{x_{\sigma},K}\}$ and the corresponding isotopy $\psi_{\sigma}\in\mathfrak{Is}(A_{\sigma})$ with the following property. For every pair of distinct cells $\sigma,\tau$ of $X_{b}$, if they are faces of a common cell $\gamma$ of $X_{b}$, then
$\operatorname{dist}(A_{\sigma},A_{\tau})>0\,.$
Additionally, for every cell $\sigma$ of $X_{b}$, we can choose a cut-off function
$\eta_{\sigma}\mathrel{\mathop{\ordinarycolon}}X\to[0,1]\,,$
such that:
  • $\eta_{\sigma}(x)\equiv 0$ or every $x\in X$ with $\|x-c_{\sigma}\|_{\ell_{\infty}}\geq 3^{-k^{\prime}}/2$,
  • $\eta_{\sigma}(x)\equiv 1$ for every $x\in X$ with $\|x-c_{\sigma}\|_{\ell_{\infty}}\leq 3^{-k^{\prime}}/3$,
where $c_{\sigma}$ is the center of $\sigma$. Consequently, we obtain a continuous map
$\hat{\psi}_{\sigma}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\operatorname{Diff}^{\infty}(M),\quad(t,x)\mapsto\psi_{\sigma}(t\eta_{\sigma}(x))\,.$
Note that $\hat{\psi}_{\sigma}(t,x)=\operatorname{Id}$ outside $A_{\sigma}$.
Finally, we define a continuous map
$\varphi\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\operatorname{Diff}^{\infty}(M)$
by composing all the $\hat{\psi}_{\sigma}$
$\varphi(t,x)\mathrel{\mathop{\ordinarycolon}}=\circ_{\sigma\text{ is a cell of }X_{b}}\left(\hat{\psi}_{\sigma}(t,x)\right)\,.$
Note that no two annuli $A_{\sigma}$ and $A_{\tau}$ intersect unless either $\hat{\psi}_{\sigma}(t,x)$ or $\hat{\psi}_{\tau}(t,x)$ is an identity map. Consequently, the order of $\sigma$’s does not affect the definition of $\varphi$. And we can define
$\Phi^{*}\mathrel{\mathop{\ordinarycolon}}=\varphi(1,\cdot)(\Phi_{i}(\cdot))\in\Lambda(\Phi)\,.$
Consequently, by the definition of $\eta_{\sigma}$, for every $x\in X$, we have:
  • There are at most $3^{m}$ cells $\sigma$ of $X_{b}$ such that $\eta_{\sigma}(x)>0$.
  • If two cells $\sigma,\tau$ of $X_{b}$ has $\eta_{\sigma}(x),\eta_{\tau}(x)>0$, then
    $\operatorname{dist}(A_{\sigma},A_{\tau})>0\,.$
  • If $x\in X_{b}$, then there exists a cell $\sigma_{0}$ of $X_{b}$ such that $\eta_{\sigma_{0}}(x)=1$.
Therefore, if $x\in X\setminus D$, we have
$\mathcal{H}^{2}(\Phi^{*}(x))$ $=\sum_{\sigma\mathrel{\mathop{\ordinarycolon}}\eta_{\sigma}(x)>0}\mathcal{H}^{2}(\Phi^{*}(x)\cap A_{\sigma})+\mathcal{H}^{2}\left(\Phi(x)\setminus\bigcup_{\sigma\mathrel{\mathop{\ordinarycolon}}\eta_{\sigma}(x)>0}A_{\sigma}\right)$
$\leq\sum_{\sigma\mathrel{\mathop{\ordinarycolon}}\eta_{\sigma}(x)>0}(\mathcal{H}^{2}(\Phi(x)\cap A_{\sigma})+\delta)+\mathcal{H}^{2}\left(\Phi(x)\setminus\bigcup_{\sigma\mathrel{\mathop{\ordinarycolon}}\eta_{\sigma}(x)>0}A_{\sigma}\right)$
$\leq\mathcal{H}^{2}(\Phi(x))+3^{m}\delta\,.$
On the other hand, if $x\in D\subset X_{b}$, let $\sigma_{0}$ of $X_{b}$ be a cell such that $\eta_{\sigma_{0}}(x)=1$, and then we have
$\mathcal{H}^{2}(\Phi^{*}(x))$ $=\sum_{\sigma\mathrel{\mathop{\ordinarycolon}}\eta_{\sigma}(x)>0}\mathcal{H}^{2}(\Phi^{*}(x)\cap A_{\sigma})+\mathcal{H}^{2}\left(\Phi(x)\setminus\bigcup_{\sigma\mathrel{\mathop{\ordinarycolon}}\eta_{\sigma}(x)>0}A_{\sigma}\right)$
$\leq\sum_{\sigma\mathrel{\mathop{\ordinarycolon}}\sigma\neq\sigma_{0},\ \eta_{\sigma}(x)>0}(\mathcal{H}^{2}(\Phi(x)\cap A_{\sigma})+\delta)+(\mathcal{H}^{2}(\Phi(x)\cap A_{\sigma_{0}})-\varepsilon)$
$\quad+\mathcal{H}^{2}\left(\Phi(x)\setminus\bigcup_{\sigma\mathrel{\mathop{\ordinarycolon}}\eta_{\sigma}(x)>0}A_{\sigma}\right)$
$\leq\mathcal{H}^{2}(\Phi(x))+(3^{m}-1)\delta-\varepsilon\,.$
In addition, for $x\in D^{\prime}\subset X^{\prime}_{b}$, we have that $\eta_{\sigma}(x)=0$ holds for any cell $\sigma$ of $X_{b}$, and thus,
$\Phi^{*}(x)=\Phi(x)\,,$
which confirms the statement (4).
The statement (3) follows from the construction of the map $\varphi$. ∎
4.2. Existence and regularity of a smooth minimal surface
From the definition of width, it is clear that one can choose a minimizing sequence $\{\Phi_{i}\}$ in $\Lambda(\Phi)$.
Recall that any minimizing sequence can be pulled-tight via a family of diffeomorphisms.
Proposition 4.4 (Existence of pulled-tight sequence, [CD03, Proposition 3.1][Pit81, Theorem 4.3]).
Given any minimizing sequence $\{\Phi_{i}\}$ in a homotopy class $\Lambda$, there exists a pulled-tight minimizing sequence $\{\Phi^{*}_{i}\}$ in $\Lambda$ such that
$\mathbf{C}(\{\Phi^{*}_{i}\})\subset\mathbf{C}(\{\Phi_{i}\})\,.$
Proof.
Following Pitts [Pit81, p. 153] (See also [CD03, MN14]), there exists a continuous map
$f\mathrel{\mathop{\ordinarycolon}}[0,1]\times\{V\in\mathcal{V}_{2}(M)\mathrel{\mathop{\ordinarycolon}}\|V\|(M)\leq 2L\}\to\operatorname{Diff}^{\infty}(M)$
such that
  1. (1)
    $f(0,\cdot)=\mathrm{id}$;
  2. (2)
    $f(t,V)=\mathrm{id}$ for all $t\in[0,1]$ if $V$ is stationary;
  3. (3)
    $\|f(1,V)_{\#}V\|(M)<\|V\|(M)$ if $V$ is not stationary.
By Proposition 2.8 ((2)), each $[\Phi_{i}]\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{Z}_{2}(M;\mathbf{F};\mathbb{Z}_{2})$ is continuous, and thus, $|\Phi_{i}|\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{V}_{2}(M)$ is continuous. Hence, for each $i$, we have a continuous map $\varphi_{i}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\operatorname{Diff}^{\infty}(M)$ defined by
$\varphi_{i}(t,x)\mathrel{\mathop{\ordinarycolon}}=f(t,|\Phi_{i}(x)|)\,.$
Let us define $\Phi^{*}_{i}$ by
$\Phi^{*}_{i}(x)\mathrel{\mathop{\ordinarycolon}}=\varphi_{i}(1,\|\Phi_{i}\|(x))(\Phi_{i}(x))\,.$
It follows immediately from the construction that each $\Phi^{*}_{i}$ is homotopic to $\Phi_{i}$, $\mathbf{C}(\{\Phi^{*}_{i}\})\subset\mathbf{C}(\{\Phi_{i}\})$ and every varifold in $\mathbf{C}(\{\Phi^{*}_{i}\})$ is stationary. ∎
For simplicity, let us assume that the minimizing sequence $\{\Phi_{i}\}$ is already pulled-tight. Now we aim to show that
$\mathbf{C}(\{\Phi_{i}\})\cap\mathcal{W}_{L}\neq\emptyset.$
Indeed, by Proposition 4.8, it suffices to show that at least one varifold in $\mathbf{C}(\{\Phi_{i}\})$ is almost minimizing in admissible annuli as defined in Definition 4.6.
Definition 4.5 (Almost minimizing).
For an open subset $U\subset M$ and a sequence of punctate surfaces $\{\Sigma_{j}\}\subset\mathcal{S}^{*}(M)$, a varifold $V\in\mathcal{V}_{2}(M)$ is called almost minimizing with respect to $\{\Sigma_{j}\}$ if there exist two sequences of positive real numbers $\varepsilon_{j}\to 0$, $\delta_{j}\to 0$ such that:
  1. (1)
    $\mathbf{F}(|\Sigma_{j}|,V)<\varepsilon_{j}$.
  2. (2)
    $\Sigma_{j}\in\mathfrak{a}(U;\varepsilon_{i},\delta_{i})$.
Definition 4.6 (Almost minimizing in admissible annuli).
Given $R\in(0,\infty]$, a sequence of punctate surfaces $\{\Sigma_{j}\}\subset\mathcal{S}^{*}(M)$ and a varifold $V\in\mathcal{V}_{2}(M)$, we say that a varifold $V$ is almost minimizing in every $K$-admissible collection of annuli of outer radius at most $R$ with respect to $\{\Sigma_{j}\}$, if there exists $\varepsilon_{j}\to 0$, $\delta_{j}\to 0$ such that:
  1. (1)
    $\mathbf{F}(|\Sigma_{j}|,V)<\varepsilon_{j}$.
  2. (2)
    for any $K$-admissible collection of annuli $\{A_{i}\}^{K}_{i=1}$ each of outer radius at most $R$, and any $\Sigma_{j}$,
    $\Sigma_{j}\in\bigcup^{K}_{i=1}\mathfrak{a}(A_{i};\varepsilon_{j},\delta_{j})\,.$
Remark 4.7.
Suppose that a varifold $V$ is almost minimizing in every $K$-admissible collection of annuli of outer radius at most $R$. Then from the definition, it is clear that for any $K$-admissible collection of annuli $\{A_{i}\}^{K}_{i=1}$ each of radius at most $R$, $V$ is stable in at least one $A_{i}$.
In particular, if $\Sigma=\operatorname{spt}\|V\|$ is a smooth surface, we can say that $\Sigma$ is stable in every $K$-admissible collection of annuli of outer radius at most $R$.
Proposition 4.8 (Regularity of varifolds almost minimizing in admissible annuli).
Given a closed Riemannian manifold $(M,g)$, $K\in\mathbb{N}^{+}$, $R\in(0,+\infty]$, a sequence $\{\Sigma_{j}\}\subset\mathcal{S}(M)$, and a stationary varifold $0\neq V\in\mathcal{V}_{2}(M)$, suppose that:
  1. (i)
    We can choose for each $j\in\mathbb{N}^{+}$ a finite set $P_{j}$ such that $\Sigma_{j}\setminus P_{j}$ is a smooth surface and
    $P=\lim_{j\to\infty}P_{j}$
    in the Hausdorff sense where $P$ is a finite set in $M$.
  2. (ii)
    $V$ is almost minimizing in every $K$-admissible collection of annuli of radius at most $R$ with respect to $\{\Sigma_{j}\}$.
Then $V\in\mathcal{W}_{L}$, where $L=\|V\|(M)$.
Proof.
Let $\{\varepsilon_{j}\}$ and $\{\delta_{j}\}$ be the sequences of positive real numbers for $V$ and $\{\Sigma_{j}\}$ as in Definition 4.6.
For every fixed $p\in M$, we can choose $R(p)\in(0,R)$ such that
$B(p,R(p))\cap(P\setminus\{p\})=\emptyset\,.$
[…]
  1. (1)
    $V$ is almost minimizing in $A(p;s,r)$ with respect to $\{\Sigma_{j_{k_{l}}}\}$.
  2. (2)
    For every $l\in\mathbb{N}^{+}$, $\Sigma_{j_{k_{l}}}\cap A(p;s,r)$ is a smooth surface.
[…]
The following proposition generalizes [CD03, Proposition 5.3] to a multi-parameter Simon-Smith family, using Pitts’s combinatorial argument [Pit81, Proposition 4.9]. It is worth noting that De Lellis-Ramic has proved a similar multi-parameter generalization in [DR18]. However, their homotopy class includes families not derived from isotopies, leading to a slight different definition of almost-minimizing compared to [CD03] and ours. We provide a proof here for the sake of completeness.
Proposition 4.9 (Existence of varifolds almost minimizing in admissible annuli).
Suppose that $X$ is a finite cubical subcomplex of some cube $I(m,k)$, $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}^{*}(M)$ is a Simon-Smith family and $\{\Phi_{i}\}$ is a pulled-tight sequence in $\Lambda(\Phi)$. Then for $K=K(m)\mathrel{\mathop{\ordinarycolon}}=3^{m3^{m}}$ and any $R\in(0,\infty]$, there exists an integer and a min-max subsequence $\Sigma_{j}\mathrel{\mathop{\ordinarycolon}}=\Phi_{i_{j}}(x_{j})$ such that $\Sigma_{j}$ converges to some $W\in\mathbf{C}(\{\Phi_{i}\})$ in the varifold sense. Moreover, $W$ is almost minimizing in every $K$-admissible collection of annuli of outer radius at most $R$ with respect to $\{\Sigma_{j}\}$.
Proof.
We set
$\varepsilon_{i}=\frac{1}{i}\,,\quad\delta_{i}=\frac{\varepsilon_{i}}{2K(m)}\,.$
Then up to a subsequence, we may assume that
$\sup_{x\in X}\mathcal{H}^{2}(\Phi_{i}(x))<\mathbf{L}(\Lambda(\Phi))+\delta_{i}\,.$
Suppose for the sake of contradiction that for some large $i$ and for every $\Phi_{i}(x)$ with $\mathcal{H}^{2}(\Phi_{i}(x))\geq\mathbf{L}(\Lambda(\Phi))-2\varepsilon_{i}$ (and thus, $\Phi_{i}(x)\in\mathcal{S}(M)$), there exists a $K(m)$-admissible collection of annuli $\{A_{x,j}\}^{K}_{j=1}$ of outer radius at most $R$ such that
$\Phi_{i}(x)\notin\bigcup^{K}_{i=1}\mathfrak{a}(A_{x,j};2\varepsilon_{i},\delta_{i}/2)\,.$
It follows from Proposition 4.3 that there exists $\Phi^{*}_{i}\in\Lambda(\Phi)$ such that if $x\in X\setminus D$,
$\mathcal{H}^{2}(\Phi^{*}_{i}(x))$ $\leq\mathcal{H}^{2}(\Phi_{i}(x))+3^{m}\delta_{i}$
$\leq(\mathbf{L}(\Lambda(\Phi))-2\varepsilon_{i})+\varepsilon_{i}/2$
$<\mathbf{L}(\Lambda(\Phi))-\delta_{i}\,.$
On the other hand, if $x\in D$,
$\mathcal{H}^{2}(\Phi^{*}_{i}(x))$ $\leq\mathcal{H}^{2}(\Phi_{i}(x))+(3^{m}-1)\delta_{i}-\varepsilon_{i}$
$<\mathbf{L}(\Lambda(\Phi))-\delta_{i}\,.$
In conclusion, we have $\sup_{x\in X}\mathcal{H}^{2}(\Phi^{*}_{i}(x))<\mathbf{L}(\Lambda(\Phi))-\delta_{i}$, a contradiction. This completes the proof. ∎
Proof of Theorem 2.16 (1) without genus bound.
By Proposition 4.4, there exists a pulled-tight sequence $\{\Phi_{i}\}$ in $\Lambda(\Phi)$.
Then it follows from Proposition 4.9, there exists an integer $K\in\mathbb{N}^{+}$ and a varifold $W\in\mathbf{C}(\{\Phi_{i}\})$ which is almost minimizing in every $K$-admissible collection of annuli with respect to some $\{\Sigma_{j}=\Phi_{i_{j}}(x_{j})\}$.
Finally, as discussed in the second paragraph of Remark 2.11, we can choose $P_{j}$ for each $\Sigma_{j}$ such that
$\sup_{j}\#P_{j}<\infty\,,$
and thus, up to a subsequence, $P_{j}$ converges to a finite set $P$ in the Hausdorff sense. By Proposition 4.8, we can conlude that
$W\in\mathbf{C}(\{\Phi_{i}\})\cap\mathcal{W}_{L}\,.$
4.3. Genus bound
In the previous subsection, we have proved that the existence of a varifold $W\in\mathbf{C}(\{\Phi_{i}\})\cap\mathcal{W}_{L}$ almost minimizing in every $K$-admissible collection of annuli of outer radius at most $R$ with respect to $\{\Sigma_{j}\}$. In this and the next subsection, we take $R=+\infty$.
To justify Theorem 2.16 (1), it suffices to show that this $W$ satisfies the genus estimates as follows.
Proposition 4.10 (Genus bound of varifolds almost minimizing in admissible annuli).
Given a closed Riemannian manifold $(M,g)$, $K\in\mathbb{N}^{+}$, $\mathfrak{g}_{0}\in\mathbb{N}$, $R\in(0,\infty]$, a sequence $\{\Sigma_{j}\}\subset\mathcal{S}(M)$ and a stationary varifold $V\in\mathcal{V}_{2}(M)$, suppose that:
  1. (i)
    We can choose for each $j\in\mathbb{N}^{+}$ a finite set $P_{j}$ such that $\Sigma_{j}\setminus P_{j}$ is a smooth surface and
    $P=\lim_{j\to\infty}P_{j}$
    in the Hausdorff sense where $P$ is a finite set in $M$;
  2. (ii)
    $\sup_{j}\mathfrak{g}(\Sigma_{j})\leq\mathfrak{g}_{0}$;
  3. (iii)
    $V$ is almost minimizing in every $K$-admissible collection of annuli of outer radius at most $R$ with respect to $\{\Sigma_{j}\}$.
Then $V\in\mathcal{W}_{L,\leq\mathfrak{g}_{0}}$ for $L=\|V\|(M)$.
The genus bound result essentially relies on the improved lifting lemma with multiplicity proved by Ketover [Ket19, Proposition 2.2], which we have tailored to our setting in the following proposition. More precisely, we do not assume that the limiting sequence consists of smooth surfaces.
Proposition 4.11 (Improved lifting lemma with multiplicity).
Given a closed Riemannian manifold $(M,g)$, $K\in\mathbb{N}^{+}$, a sequence $\{\Sigma_{j}\}\subset\mathcal{S}(M)$ and a stationary varifold $V\in\mathcal{V}_{2}(M)$, suppose that:
  1. (i)
    We can choose for each $j\in\mathbb{N}^{+}$ a finite set $P_{j}$ such that $\Sigma_{j}\setminus P_{j}$ is a smooth surface and
    $P=\lim_{j\to\infty}P_{j}$
    in the Hausdorff sense where $P$ is a finite set in $M$;
  2. (ii)
    $V$ is almost minimizing in every $K$-admissible collection of annuli with respect to $\{\Sigma_{j}\}$.
Thus, by Proposition 4.8, there exists a disjoint set of smooth, connected, embedded minimal surface $\{\Gamma_{i}\}^{l}_{i=1}$ and a set of positive integers $\{m_{i}\}^{l}_{i=1}$ such that
$V=m_{1}|\Gamma_{1}|+\cdots m_{l}|\Gamma_{l}|\,.$
We denote $\Gamma\mathrel{\mathop{\ordinarycolon}}=\bigcup^{l}_{i=1}\Gamma_{i}$.
Assume that there exists a collection of points $\{q_{i}\in\Gamma_{i}\}^{l}_{i=1}$ and a collection of simple closed curves $\{\gamma_{k}\}^{t}_{k=1}$ contained in $\Gamma\setminus P$ so that for all $k_{1}\neq k_{2}\in\{1,2,\cdots,t\}$, if $\gamma_{k_{2}},\gamma_{k_{2}}\subset\Gamma_{i}$ then $\gamma_{k_{1}}\cap\gamma_{k_{2}}=\{q_{i}\}$. Then there exists $\varepsilon_{0}>0$, so that for any $\varepsilon<\varepsilon_{0}$, there exists curves $\{\widetilde{\gamma}_{k}\}^{t}_{k=1}$, a subsequence of $\{\Sigma_{j}\}$ (still labeled as $\{\Sigma_{j}\}$), and punctate surfaces $\{\widetilde{\Sigma}_{j}\}$ each of which is obtained from $\Sigma_{j}$ by finitely many neck-pinch surgeries outside $P_{j}$ with the following properties.
  1. (1)
    Each $\widetilde{\gamma}_{k}$ is homotopic to $\gamma_{k}$ in $\Gamma$ and $\widetilde{\gamma}_{k}\subset T_{\varepsilon}(\gamma_{k})$;
  2. (2)
    $|\widetilde{\Sigma}_{j}|\to V$ as varifolds;
  3. (3)
    For each $k\in\{1,2,\cdots,t\}$, if $\widetilde{\gamma}_{k}\subset\Gamma_{i}$, then exactly one of the following holds:
    1. (1)
      $\pi^{-1}(\widetilde{\gamma}_{k})\cap T_{\varepsilon}(\Gamma_{i})\cap\widetilde{\Sigma}_{j}$ is a union of $m_{i}$ closed curves, each of which projects via $p$ onto $\widetilde{\gamma}_{i}$ with degree one.
    2. (2)
      $\pi^{-1}(\widetilde{\gamma}_{k})\cap T_{\varepsilon}(\Gamma_{i})\cap\widetilde{\Sigma}_{j}$ is a union of $m_{i}/2$ closed curves, each of which projects via the closest-point projection $p$ onto $\widetilde{\gamma}_{k}$ with degree two. In this case, $\Gamma_{i}$ is non-orientable and $\pi^{-1}(\widetilde{\gamma}_{k})$ is a M obius band.
    Here, $\pi\mathrel{\mathop{\ordinarycolon}}T_{\varepsilon}(\Gamma_{i})\to\Gamma_{i}$ is the closest-point projection onto $\Gamma_{i}$.
Indeed, in [Ket19, Section 4], the proof of lifting lemma, the only point where Ketover essentially relies on the smoothness of the sequence $\{\Sigma_{i}\}$ is in the application of the following Genus Collapse Lemma by Colding-Minicozzi.
Lemma 4.12 (Genus Collapse, [CM04, Lemma I.0.14]).
[…]
$\mathfrak{g}(\Sigma_{j}\cap B_{r_{x}}(x))=0\,.$
Remark 4.13.
In the original statement in Colding-Minicozzi [CM04], it appears they require each $\Sigma_{j}$ is minimal; However, following their proof, as in Ketover’s cited statement [Ket19, Lemma 3.4], this minimality condition is unnecessary.
In addition, the requirement by Colding-Minicozzi that the ambient manifold $(M,g)$ be closed is also unnecessary.
Proof of Proposition 4.11.
We choose $\varepsilon_{0}$ as in [Ket19, Set up 4.2] and fix $\varepsilon<\varepsilon_{0}$.
We choose $s\in(0,\operatorname{dist}(P,\bigcup_{k}\gamma_{k}))$ so that for every $p\in P$, $B_{s}(p)\cap\Gamma$ is either a disk or an empty set. Then we set
$\widetilde{M}\mathrel{\mathop{\ordinarycolon}}=M\setminus B_{s}(P)\,.$
By possibly choosing a subsequence, without loss of generality, we may assume that every $\Sigma_{j}\cap\widetilde{M}$ is a smooth surface.
Note that all the arguments in [Ket19, Section 4] are confined to a neighborhood of $\bigcup_{k}\gamma_{k}$. With Lemma 4.12, we can follow these arguments verbatim to prove our proposition. ∎
Proof of Proposition 4.10.
In the previous subsection, we prove $V\in\mathcal{W}_{L}$, i.e.,
$V=m_{1}|\Gamma_{1}|+\cdots+m_{l}|\Gamma_{l}|$
for a disjoint set of smooth, connected, embedded minimal surface $\{\Gamma_{i}\}^{l}_{i=1}$ and a set of positive integers $\{m_{i}\}^{l}_{i=1}$.
We follow the proof of [Ket19, Theorem 1.2] to select a set of curves $\{\alpha_{k}\}^{t}_{k=1}\subset\Gamma$. Since #$\#P<\infty$, we may choose $\{\alpha_{k}\}^{t}_{k=1}\subset\Gamma\setminus P$. This allows us to apply our Proposition 4.11 to obtain a sequence of punctate surfaces $\{\widetilde{\Sigma}_{j}\}$. However, the only issue to continue the arguments from [Ket19] is that our $\widetilde{\Sigma}_{j}$ is not necessarily smooth everywhere.
To address this issue, for each $j$, we choose $s_{j}\to 0$ such that we can replace $\widetilde{\Sigma}_{j}\cap B_{s_{j}}(P_{j})=\Sigma_{j}\cap B_{s_{j}}(P_{j})$ in $\widetilde{\Sigma}_{j}$ with finitely many disks through surgeries and the removal of connected components, thereby obtaining a sequence of smooth surfaces $\{\widetilde{\Sigma}^{\prime}_{j}\}$. For sufficiently large $j$, we have:
  1. (1)
    $B_{s_{j}}(P_{j})\cap\pi^{-1}(\bigcup^{t}_{k=1}\alpha_{k})=\emptyset$;
  2. (2)
    $\widetilde{\Sigma}^{\prime}_{j}\setminus B_{s_{j}}(P_{j})=\widetilde{\Sigma}_{j}\setminus B_{s_{j}}(P_{j})$;
  3. (3)
    $|\widetilde{\Sigma}^{\prime}_{j}|\to W$ as varifolds;
  4. (4)
    $\mathfrak{g}(\widetilde{\Sigma}^{\prime}_{j})\leq\mathfrak{g}(\widetilde{\Sigma}_{j})\leq\mathfrak{g}(\Sigma_{j})\leq\mathfrak{g}_{0}$.
Consequently, with the sequence of smooth surfaces $\{\widetilde{\Sigma}^{\prime}_{j}\}$, we can follow the proof of [Ket19, Theorem 1.2] verbatim to obtain
(4.1) $\sum^{l}_{i=1}m_{i}\mathfrak{g}(\Gamma_{i})\leq\mathfrak{g}_{0}\,.$
4.4. Multiplicity one theorem
The Multiplicity one result is an adaption of Theorem B in [WZ24], which essentially follows from the PMC min-max theorem, Theorem 2.4 therein.
4.4.1. PMC Simon-Smith min-max theory
Let us first recall the concepts of $\mathcal{A}^{h}$-functional, $\mathcal{V}\mathcal{C}$-space, $C^{1,1}$ boundary and strong $\mathcal{A}^{h}$-stationarity introduced in [WZ24].
Definition 4.14 (Prescribed mean curvature functional).
For a function $h\in C^{\infty}(M)$, we define the prescribed mean curvature functional associated with $h$ to be
[…]
Definition 4.15 ($\mathcal{A}^{h}$-stationary pairs).
Given an open subset $U\subset M$, a pair $(V,\Omega)\in\mathcal{V}_{2}(M)\times\mathcal{C}(M)$ is $\mathcal{A}^{h}$-stationary in $U$, if for any vector field $X\in\mathcal{X}(U)$ with associated flow $\varphi^{t}$,
$\delta\mathcal{A}^{h}(V,\Omega)(X)\mathrel{\mathop{\ordinarycolon}}=\frac{d}{dt}|_{t=0}\mathcal{A}^{h}(\varphi^{t}_{\#}(V,\Omega))=0\,.$
An $\mathcal{A}^{h}$-stationary pair $(V,\Omega)$ is $\mathcal{A}^{h}$-stable in $U$ if if for any vector field $X\in\mathcal{X}(U)$ with associated flow $\varphi^{t}$,
$\delta^{2}\mathcal{A}^{h}(V,\Omega)(X,X)\mathrel{\mathop{\ordinarycolon}}=\frac{d^{2}}{dt^{2}}|_{t=0}\mathcal{A}^{h}(\varphi^{t}_{\#}(V,\Omega))\geq 0\,.$
Definition 4.16 ($\mathcal{V}\mathcal{C}$-space).
The $\mathcal{V}\mathcal{C}$-space on $M$, denoted by $\mathcal{V}\mathcal{C}(M)$, is the space of all pairs $(V,\Omega)\in\mathcal{V}_{2}(M)\times\mathcal{C}(M)$ such that their is a sequence $\{\Omega_{k}\}\subset\mathcal{C}(M)$ satisfying
$|\partial^{*}\Omega_{k}|\to V\,,\quad\Omega_{k}\to\Omega\,.$
Given two pairs $(V,\Omega)$ and $(V^{\prime},\Omega^{\prime})$ in $\mathcal{V}\mathcal{C}(M)$, the $\mathscr{F}$-distance between them is
$\mathscr{F}((V,\Omega),(V^{\prime},\Omega^{\prime}))\mathrel{\mathop{\ordinarycolon}}=\mathbf{F}(V,V^{\prime})+\mathcal{F}(\Omega,\Omega^{\prime})\,.$
Definition 4.17 ($C^{1,1}$ almost embedding).
Given an open subset $U\subset M$, a $C^{1,1}$ immersed surface $\phi\mathrel{\mathop{\ordinarycolon}}\Sigma\to U$ with $\phi(\partial\Sigma)\cap U=\emptyset$ is called a $C^{1,1}$ almost embedded surface in $U$ provided that at every point $p\in\phi(\Sigma)$ where $\phi$ is not an embedding, there exists a neighborhood $W\subset U$ of $p$, such that:
  1. (1)
    $\Sigma\cap\phi^{-1}(W)$ is a disjoint union of connected components $\bigsqcup^{\ell}_{i=1}\Gamma^{i}$.
  2. (2)
    $\phi\mathrel{\mathop{\ordinarycolon}}\Gamma^{i}\to W$ is a $C^{1,1}$ embedding for each $i$.
  3. (3)
    For each pair $i\neq j$ , $\phi(\Gamma^{j})$ lies on one-side of $\phi(\Gamma_{i})$ in $W$.
We call the set of all such $p$ as the touching set of $\Sigma$. For simplicity, we will denote $\phi(\Sigma)$ by $\Sigma$ and $\phi(\Gamma^{i})$ by $\Gamma^{i}$.
Definition 4.18 ($C^{1,1}$ boundary).
Given an open subset $U$ of $M$, a $C^{1,1}$ almost embedded surface $\phi\mathrel{\mathop{\ordinarycolon}}\Sigma\to U$ and $\Omega\in\mathcal{C}(U)$, $(\Sigma,\Omega)$ is called a $C^{1,1}$ boundary in $U$, provided that $\Sigma$ is orientable and
$\phi_{\#}([\Sigma])=\partial^{*}\Omega\,.$
For a function $h\in C^{\infty}(M)$, a $C^{1,1}$ boundary $(\Sigma,\Omega)$ in $U$ is called a $C^{1,1}$ $h$-boundary if $(|\Sigma|,\Omega)$ is $\mathcal{A}^{h}$-stationary in $U$.
Definition 4.19 (Strong $\mathcal{A}^{h}$-stationarity).
Given an open subset $U$ of $M$, a $C^{1,1}$ $h$-boundary $(\Sigma,\Omega)$ is said to be strongly $\mathcal{A}^{h}$-stationary in $U$ provided the following holds.
For every $p\in U$ in the touching set of $\Sigma$, there exists a neighborhood $W\subset U$ of $p$, and a decomposition $\Sigma\cap W=\bigcup^{\ell}_{i=1}\Gamma^{i}$ into :$\ell\mathrel{\mathop{\ordinarycolon}}=\Theta^{2}(\Sigma,p)\geq 2$ connected disks with a natural ordering
$\Gamma^{1}\leq\Gamma^{2}\leq\cdots\leq\Gamma^{\ell}\,.$
Let $W^{1}$ and $W^{\ell}$ be the bottom and the top components of $W\setminus\Sigma$. For $i=1$ or $\ell$ and all $X\in\mathcal{X}(W)$ pointing into $W^{i}$ along $\Gamma^{i}$, if $W^{i}\subset\Omega$, then
$\delta\mathcal{A}^{h}(\Gamma^{i},W^{i})(X)\geq 0\,;$
otherwise, $W^{i}\cap\Omega=\emptyset$ and
$\delta\mathcal{A}^{h}(\Gamma^{i},W\setminus W^{i})(X)\geq 0\,.$
To adapt their PMC Simon-Smith min-max theory to our setting, note that for every $\Sigma\in\mathcal{S}^{*}(M)$, by definitions, $[\Sigma]$ can be viewed as the reduced boundary of a Caccioppoli set $\Omega\in\mathcal{C}(M)$ and its complement $M\setminus\Omega\in\mathcal{C}(M)$. Therefore, we can extend the definition of $\mathscr{E}$ in [WZ24, Section 2] to
~:=$\widetilde{\mathscr{E}}\mathrel{\mathop{\ordinarycolon}}=\{(\Sigma,\Omega)\mathrel{\mathop{\ordinarycolon}}\Sigma\in\mathcal{S}^{*}(M),\ [\Sigma]=\partial^{*}\Omega\}\,.$
Definition 4.20 (PMC Simon-Smith family).
Let $X$ be a cubical subcomplex of some $I(m,j)$. A map $(\Phi,\Omega)\mathrel{\mathop{\ordinarycolon}}X\to\widetilde{\mathscr{E}}$ is called a PMC Simon-Smith family, provided that:
  1. (1)
    Its first component $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}^{*}(M)$ is a Simon-Smith family.
  2. (2)
    Its second component $\Omega\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{C}(M)$ is continuous.
Definition 4.21 (Relative PMC homotopy class).
Let $Z\subset X$ be cubical subcomplexes of some $I(m,j)$. Two PMC Simon-Smith families $(\Phi_{0},\Omega_{0})$ and $(\Phi_{1},\Omega_{1})$ parametrized by the same parameter space $X$ with $(\Phi_{0},\Omega_{0})|_{Z}=(\Phi_{1},\Omega_{1})|_{Z}$ are said to be homotopic relative to $(\Phi_{0},\Omega_{0})|_{Z}$ to each other if there exists a continuous map
$\varphi\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\operatorname{Diff}^{\infty}(M)$
such that
  1. (1)
    $\varphi(0,x)=\operatorname{Id}$ for all $x\in X$.
  2. (2)
    $\varphi(t,z)=\operatorname{Id}$ for all $t\in[0,1]$ and $z\in Z$.
  3. (3)
    $(\varphi(1,x)(\Phi_{0}(x)),\varphi(1,x)(\Omega_{0}(x)))=(\Phi_{1}(x),\Omega_{1}(x))$ for all $x\in X$.
[…]
Definition 4.22.
Given a $(X,Z)$-relative homotopy class $\Lambda_{Z}$ of a PMC Simon-Smith family and a function $h\in C^{\infty}(M)$, its $h$-width is defined by
$\mathbf{L}^{h}(\Lambda_{Z})\mathrel{\mathop{\ordinarycolon}}=\inf_{(\Phi,\Omega)\in\Lambda_{Z}}\sup_{x\in X}\mathcal{A}^{h}(|\Phi|(x),\Omega(x))\,.$
Definition 4.23.
A sequence $\{(\Phi_{i},\Omega_{i})\}$ in $\Lambda_{Z}$ is said to be minimizing associated with $h\in C^{\infty}(M)$ if
$\lim_{i\to\infty}\sup_{x\in X}\mathcal{A}^{h}(|\Phi|(x),\Omega(x))=\mathbf{L}^{h}(\Lambda_{Z})\,.$
If $\{(\Phi_{i}\Omega_{i})\}$ is a minimizing sequence associated with $h$ and $\{x_{i}\}\subset X$ satisfies
$\lim_{i\to\infty}\mathcal{A}^{h}(|\Phi_{i}(x_{i})|,\Omega_{i}(x_{i}))=\mathbf{L}^{h}(\Lambda_{Z})\,,$
then $\{\Phi_{i}(x_{i})\}$ is called a min-max sequence associated with $h$.
For a minimizing sequence $\{(\Phi_{i},\Omega_{i})\}$, we define its critical set associated with $h$ to be the set of all subsequential varifold-limit of its min-max sequences:
$\mathbf{C}^{h}(\{(\Phi_{i},\Omega_{i})\})\mathrel{\mathop{\ordinarycolon}}=\{(V,\Omega)=\lim_{j}(|\Phi_{i_{j}}(x_{j})|,\Omega_{i_{j}}(x_{j}))\in$ $\mathcal{V}\mathcal{C}(M)\mathrel{\mathop{\ordinarycolon}}x_{j}\in X,\,$
$\mathcal{A}^{h}(V,\Omega)=\mathbf{L}^{h}(\Lambda_{Z})\}\,.$
Definition 4.24 ($(\mathcal{A}^{h},\varepsilon,\delta)$-deformation).
Given $\varepsilon,\delta>0$, an open set $U\subset M$, a function $h\in C^{\infty}(M)$ and a pair $(\Sigma,\Omega)\in\widetilde{\mathscr{E}}$, we call an isotopy $\psi\in\mathfrak{Is}(U)$ an $(\mathcal{A}^{h},\varepsilon,\delta)$-deformation of $\Sigma$ in $U$ associated with $h$ provided that:
  1. (1)
    $\mathcal{A}^{h}(\psi(t)_{\#}(|\Sigma|,\Omega))\leq\mathcal{A}^{h}(|\Sigma|,\Omega)+\delta$ for all $t\in[0,1]$.
  2. (2)
    $\mathcal{A}^{h}(\psi(t)_{\#}(|\Sigma|,\Omega))\leq\mathcal{A}^{h}(|\Sigma|,\Omega)-\varepsilon$.
We define $\mathfrak{a}^{h}(U;\varepsilon,\delta)$ to be the set of all pairs in $\widetilde{\mathscr{E}}$ that do not admit $(\mathcal{A}^{h},\varepsilon,\delta)$-deformations in $U$.
Definition 4.25 (Almost minimizing in admissible annuli).
Given a sequence of pairs $\{(\Sigma_{j},\Omega_{j})\}\subset\widetilde{\mathscr{E}}$ and a pair $(V,\Omega)\in\mathcal{V}\mathcal{C}(M)$, we say that $(V,\Omega)$ is $\mathcal{A}^{h}$-almost minimizing in every $K$-admissible collection of annuli of outer radius at most $R$ with respect to $\{(\Sigma_{j},\Omega_{j})\}$, if there exists $\varepsilon_{j}\to 0$, $\delta_{j}\to 0$ such that:
  1. (1)
    $\mathscr{F}((|\Sigma_{j}|,\Omega_{j}),(V,\Omega))<\varepsilon_{j}$.
  2. (2)
    for any $K$-admissible collection of annuli $\{A_{i}\}^{K}_{i=1}$, and any pair $(\Sigma_{j},\Omega_{j})$,
    $(\Sigma_{j},\Omega_{j})\in\bigcup^{K}_{i=1}\mathfrak{a}^{h}(A_{i};\varepsilon_{j},\delta_{j})\,.$
Remark 4.26.
Similar to Remark 4.7, if $\Sigma=\operatorname{spt}\|V\|$ is a minimal surface, then for every $K$-admissible collection of annuli, $(|\Sigma|,\Omega)$ is $\mathcal{A}^{h}$-stable in at least one annulus.
Theorem 4.27 (Relative PMC min-max theorem, [WZ24, Theorem 2.4]).
Given $Z\subset X$ cubical subcomplexes of some $I(m,j)$, a function $h\in C^{\infty}(M)$ and a PMC Simon-Smith family $(\Phi_{0},\Omega_{0})$, suppose that
$\mathbf{L}^{h}(\Lambda_{Z}(\Phi_{0},\Omega_{0}))>\max\{\sup_{z\in Z}\mathcal{A}^{h}(|\Phi_{0}(z)|,\Omega_{0}(z)),0\}\,.$
Then there exists a minimizing sequence $\{(\Phi_{i},\Omega_{i})\}\subset\Lambda_{Z}(\Phi_{0},\Omega_{0})$, and a strongly $\mathcal{A}^{h}$-stationary, $C^{1,1}$ $h$-boundary $(\Sigma,\Omega)$ with $(|\Sigma|,\Omega)\in\mathbf{C}^{h}(\{(\Phi_{i},\Omega_{i})\})$, such that
$\mathcal{A}^{h}(\Sigma,\Omega)=\mathbf{L}^{h}(\Lambda_{Z}(\Phi_{0},\Omega_{0}))\,.$
Moreover, there exists an integer $K=K(m)$ and a min-max subsequence $(\Sigma_{j},\Omega_{j})\mathrel{\mathop{\ordinarycolon}}=(\Phi_{i_{j}}(x_{j}),\Omega_{i_{j}}(x_{j})$ such that $(|\Sigma|,\Omega)$ is $\mathcal{A}^{h}$-almost minimizing in every $K$-admissible collection of annuli with respect to $\{(\Sigma_{j},\Omega_{j})\}$.
Proof.
Firstly, as in the proof of our Proposition 4.9, one can adapt the proof of [WZ24, Theorem 3.8] to obtain a pair $(V_{0},\Omega_{0})\in\mathcal{V}\mathcal{C}(M)$ which is $\mathcal{A}^{h}$-almost minimizing in small annuli with respect to $\{(\Sigma_{j},\Omega_{j})=(\Phi_{i_{j}}(x_{j}),(\Omega_{i_{j}}(x_{j}))\}$.
Then, as discussed in the second paragraph of Remark 2.11, we can choose $P_{j}$ for each $\Sigma_{j}$ such that
$\sup_{j}\#P_{j}<\infty$
and thus, up to a subsequence, $P_{j}$ converges to a finite set $P$ in the Hausdorff sense. Since $\Sigma_{j}\setminus P_{j}$ is a smooth surface, one can follow verbatim the proof of [WZ24, Theorem 4.7] and conclude that $(V_{0},\Omega_{0})$ is $C^{1,1}$ and strongly $\mathcal{A}^{h}$-stationary boundary in $M\setminus P^{\prime}$ for some finite set $P^{\prime}\supset P$.
Finally, one can follow their $C^{1,1}$-version of removal singularity as in the proof of [WZ24, Theorem 2.4] to conclude the theorem. ∎
4.4.2. Proof of Theorem 2.16 (2)
Let $\mathcal{W}^{m}_{L,\leq\mathfrak{g}_{0}}$ be the set of all varifolds $W\in\mathcal{W}_{L,\leq\mathfrak{g}_{0}}$, whose support is a smooth embedded minimal surface $\Sigma$, such that for every $K(m)\mathrel{\mathop{\ordinarycolon}}=3^{m3^{m}}$-admissible collection of annuli, $\Sigma$ is stable in at least one annulus. Note that this is just a reformulation of Property (R’) in [WZ24]. Proposition 4.8 with $R=+\infty$ implies that
$W\in\mathbf{C}(\{\Phi_{i}\})\cap\mathcal{W}^{m}_{L,\leq\mathfrak{g}_{0}}\,.$
By Sharp’s compactness theorem [Sha17] (See also Theorem [WZ24, Theorem 7.1]), the set $\mathcal{W}^{m}_{L,\leq\mathfrak{g}_{0}}$ is compact. In particular, if $g$ is a bumpy metric on $M$, then $\mathcal{W}^{m}_{L,\leq\mathfrak{g}_{0}}$ is a finite set.
As in the proof of [WZ24, Theorem 7.3], it suffices to work on bumpy metrics $g$ and the general results follow from a approximation by bumpy metrics.
By Proposition 2.8, for any Simon-Smith family $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}^{*}(M)$, the map $[\Phi]$ is continuous in the $\mathbf{F}$-metric, and thus, it can be lifted to a PMC Simon-Smith family
$(\widetilde{\Phi},\widetilde{\Omega})\mathrel{\mathop{\ordinarycolon}}\widetilde{X}\to\widetilde{\mathscr{E}}\,,$
where $\widetilde{X}$ is a double-cover of $X$.
Hence, one can follow verbatim the proofs of Theorem 7.2 and Theorem 7.3 of [WZ24] through replacing Theorem 2.4 therein by our Theorem 4.27. Note that by Remark 4.26, Theorem 4.27 also implies the required Property (R’) as defined in [WZ24, Corollary 3.11]. This completes the proof of Theorem 2.16 (2).
4.5. Refinement of the critical set by Marques-Neves
Conclusion (3) is adapted from [MN21, Theorem 4.7], which essentially follows from the combinatorial argument of [Pit81, Theorem 4.10]. Note that our Definition 4.1 of $(\varepsilon,\delta)$ deformation is continuous, so we do not need to appeal to the intricate interpolation lemmas as in [MN21].
Let us first recall a useful lemma proved in Marques-Neves [MN21, Lemma 4.5].
Lemma 4.28.
For every $r,L>0$ and $m\in\mathbb{N}$, there exists $\widetilde{\eta}=\widetilde{\eta}(r,L,m)>0$ so that for every $V,T\in\mathcal{V}_{2}(M)$ with $\mathbf{M}(V)\leq 2L,\mathbf{M}(T)\leq 2L$, $V$ stationary,
$V\llcorner(M\setminus(\overline{B}_{2\widetilde{\eta}}(p_{1})\cup\cdots\cup\overline{B}_{2\widetilde{\eta}}(p_{t})))=T\llcorner(M\setminus(\overline{B}_{2\widetilde{\eta}}(p_{1})\cup\cdots\cup\overline{B}_{2\widetilde{\eta}}(p_{t})))$
for some collection $\{p_{1},\cdots,p_{t}\}\subset M$, $t\leq K(m)\mathrel{\mathop{\ordinarycolon}}=3^{m3^{m}}$, and
$\|V\|(M)-\widetilde{\eta}\leq\|T\|(M)\leq\|V\|(M)+\widetilde{\eta}\,,$
then $\mathbf{F}(V,T)<r/4$.
With this lemma, we can fix $\widetilde{\eta}$ from Lemma 4.28 using the parameters $r,L,m$ from Theorem 2.16.
As discussed in Remark 4.7, the support of the varifold generated from the min-max theory is also stable in admissible annuli. Let $\mathcal{W}^{m,\widetilde{\eta}}_{L,\leq\mathfrak{g}_{0}}$ be the set of all varifolds $W\in\mathcal{W}_{L,\leq\mathfrak{g}_{0}}$, whose support is a smooth embedded minimal surface $\Sigma$, such that for every $K(m)\mathrel{\mathop{\ordinarycolon}}=3^{m3^{m}}$-admissible collection of annuli of outer radius at most $\widetilde{\eta}$, $\Sigma$ is stable in at least one annulus.
Let $\{\Phi_{i}\}$ be a pulled-tight sequence from Theorem 2.16 (1). In the following, we will deform $\{\Phi_{i}\}$ and obtain a new pulled-tight sequence $\{\widetilde{\Phi}_{i}\}$ such that there exists $\eta>0$ such that for all sufficiently large $i$,
(4.2) $\mathcal{H}^{2}(\Phi^{*}_{i}(x))\geq L-\eta\implies|\Phi^{*}_{i}(x)|\in\mathbf{B}^{\mathbf{F}}_{r}(\mathcal{W}^{m,\widetilde{\eta}}_{L,\leq\mathfrak{g}_{0}})\,.$
Note that
$\mathcal{W}^{m}_{L,\leq\mathfrak{g}_{0}}\subset\mathcal{W}^{m,\widetilde{\eta}}_{L,\leq\mathfrak{g}_{0}}\subset\mathcal{W}_{L,\leq\mathfrak{g}_{0}}\,.$
It follows from Remark 4.7 and Proposition 4.10 (with $R=\widetilde{\eta}$) that for each $V\in\mathbf{C}(\{\Phi_{i}\})\setminus\mathcal{W}^{m,\widetilde{\eta}}_{L,\leq\mathfrak{g}_{0}}$, there exists a $K(m)\mathrel{\mathop{\ordinarycolon}}=3^{m3^{m}}$-admissible collection of annuli of outer radius at most $\widetilde{\eta}$, denoted by $\{A_{V,j}\}^{K(m)}_{j=1}$, such that $V$ is not almost-minimizing in any $A_{V,j}$ with respect to any min-max subsequence. In particular, for every $V\in\mathbf{C}(\{\Phi_{i}\})\setminus\mathcal{W}^{m,\widetilde{\eta}}_{L,\leq\mathfrak{g}_{0}}$, there exists $\varepsilon_{V}>0$ and $N_{V}\in\mathbb{N}^{+}$ with the following property: For every $\delta>0$ and every $\Phi_{i}(x)$ with $i\geq N_{V}$, if $\mathbf{F}(V,|\Phi_{i}(x)|)<\varepsilon_{V}$, then
$\Phi_{i}(x)\notin\bigcup^{K}_{j=1}\mathfrak{a}(A_{V,j};\varepsilon_{V},\delta)\,.$
Let $\mathcal{K}\mathrel{\mathop{\ordinarycolon}}=\mathbf{C}(\{\Phi_{i}\})\setminus\mathbf{B}^{\mathbf{F}}_{r/2}(\mathcal{W}^{m,\widetilde{\eta}}_{L,\leq\mathfrak{g}_{0}})$. Since $\mathcal{K}$ is compact, we can find a finite set $\{V_{k}\}^{\nu}_{k=1}\subset\mathcal{K}$ such that
$\mathcal{K}\subset\bigcup^{\nu}_{k=1}\mathbf{B}^{\mathbf{F}}_{\frac{\varepsilon_{k}}{4}}(V_{k})\,,$
where $\varepsilon_{k}\mathrel{\mathop{\ordinarycolon}}=\varepsilon_{V_{k}}$. We set
$\varepsilon^{(1)}\mathrel{\mathop{\ordinarycolon}}=\frac{\min_{k}\varepsilon_{k}}{10}\,,\quad N^{(1)}\mathrel{\mathop{\ordinarycolon}}=\max_{k}N_{V_{k}}\,.$
Since $\{\Phi_{i}\}$ is pulled-tight, we can choose an integer $N^{(2)}\geq N^{(1)}$ and $\varepsilon^{(2)}\in(0,\varepsilon^{(1)})$ such that for each $i\geq N^{(2)}$, and any $x\in X$, if
$\mathcal{H}^{2}(\Phi_{i}(x))\geq\sup_{y\in X}\mathcal{H}^{2}(\Phi_{i}(y))-2\varepsilon^{(2)}\,,\quad\mathbf{F}(|\Phi_{i}(x)|,\mathcal{W}^{m,\widetilde{\eta}}_{L,\leq\mathfrak{g}_{0}})\geq\frac{r}{2}\,,$
then $\mathbf{F}(|\Phi_{i}(x)|,V_{k_{i,x}})<\frac{\varepsilon_{k_{i,x}}}{2}$ for some $k_{i,x}\in\{1,2,\cdots,\nu\}$, and thus, for every $\delta>0$, we have
(4.3) $\Phi_{i}(x)\notin\bigcup^{K}_{j=1}\mathfrak{a}(A_{V_{k_{i,x}},j};\varepsilon_{V_{k_{i,x}}},\delta)\,.$
We denote by $D_{i}$ the set of all such $x$ (“bad points”) for each $i\geq N^{(2)}$.
We set
$\varepsilon\mathrel{\mathop{\ordinarycolon}}=\frac{\min\left\{\varepsilon^{(1)},\varepsilon^{(2)},r,\widetilde{\eta}\right\}}{10}\,.$
For each $i\geq N^{(2)}$, we also choose $\delta_{i}>0$ such that
$\lim_{i}\delta_{i}=0\,.$
By (4.3), for every $i\geq N^{(2)}$, if $x\in D_{i}$, then
(4.4) $\Phi_{i}(x)\notin\bigcup^{K}_{j=1}\mathfrak{a}(A_{V_{k_{i,x}},j};2\varepsilon^{(1)},\delta_{i}/2)\,.$
Hence, by Proposition 4.3 with the parameters $R=\widetilde{\eta}$, $\varepsilon=\varepsilon^{(1)}$, $\delta=\delta_{i}$ and $D=D_{i}$, we obtain $\Phi^{*}_{i}\in\Lambda(\Phi)$. Moreover, for every $x\in X$,
$\mathcal{H}^{2}(\Phi^{*}_{i}(x))\leq\mathcal{H}^{2}(\Phi^{*}_{i}(x))+3^{m}\cdot\delta_{i}\,.$
Since $\delta_{i}\to 0$, $\{\Phi^{*}_{i}\}$ is also a minimizing sequence.
To show that $\mathbf{C}(\{\Phi^{*}_{i}\})\subset\mathbf{B}^{\mathbf{F}}_{\frac{3}{4}r}(\mathcal{W}^{m,\widetilde{\eta}}_{L,\leq\mathfrak{g}_{0}})$, we need to prove the following claim.
Claim 4.29.
For sufficiently large $i$, if $\mathcal{H}^{2}(\Phi^{*}_{i}(x))\geq\sup_{y\in X}\mathcal{H}^{2}(\Phi_{i}(y))-\varepsilon/4$, then
$\mathbf{F}(|\Phi_{i}(x)|,\mathcal{W}^{m,\widetilde{\eta}}_{L,\leq\mathfrak{g}_{0}})<r/2\,.$
Proof.
Since $\lim_{i}\delta_{i}=0$, for sufficiently large $i$, we have
$\delta_{i}<\frac{\varepsilon}{1000K(m)}\,.$
By the definition of $D_{i}$, for $x\in X$, if $\mathbf{F}(|\Phi_{i}(x)|,\mathcal{W}^{m,\widetilde{\eta}}_{L,\leq\mathfrak{g}_{0}})\geq r/2$, we have either $\Phi_{i}(x)\in D_{i}$ or $\mathcal{H}^{2}(\Phi_{i}(x))<\sup_{y\in X}\mathcal{H}^{2}(\Phi_{i}(y))-2\varepsilon^{(2)}$.
If $\Phi_{i}(x)\in D_{i}$, i.e.,
$\mathcal{H}^{2}(\Phi_{i}(x))\geq\sup_{y\in X}\mathcal{H}^{2}(\Phi_{i}(y))-2\varepsilon^{(2)}\,,\quad\mathbf{F}(|\Phi_{i}(x)|,\mathcal{W}^{m,\widetilde{\eta}}_{L,\leq\mathfrak{g}_{0}})\geq r/2\,,$
by Proposition 4.3, we have
$\mathcal{H}^{2}(\Phi^{*}_{i}(x))\leq\mathcal{H}^{2}(\Phi_{i}(x))-\varepsilon^{(2)}+(3^{m}-1)\delta_{i}<\sup_{y}\mathcal{H}^{2}(\Phi_{i}(y))-\varepsilon/4\,.$
If $\mathcal{H}^{2}(\Phi_{i}(x))<\sup_{y\in X}\mathcal{H}^{2}(\Phi_{i}(y))-2\varepsilon^{(2)}$, then by Proposition 4.3 again, we have
$\mathcal{H}^{2}(\Phi^{*}_{i}(x))\leq\mathcal{H}^{2}(\Phi_{i}(x))+3^{m}\cdot\delta_{i}<\sup_{y\in X}\mathcal{H}^{2}(\Phi_{i}(y))-2\varepsilon^{(2)}+3^{m}\cdot\delta_{i}<\mathcal{H}^{2}(\Phi_{i}(x))-\varepsilon/4\,.$
This finishes the proof of the claim. ∎
Now, every $W\in\mathbf{C}(\{\Phi^{*}_{i}\})$ is a varifold limit of some subsequence $\Phi^{*}_{i_{j}}(x_{i_{j}})$ for $j\to\infty$. After passing to a subsequence, $\Phi_{i_{j}}(x_{i_{j}})$ converges to a varifold $V$. By Proposition 4.3, we know that there exist some points $p_{1},\cdots,p_{t}\in M$ ($t\leq K(m)$) such that
$V\llcorner(M\setminus(\overline{B}_{2\widetilde{\eta}}(p_{1})\cup\cdots\cup\overline{B}_{2\widetilde{\eta}}(p_{t})))=W\llcorner(M\setminus(\overline{B}_{2\widetilde{\eta}}(p_{1})\cup\cdots\cup\overline{B}_{2\widetilde{\eta}}(p_{t})))\,.$
Furthermore, since
$L\geq\|V\|(M)=\lim_{j\to\infty}\mathcal{H}^{2}(\Phi_{i_{j}}(x_{i_{j}}))$ $\geq\lim_{j\to\infty}\mathcal{H}^{2}(\Phi^{*}_{i_{j}}(x_{i_{j}}))-K(m)\cdot\delta_{i_{j}}$
$=\|W\|(M)=L\,,$
by Lemma 4.28, we can conlude that
$\mathbf{F}(V,W)<r/4\,.$
The previous claim implies that $\mathbf{F}(V,\mathcal{W}^{m,\widetilde{\eta}}_{L,\leq\mathfrak{g}_{0}})<r/2$, so we have
$\mathbf{F}(W,\mathcal{W}^{m,\widetilde{\eta}}_{L,\leq\mathfrak{g}_{0}})<\frac{3}{4}r\,.$
Note that the new minimizing sequence $\{\Phi^{*}_{i}\}$ is not necessarily pulled-tight. However, by Proposition 4.4, we can obtain a pulled-tight minimizing sequence $\{\widetilde{\Phi}_{i}\}$ with
$\mathbf{C}(\{\widetilde{\Phi}_{i}\})\subset\mathbf{C}(\{\Phi^{*}_{i}\})\subset\mathbf{B}^{\mathbf{F}}_{\frac{3}{4}r}(\mathcal{W}^{m,\widetilde{\eta}}_{L,\leq\mathfrak{g}_{0}})\,.$
The conclusion follows immediately from a suitable choice of $\eta$.
4.6. Proof of Theorem 2.21
Since $\mathbf{L}(\Lambda_{Z}(\Phi))>\sup_{z\in Z}\mathcal{H}^{2}(\Phi(z))$, for each $\Phi^{\prime}\in\Lambda_{Z}(\Phi)$, one can restrict all the deformations in the previous subsections to occur away from the compact set
$X^{\prime}\mathrel{\mathop{\ordinarycolon}}=\left\{x\in X\mathrel{\mathop{\ordinarycolon}}\mathcal{H}^{2}(\Phi(x))\leq\frac{\mathbf{L}(\Lambda_{Z}(\Phi))+\sup_{z\in Z}\mathcal{H}^{2}(\Phi(z))}{2}\right\}\,.$
For example, all deformations can be composed with a cut-off function on $X$ that vanishes on $X^{\prime}$.
Consequently, all the results follow immediately.
5. The family $\Psi$
5.1. The space of Clifford Tori
Let us recall the parametrization of the space $\mathcal{C}$ of unoriented Clifford tori. A Clifford torus is uniquely determined by an unordered pair of orthogonal unoriented 222-subspaces in $\mathbb{R}^{4}$. Namely, given two orthogonal 222-subspaces $P$ and $Q$, the Clifford torus they define consists of the points on the unit sphere $\mathbb{S}^{3}$ that are equidistant from the equators $P\cap\mathbb{S}^{3}$ and $Q\cap\mathbb{S}^{3}$. Note that for every 222-subspace of $\mathbb{R}^{4}$, there exists exactly one orthogonal 222-subspace $Q$. Hence, to choose an inside direction for a Clifford torus, i.e. to assign an orientation, one simply needs to specify one 222-plane $P$, so the space of oriented Clifford tori is $G_{2}(\mathbb{R}^{4})$, the space of unoriented 2-planes in $\mathbb{R}^{4}$.
By [HO80, §1], the set $G_{2}^{+}(\mathbb{R}^{4})$ of oriented 2-planes in $\mathbb{R}^{4}$ can be identified with the set
[…]
Namely, for an oriented orthonormal basis $(u,v)$ of a 2-plane in $\mathbb{R}^{4}$, we can define
$(z_{1},z_{2},z_{3},z_{4})\mathrel{\mathop{\ordinarycolon}}=u+iv\,.$
Moreover, there is a biholomorphism
$F\mathrel{\mathop{\ordinarycolon}}\mathbb{CP}^{1}\times\mathbb{CP}^{1}\rightarrow Q_{2}\,,$
defined in [HO80, §2]:
$F([w_{1}\mathrel{\mathop{\ordinarycolon}}1],[w_{2}\mathrel{\mathop{\ordinarycolon}}1])$ $=(1+w_{1}w_{2},i(1-w_{1}w_{2}),w_{1}-w_{2},-i(w_{1}+w_{2}))\,,$
$F([w_{1}\mathrel{\mathop{\ordinarycolon}}0],[w_{2}\mathrel{\mathop{\ordinarycolon}}1])$ $=(w_{2},-iw_{2},1,-i)\,,$
$F([w_{1}\mathrel{\mathop{\ordinarycolon}}1],[w_{2}\mathrel{\mathop{\ordinarycolon}}0])$ $=(w_{1},-iw_{1},-1,-i)\,,$
$F([w_{1}\mathrel{\mathop{\ordinarycolon}}0],[w_{2}\mathrel{\mathop{\ordinarycolon}}0])$ $=(1,-i,0,0)\,.$
[…]
$(x,y)\sim(-x,-y)\sim(x,-y)\sim(-x,y)\,,$
which is an ${\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$ (see also [Whi91, §4]).
5.2. Definition of $\Psi$
In this subsection, we define the 9-parameter family $\Psi$. We will reuse many of the notations and presentations in [MN14] by Marques-Neves, [Nur16] by C. Nurser, and an unpublished manuscript [Mar23] by F. C. Marques, in which he constructed a 9-parameter family (in a slightly different way than us) that is an 8-sweepout to prove that the 888-width of the unit 333-sphere is $2\pi^{2}$.
Let $\mathbb{S}^{3}$ be the unit 3-sphere. Let $Z={\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$ parametrize the space of unoriented Clifford tori, and $\widetilde{Z}$ denote the space of oriented Clifford tori. Then $\widetilde{Z}$ is connected, and there is a natural double-cover map $q\mathrel{\mathop{\ordinarycolon}}\widetilde{Z}\to Z$ given by forgetting the orientation. Let $\sigma\mathrel{\mathop{\ordinarycolon}}\widetilde{Z}\to\widetilde{Z}$ be the nontrivial deck transformation. Note that $\sigma(\sigma(z))=z$.
For each $z\in\widetilde{Z}$, let $\Sigma(z)$ be the corresponding oriented Clifford torus, and let $A(z)$ and $A^{*}(z)$ denote the two connected components of $\mathbb{S}^{3}\backslash\Sigma(z)$ such that $A(z)$ and $A^{*}(z)$ vary continuously in $z$. Let $N(z)$ denote the unit normal vector field on $\Sigma(z)$ that points into $A^{*}(z)$. Note that for every $z\in\widetilde{Z}$, and $p\in\Sigma(z)$,
$A(z)=A^{*}(\sigma(z))\,,\quad N(z)_{-p}=-N(z)_{p}\,,\quad N(z)=-N(\sigma(z))\,.$
[…]
Proposition 5.1.
For each oriented Clifford torus $\Sigma(z)$, with $z\in\widetilde{Z}$, there exists a family $\Psi^{\Sigma(z)}_{5}$ of closed subsets of $\mathbb{S}^{3}$, parametrized by ${\mathbb{RP}}^{5}$, such that:
  1. (1)
    The maps $[\Psi^{\Sigma(z)}_{5}]$ and $\Phi^{\Sigma(z)}_{5}$ (defined by C. Nurser) are the same.
  2. (2)
    For every $[(v,t)]\in{\mathbb{RP}}^{5}$ and $z\in\widetilde{Z}$,
    $\Psi^{\Sigma(z)}_{5}([(v,t)])=\Psi^{\Sigma(\sigma(z))}_{5}([(v,-t)])\,.$
We will prove this proposition in the following subsections.
Let us assume this proposition for now, and consider the ${\mathbb{RP}}^{5}\times\widetilde{Z}$-family $\widetilde{\Psi}$ of closed subsets of $\mathbb{S}^{3}$ defined by
$\widetilde{\Psi}([(v,t)],z)\mathrel{\mathop{\ordinarycolon}}=\Psi^{\Sigma(z)}([(v,t)])\,.$
Moreover, consider the $\mathbb{Z}_{2}$-action on ${\mathbb{RP}}^{5}\times\widetilde{Z}$ where the non-trivial element of $\mathbb{Z}_{2}$ acts by
(5.1) $([(v,t)],z)\mapsto([(v,-t)],\sigma(z))\,,$
and let $Y$ be the quotient space. Using Proposition 5.1, the family $\widetilde{\Psi}$ immediately induces a family $\Psi$ of closed subsets of $\mathbb{S}^{3}$, with parameter space $Y$. Furthermore, $Y$ is an ${\mathbb{RP}}^{5}$-bundle over $Z={\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$.
5.2.1. Conformal maps and level surfaces
For each $v\in{\mathbb{B}^{4}}$, we consider the conformal map
$F_{v}\mathrel{\mathop{\ordinarycolon}}\mathbb{S}^{3}\to\mathbb{S}^{3}\,,\quad x\mapsto\frac{(1-|v|^{2})}{|x-v|^{2}}(x-v)-v\,.$
For each $(v,z)\in{\mathbb{B}^{4}}\times\widetilde{Z}$, we define $d(v,z)\mathrel{\mathop{\ordinarycolon}}\mathbb{S}^{3}\to\mathbb{R}$ to be the signed distance to $F_{v}(\Sigma(z))$, which is positive in $F_{v}(A^{*}(z))$ and negative in $F_{v}(A(z))$. Then for each $v\in{\mathbb{B}^{4}}$ and $t\in\mathbb{R}$, we define the open set
$A(v,t,z)=\{x\in\mathbb{S}^{3}\mathrel{\mathop{\ordinarycolon}}d(v,z)(x)<t\}\,,$
and the closed set
$\Sigma(v,t,z)=\partial A(v,t,z)\,.$
Note that $\Sigma(v,t,z)$ has a natural orientation and induces an element in $\mathcal{Z}(\mathbb{S}^{3};\mathbb{Z})$, and
$A(v,t,\sigma(z))=\mathbb{S}^{3}\setminus(\Sigma(z)\cup A(v,-t,z))\,.$
In order to define an ${\mathbb{RP}}^{5}$-family, we need to reparametrize $A$ by performing blow-up suitably, because $A$ does not give a continuous family near $v\in\Sigma(z)$.
5.2.2. Reparametrizing $A$ via blow-up
Define the right-half disk on $\mathbb{R}^{2}$ as
$D^{2}_{+}(r)\mathrel{\mathop{\ordinarycolon}}=\{s=(s_{1},s_{2})\in\mathbb{R}^{2}\mathrel{\mathop{\ordinarycolon}}|s|<r,s_{1}\geq 0\}\,.$
For $\varepsilon>0$ sufficiently small, we define the map $\Lambda(z)\mathrel{\mathop{\ordinarycolon}}\Sigma(z)\times D^{2}_{+}(3\varepsilon)\to\overline{{\mathbb{B}^{4}}}$ by
$\Lambda(z)(p,s)=(1-s_{1})(\cos(s_{2})p+\sin(s_{2})N(z)_{p})\,,$
which is a diffeomorphism onto a neighborhood of $\Sigma(z)$ in $\overline{{\mathbb{B}^{4}}}$. Notice $\Lambda(z)$ maps into $\mathbb{S}^{3}$ if $s_{1}=0$.
Let $\Omega_{r}(z)$ be the image $\Lambda(z)(\Sigma(z)\times D^{2}_{+}(r))$ for all $r\leq 3\varepsilon$. Then as $s$ approaches $(0,3\varepsilon)$ (resp. $(0,-3\varepsilon)$), $\Lambda(z)(p,s)$ approaches $A^{*}(z)\backslash\Omega_{3\varepsilon}(z)$ (resp. $A(z)\backslash\Omega_{3\varepsilon}(z)$).
Now, we define a continuous map $T(z)\mathrel{\mathop{\ordinarycolon}}\overline{{\mathbb{B}^{4}}}\to\overline{{\mathbb{B}^{4}}}$ which collapses the tubular neighborhood $\Omega_{\varepsilon}(z)$ of $\Sigma(z)$ onto $\Sigma(z)$:
  • $T(z)$ is the identity on $\overline{{\mathbb{B}^{4}}}\setminus\Omega_{3\varepsilon}(z)$.
  • On $\Omega_{3\varepsilon}(z)$,
    $T(z)(\Lambda(z)(p,s))=\Lambda(z)(p,\psi(|s|)s)\,,$
    where $\psi$ is smooth, 0 on $[0,\varepsilon]$, strictly increasing on $[\varepsilon,2\varepsilon]$, and 1 on $[2\varepsilon,3\varepsilon]$.
Note that $T(z)=T(\sigma(z))$.
For every $p\in\Sigma(z)$ and $k\in[-\infty,+\infty]$, define
(5.2) $\overline{Q}_{p,k,z}=-\frac{k}{\sqrt{1+k^{2}}}p-\frac{1}{\sqrt{1+k^{2}}}N(z)_{p}\in\mathbb{S}^{3}\,.$
(Note we use the convention that $\frac{k}{\sqrt{1+k^{2}}}=\pm 1$ and $\frac{1}{\sqrt{1+k^{2}}}=0$ for $k=\pm\infty$.) We define the generalized Gauss map $\overline{Q}(z)\mathrel{\mathop{\ordinarycolon}}\mathbb{S}^{3}\cup\overline{\Omega}_{\varepsilon}(z)\to\mathbb{S}^{3}$ by:
(5.3) $\overline{Q}(z)(v)=\begin{cases}-T(v)&v\in A^{*}(z)\setminus\overline{\Omega}_{\varepsilon}(z)\\ T(v)&v\in A(z)\setminus\overline{\Omega}_{\varepsilon}(z)\\ \overline{Q}_{p,k(s),z}&v=\Lambda(z)(p,s)\in\overline{\Omega}_{\varepsilon}(z)\,,\end{cases}$
where $s=(s_{1},s_{2})\in D^{2}_{+}(\varepsilon)$ and
(5.4) $k(s)=\frac{s_{2}}{\sqrt{\varepsilon^{2}-s_{2}^{2}}}\in[-\infty,+\infty]\,.$
The key property of $\overline{Q}$ is that, as $v\in\mathbb{S}^{3}$ moves from $A(z)\backslash{\Omega}_{3\varepsilon}(z)$, crosses $\Sigma(z)$, and arrives in $A^{*}(z)\backslash{\Omega}_{3\varepsilon}(z)$ (so that $v\in\mathbb{S}^{3}$ crosses from one side of $\Sigma(z)$ into another), $\overline{Q}(z)$ flips from being the identity map to the antipodal map, continuously. And the continuous function $k(s)$ is such that $k(s)$ goes to +$\infty$+ ∞ (resp. $\infty$- ∞) if $\Lambda(z)(p,s)$ goes towards $A^{*}(z)\backslash\Omega_{\varepsilon}(z)$ (resp. $A(z)\backslash\Omega_{\varepsilon}(z)$).
We define $\overline{r}(z)\mathrel{\mathop{\ordinarycolon}}\mathbb{S}^{3}\cup\overline{\Omega}_{\varepsilon}(z)\to[0,\pi]$ by:
(5.5) $\overline{r}(z)(v)=\begin{cases}0&v\in A^{*}(z)\setminus\overline{\Omega}_{\varepsilon}(z)\\ \pi&v\in A(z)\setminus\overline{\Omega}_{\varepsilon}(z)\\ \overline{r}_{k(s)}&v=\Lambda(z)(p,s)\in\overline{\Omega}_{\varepsilon}(z)\,,\end{cases}$
where $\bar{r}_{k}=\pi/2-\arctan k$. The key point is that $\overline{r}(z)$ changes from 0 to $\pi$ continuously across the intermediate region $\overline{\Omega}_{\varepsilon}(z)$.
With the above preparation, we can start defining $\Psi^{\Sigma(z)}_{5}$.
We know the map $\overline{Q}(z)$ satisfies $\overline{Q}(\sigma(z))=-\overline{Q}(z)$ from $N(z)=-N(\sigma(z))$, and the function $\overline{r}(z)$ satisfies $r(\sigma(z))=\pi-\overline{r}(z)$. Notice that the subset $\overline{\Omega}_{\varepsilon}(z)$ is invariant under the antipodal map, and $\overline{\Omega}_{\varepsilon}(\sigma(z))=\overline{\Omega}_{\varepsilon}(z)$. The function $\overline{r}$ can initially be considered as a continuous function defined on
$\{(z,v)\mathrel{\mathop{\ordinarycolon}}z\in\widetilde{Z},v\in\mathbb{S}^{3}\cup\overline{\Omega}_{\varepsilon}(z)\}\,,$
which is a closed subset of $\widetilde{Z}\times{\mathbb{B}^{4}}$. We extend it continuously to a function
$\overline{r}\mathrel{\mathop{\ordinarycolon}}\widetilde{Z}\times\overline{\mathbb{B}^{4}}\to[0,\pi]$
that still satisfies $\overline{r}(\sigma(z))=\pi-\overline{r}(z)$. For each $(v,t,z)\in\overline{\mathbb{B}^{4}}\times\mathbb{R}\times\widetilde{Z}$, we define the set
$U(v,t,z)=\begin{cases}A(T(z)(v),t,z)&\text{if }v\in{\mathbb{B}^{4}}\setminus\overline{\Omega}_{\varepsilon}(z)\,,\\ B_{\overline{r}(z)(v)+t}(\overline{Q}(z)(v))&\text{if }v\in\mathbb{S}^{3}\cup\overline{\Omega}_{\varepsilon}(z)\,.\end{cases}$
Notice that $U(v,\pi,z)=\mathbb{S}^{3}$ and $U(v,-\pi,z)$ is empty for every $v\in{\mathbb{B}^{4}}$. And for $v\in\mathbb{S}^{3}\cup\overline{\Omega}_{\varepsilon}(z)$, $U$ gives geodesic spheres of varying radius, and the radius always covers the full range $[0,\pi]$ as $t$ ranges over $[-\pi,\pi]$.
Marques-Neves [MN14, Theorem 5.1] showed that the family $\partial U$ is continuous in the flat topology.
5.2.3. Closing up the parameter space
Let $\gamma\mathrel{\mathop{\ordinarycolon}}[0,1]\to[0,1]$ be the continuous function satisfying $\gamma(t)=0$ if $t\leq\frac{1}{2}$ and $\gamma(t)=2t-1$ if $t\geq\frac{1}{2}$. For each
$(v,t,z)\in\overline{\mathbb{B}^{4}}\times[-1,1]\times\widetilde{Z}$
we define the open set
$U^{\prime}(v,t,z)=U\left(v,2\pi t+\gamma(|v|)\left(\frac{\pi}{2}-\overline{r}(z)(v)\right),z\right)\,.$
Reparametrizing $t$ as above allows us to close up the boundary of $\overline{\mathbb{B}^{4}}\times[-1,1]\times\{z\}$ via the antipodal map, as follows.
For
$(v,t)\in\partial(\overline{\mathbb{B}^{4}}\times[-1,1])=(\mathbb{S}^{3}\times(-1,1))\cup(\overline{\mathbb{B}^{4}}\times\{-1,1\})\,,$
it is straightforward to check that
$U^{\prime}(v,t,z)=\begin{cases}B_{\frac{\pi}{2}+2\pi t}(\overline{Q}(z)(v))&\text{if }(v,t)\in\mathbb{S}^{3}\times(-1,1)\,,\\ 0&\text{if }(v,t)\in\overline{\mathbb{B}^{4}}\times\{-1\}\,,\\ \mathbb{S}^{3}&\text{if }(v,t)\in\overline{\mathbb{B}^{4}}\times\{1\}\,.\end{cases}$
Note that the smallness of $\varepsilon$ is used here for the case where $v\in\overline{\Omega}_{\varepsilon}(z)$. Now, using the fact that $B_{\pi-r}(-p)=\mathbb{S}^{3}\backslash\overline{B_{r}(p)}$, $T(z)(-v)=-T(z)(v)$, and $\overline{Q}(z)(-v)=-\overline{Q}(z)(v)$ by [Nur16, §3.4.3], it follows immediately that for every $(v,t)\in\partial(\overline{\mathbb{B}^{4}}\times[-\pi,\pi])$,
(5.6) $U^{\prime}(-v,-t,z)=\mathbb{S}^{3}\setminus\overline{U^{\prime}(v,t,z)}\,,$
This allows us to define an ${\mathbb{RP}}^{5}$-family $\Psi^{\Sigma(z)}_{5}$ of closed subsets by
$\Psi^{\Sigma(z)}_{5}([(v,t)])=\partial U^{\prime}(v,t,z)\,.$
Note that we have disregarded orientation, and a geodesic sphere of radius 0 or $\pi$ should be both treated as a single point, ensuring $\Psi_{5}^{\Sigma(z)}$ satisfies Definition 2.4 (2).
5.2.4. Symmetry of $\Psi^{\Sigma(z)}_{5}$ in $z$
Furthermore, it follows from a straightforward computation that the map $U^{\prime}$ satisfies the identity
$U^{\prime}(v,t,z)=\mathbb{S}^{3}\setminus\overline{U^{\prime}(v,-t,\sigma(z))}\,,$
for every $(v,t,z)\in{\mathbb{B}^{4}}\times[-1,1]\times\widetilde{Z}$. Thus, we have
$\Psi^{\Sigma(z)}_{5}([(v,t)])=\Psi^{\Sigma(\sigma(z))}_{5}([(v,-t)])\,.$
This finishes the proof of Proposition 5.1.
5.3. Proof of Theorem 3.1
From the above construction of $\Psi$, it is clear that for each oriented Clifford torus $\Sigma$, there exists a homeomorphism $f\mathrel{\mathop{\ordinarycolon}}Y_{[\Sigma]}\to{\mathbb{RP}}^{5}$ such that $\Phi^{\Sigma}_{5}\circ f=[\Psi|_{Y_{[\Sigma]}}]$. So to prove Theorem 3.1, it suffices to show that $\Psi$ is a Simon-Smith family of genus 1$\leq$ 1≤ 1. We start with studying the topology of the surfaces in $\Psi$.
5.3.1. Topology of $\Sigma(v,t,z)$
Fix a $z_{0}\in\widetilde{Z}$ which corresponds to the Clifford torus $\Sigma_{0}$ given by
$X(\alpha,\beta)=\frac{1}{\sqrt{2}}(\cos\alpha,\sin\alpha,\cos\beta,\sin\beta)\,,$
with $\alpha,\beta\in S^{1}$ (where $S^{1}$ is viewed as $[0,2\pi]$ with endpoints identified), oriented such that the distance function $d(0,z_{0})$ is negative on the “core circle”
$w(\beta)=\frac{1}{\sqrt{2}}(0,0,\cos\beta,\sin\beta),\quad\beta\in S^{1}\,.$
Proposition 5.2.
Let $(v,t)\in{\mathbb{B}^{4}}\times[-\pi,\pi]$. We have the following descriptions regarding the closed set
$\Sigma(v,t)\mathrel{\mathop{\ordinarycolon}}=\partial\{x\in\mathbb{S}^{3}\mathrel{\mathop{\ordinarycolon}}d(v,z_{0})(x)<t\}\,.$
If $v=0$, then $\Sigma(v,t)$ is a smooth torus for $t\in(-\pi/4,\pi/4)$, a great circle for $t=\pm\pi/4$, and empty for other $t$.
If $v\neq 0$ and $v/|v|$ lies on the circle
(5.7) $\frac{1}{\sqrt{2}}(\cos\alpha,\sin\alpha,0,0)\,,\quad\alpha\in S^{1}\,,$
then there exist $-\pi<t_{1}<0<t_{2}<\pi,$ depending on $v$, such that:
  1. (1)
    For $t_{1}<t<t_{2}$, $\Sigma(v,t)$ is a smooth torus.
  2. (2)
    For $t=t_{1}$ and $t=t_{2}$, $\Sigma(v,t)$ is a great circle.
  3. (3)
    For $t<t_{1}$ and $t>t_{2}$, $\Sigma(v,t)$ is empty.
If $v$ does not fall into the above two cases, then there exist
$-\pi<t_{1}<t_{2}<0<t_{3}<t_{4}<\pi\,,$
depending on $v$, such that:
  1. (1)
    For $t_{2}<t<t_{3}$, $\Sigma(v,t)$ is a smooth torus.
  2. (2)
    For $t=t_{2}$ and $t=t_{3}$, $\Sigma(v,t)$ is topologically a sphere with two points identified such that it is smooth except at a point.
  3. (3)
    For $t_{1}<t<t_{2}$ and $t_{3}<t<t_{4}$, $\Sigma(v,t)$ is topologically a sphere, smooth except at two points.
  4. (4)
    For $t=t_{1}$ and $t=t_{4}$, $\Sigma(v,t)$ is a point.
  5. (5)
    For $t<t_{1}$ and $t>t_{4}$, $\Sigma(v,t)$ is empty.
Proof.
First, the case $v=0$ is clear.
Second, in the case where $v\neq 0$ and $v/|v|$ lies on the circle (5.7), note that both the Clifford torus $\Sigma_{0}$ and the map $F_{v}$, are symmetric about the antipodal points ±$\pm v/|v|$. Therefore, the image $F_{v}(\Sigma_{0})$, as well as the level surfaces is parallel to $F_{v}(\Sigma_{0})$, share this symmetry, and the desired result follows immediately.
Finally, consider the case where $v\neq 0$ with $v/|v|$ not lying on the circle (5.7). The Clifford torus $\Sigma_{0}$ is the envelope of the 2-spheres $S_{\beta}\subset\mathbb{S}^{3}$ with center $w(\beta)$ and radius $\pi/4$ under the standard metric on $\mathbb{S}^{3}$, where $\beta\in S^{1}$. Thus, $\Sigma_{0}$ is a channel surface. The sphere $S_{\beta}$ is given by the following equations on $x\in\mathbb{R}^{4}$:
$|x|=1\,,\quad x\cdot w(\beta)=\cos(\pi/4)\,.$
Now, when we apply the conformal diffeomorphism $F_{v}$ to $S_{\beta}$, we obtain another 222-sphere $F_{v}(S_{\beta})$ in $\mathbb{S}^{3}$, with a new center $\widetilde{w}(\beta)$ and a new radius $\widetilde{r}(\beta)\in(0,\pi)$ under the standard metric on $\mathbb{S}^{3}$.
We see that the function $\widetilde{r}\mathrel{\mathop{\ordinarycolon}}S^{1}\to\mathbb{R}$ has the following property.
Claim 5.3.
If $v\in{\mathbb{B}^{4}}$ is non-zero and $v/|v|$ does not lie on the circle (5.7), then the function $\widetilde{r}\mathrel{\mathop{\ordinarycolon}}S^{1}\to\mathbb{R}$ has exactly one local maximum and one local minimum.
We will postpone the proof to the end, but first, let us note that this lemma is sufficient to justify the final case of our proposition.
Without loss of generality, we can focus on the case $t<0$, as the case $t>0$ follows from an identical analysis on the Clifford torus $-\Sigma_{0}$, i.e. the Clifford torus with the opposite orientation to $\Sigma_{0}$.
Let us consider the closed envelop given by the collection of spheres
$\{\partial B_{\widetilde{r}(\beta)+t}(\widetilde{w}(\beta))\}_{\beta\in S^{1}}\,,$
where we use the convention that $\partial B_{0}(p)=\{p\}$ and $\partial B_{r}=\emptyset$ for $r<0$. Given that conformal diffeomorphisms preserve channel surfaces, it is straightforward to verify that this envelop is exactly the set $\Sigma(v,t)$.
[…]
Proof of Claim 5.3.
Recalling that $F_{v}$ has a symmetry about the points ±$\pm v/|v|$, and $F_{v}$ pushes everything away from $\frac{v}{|v|}$ and towards $-\frac{v}{|v|}$, we know the center $\widetilde{w}(\beta)$ of the sphere $F_{v}(S_{\beta})$ takes the form
$\widetilde{w}(\beta)=\frac{-\cos(\theta(\beta))\frac{v}{|v|}+\sin(\theta(\beta))w(\beta)}{\left|-\cos(\theta(\beta))\frac{v}{|v|}+\sin(\theta(\beta))w(\beta)\right|}$
for some function $\theta(\beta)\in(0,\pi/2)$. For convenience, let us denote the above denominator by $l(\beta)$. Additionally, the new sphere $F_{v}(S_{\beta})$ is characterized by the following equations for $y\in\mathbb{R}^{4}$:
$|y|=1\quad\textrm{ and }\quad y\cdot\widetilde{w}(\beta)=\textrm{constant}\,,$
where the constant equals to $\cos$ of the radius of the new sphere.
Thus, let us compute $F_{v}(x)\cdot\widetilde{w}(\beta)$ for $x\in S_{\beta}$. For simplicity, we omit writing the dependence of $l,\theta$ and $w$ on $\beta$.
$\quad F_{v}(x)\cdot\widetilde{w}(\beta)$
$=\left(\frac{1-|v|^{2}}{|x-v|^{2}}(x-v)-v\right)\cdot\frac{-\cos\theta\frac{v}{|v|}+\sin\theta\;w}{l}$
(5.8) $=\frac{1}{l}\frac{1-|v|^{2}}{1+|v|^{2}-2x\cdot v}\left(-\cos\theta\frac{x\cdot v}{|v|}+\cos\theta\;|v|+\sin\theta\frac{1}{\sqrt{2}}-\sin\theta\;w\cdot v\right)$
$\;\;+\frac{1}{l}(\cos\theta\;|v|-\sin\theta\;w\cdot v)\,.$
Note that in the last equality, we use $w\cdot x=1/\sqrt{2}$, which holds because $x\in S_{\beta}$. Hence, for some constant $C=C(v,\beta)$, independent of $x\in S_{\beta}$, we have
$-\cos\theta\frac{x\cdot v}{|v|}+\cos\theta\;|v|+\sin\theta\frac{1}{\sqrt{2}}-\sin\theta\;w\cdot v=C\left({1+|v|^{2}-2x\cdot v}\right)\,.$
Now, by setting both the zeroth-order term and the first-order term in $x$ to zero, and canceling $C$ from these two equations, we obtain
(5.9) $\frac{1-|v|^{2}}{2|v|}\cos\theta=\left(\frac{1}{\sqrt{2}}-w\cdot v\right)\sin\theta\,,$
which leads to
(5.10) $\cot\theta=\frac{2|v|}{1-|v|^{2}}\left(\frac{1}{\sqrt{2}}-w\cdot v\right)\,.$
Hence, substituting (5.9) into (5.8), the radius $\widetilde{r}(\beta)\in(0,\pi)$ satisfies:
$\cos\widetilde{r}(\beta)$ $=F_{v}(x)\cdot\widetilde{w}(\beta)$
$=\frac{1}{l}\frac{1-|v|^{2}}{|x-v|^{2}}(x-v)\cos\theta\left(-\frac{x\cdot v}{|v|}+|v|+\frac{1-|v|^{2}}{2|v|}\right)$
$\;\;+\frac{1}{l}(\cos\theta\;|v|-\sin\theta\;w\cdot v)$
$=\frac{1}{l}\left(\frac{1-|v|^{2}}{2|v|}\cos\theta+(\cos\theta\;|v|-\sin\theta\;w\cdot v)\right)$
$=\frac{1}{l}\left(\frac{1}{|v|}\cos\theta-\frac{1}{\sqrt{2}}\sin\theta\right)\,.$
Now, by a straightforward computation of $l$ using (5.9), we find that
$l=\sqrt{\frac{1}{2}\sin^{2}\theta+\left(\frac{1}{|v|}\cos\theta-\frac{1}{\sqrt{2}}\sin\theta\right)^{2}}\,.$
It follows straightforward that
$\cos^{2}\widetilde{r}=\left(1+\frac{1}{(\frac{\sqrt{2}}{|v|}\cot\theta-1)^{2}}\right)\,.$
Since $\widetilde{r}$ depends smoothly on $\theta$, the equation can be extended to the case where the denominator is zero. Together with (5.10) one can derive that
(5.11) $\cot\widetilde{r}(\beta)=\frac{\sqrt{2}}{|v|}\cot\theta-1=\frac{2(1-\sqrt{2}w\cdot v)}{1-|v|^{2}}-1\,.$
Note that the sign of the right-hand side is positive sign due to the monotonicity of $\cot$ and the geometric behavior of the conformal map $F_{v}$: when $w\cdot v$ is larger, $\widetilde{r}$ is larger.
Since $\cot$ is a strictly decreasing function with non-zero slope on $(0,\pi)$, by (5.11) and the fact that $\widetilde{r}$ depends smoothly on $w\cdot v$, it suffices to show that $w\cdot v(\beta)$, which is a function of $\beta\in S^{1}$, has exactly one local minimum and one local maximum. This holds true because $w(\beta)$ describe a circles in $\mathbb{S}^{3}$ that is not perpendicular to $v$. This completes the proof of the claim. ∎
This finishes the proof of Proposition 5.2. ∎
5.3.2. $\Psi$ is a Simon-Smith family of genus 1$\leq$ 1≤ 1
Let us finish the proof of Theorem 3.1, by showing that $\Psi$ is a Simon-Smith family of genus 1$\leq$ 1≤ 1.
By the definition of $\Psi$ in the previous section and Proposition 5.2, we can conclude that $\Psi$ consists of the following types of closed subsets:
  1. (1)
    A smooth torus: This arises as the nearby level surface of the signed distance function to $F_{v}(\Sigma(z))$, where $v\notin\Sigma(z)$.
  2. (2)
    A smooth sphere: They are geodesic spheres on $\mathbb{S}^{3}$ with radius strictly between 0 and $\pi$.
  3. (3)
    A great circle: This occurs in the second case of $v$ in Proposition 5.2.
  4. (4)
    A topological sphere with two points identified, so that it is smooth except at a point: This occurs in the third case of $v$ in Proposition 5.2.
  5. (5)
    A topological sphere, smooth except at two points: This occurs in the third case of $v$ in Proposition 5.2.
  6. (6)
    A single point: It has several forms. (1) A geodesic sphere on $\mathbb{S}^{3}$ with radius 0. (2) A geodesic sphere on $\mathbb{S}^{3}$ with radius $\pi$. (3) It also occurs in the third case of $v$ in Proposition 5.2.
  7. (7)
    The empty set.
Then, it follows easily that each member of $\Psi$ is an element of $\mathcal{S}(S^{3})$, and $\Psi$ satisfies Definition 2.4, (1), (2), (3), and also the genus bound (5).
As for Definition 2.4 (4), regarding the local $C^{\infty}$-convergence, we know from the definition of $\Psi$ that it suffices to fix a $z=z_{0}$ and show that $\Psi^{\Sigma(z_{0})}_{5}$ satisfies Definition 2.4 (4). In fact, if we define the family
$\Psi^{\prime}_{5}(v,t)\mathrel{\mathop{\ordinarycolon}}=\partial U(v,t,z_{0}),\;\;(v,t)\in\overline{\mathbb{B}^{4}}\times[-2\pi,2\pi]$
(with $\Psi^{\prime}_{5}(v,t)$ treated as a single point if it is a geodesic sphere of radius 0 or $\pi$), and notice that $\Psi^{\Sigma(z_{0})}_{5}$ was defined as a subfamily of $\Psi^{\prime}_{5}$ with suitable closing-up of the boundary of the parameter space, it becomes clear that it suffices to show that $\Psi^{\prime}_{5}$ satisfies Definition 2.4 (4).
Let us first show that $\Psi^{\prime}_{5}(\cdot,0)$ satisfies Definition 2.4 (4). We now explicitly describe, for each $\Psi^{\prime}_{5}(v,0)$, what the associated set $P(v,0)$ of points (according to Definition 2.4) should be. By definition, $\Psi^{\prime}_{5}(v,0)$ takes on one of the following forms, for $v\in\overline{B}^{4}$.
  1. (1)
    If $v\in{\mathbb{B}^{4}}\backslash\overline{\Omega_{\varepsilon}(z_{0})}$, $\Psi^{\prime}_{5}(v,0)$ is a smooth torus.
  2. (2)
    If $v\in\mathbb{S}^{3}\backslash\overline{\Omega_{\varepsilon}(z_{0})}$, $\Psi^{\prime}_{5}(v,0)$ is empty.
  3. (3)
    If $v\in\mathbb{S}^{3}\cap\overline{\Omega_{\varepsilon}(z_{0})}$, $\Psi^{\prime}_{5}(v,0)$ is a single point.
  4. (4)
    If $v\in\overline{\Omega_{\varepsilon}(z_{0})}\backslash\mathbb{S}^{3}$, $\Psi^{\prime}_{5}(v,0)$ is a smooth geodesic sphere.
[…]
Now, consider the case $t\neq 0$. By the definition of $U$, each $\Psi^{\prime}_{5}(v,t)$ is a level surface of the signed distance function to $\Psi^{\prime}_{5}(v,0)$. Thus, using Proposition 5.2 and the fact that $\Psi^{\prime}_{5}(\cdot,0)$ satisfies Definition 2.4 (4), it is easy to check that $\Psi^{\prime}_{5}$ also satisfies Definition 2.4 (4), with $P(v,t)$ be the union of projection of $P(v,0)$ to $\Psi^{\prime}_{5}(v,t)$ and finite singular points as described in Proposition 5.2. This finishes the proof that $\Psi^{\prime}_{5}$ satisfies Definition 2.4 (4).
Finally, to show that the closing-up $\Psi^{\Sigma(z_{0})}_{5}$ satisfies Definition 2.4 (4), one simply need to merge the associated set $P(v,t)$ and $P(v^{\prime},t^{\prime})$ if $(v,t)\sim(v^{\prime},t^{\prime})$. Since $\Psi$ is induced from $\Psi^{\Sigma(z_{0})}_{5}$ via $SO(4)$-action and another closing-up, it follows that $\Psi$ also satisfies Definition 2.4 (4).
As a result, $\Psi$ is a Simon-Smith family of genus 1$\leq$ 1≤ 1.
5.4. A 7-sweepout in $\Psi$
As part of the setup for the proof of Theorem 3.13, we consider a subset $C\subset Y=\mathrm{dmn}(\Psi)$, where $C\cong{\mathbb{RP}}^{4}\times{\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$, such that $[\Psi|_{C}]$ is a 7-sweepout.
First, we define the submanifold
$\widetilde{C}\mathrel{\mathop{\ordinarycolon}}=\{([(v,t)],z)\in{\mathbb{RP}}^{5}\times\widetilde{Z}\mathrel{\mathop{\ordinarycolon}}t=0\}\,.$
[…]
Now,
$H_{1}(C;\mathbb{Z}_{2})\cong H_{1}({\mathbb{RP}}^{4};\mathbb{Z}_{2})\oplus H_{1}({\mathbb{RP}}^{2};\mathbb{Z}_{2})\oplus H_{1}({\mathbb{RP}}^{2};\mathbb{Z}_{2})\cong\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}\,.$
Let $\omega_{C},\alpha_{C},\beta_{C}\in H^{1}(C;\mathbb{Z}_{2})$ be respectively the Hom-duals of
(5.12) $(1,0,0),(0,1,0),(0,0,1)\in H_{1}(C;\mathbb{Z}_{2}).$
By the Künneth formula and the fact that the cohomology ring $H^{*}({\mathbb{RP}}^{k};\mathbb{Z}_{2})$ is $\mathbb{Z}_{2}[x]/(x^{k+1})$, we have
$H^{*}(C;\mathbb{Z}_{2})\cong\mathbb{Z}_{2}[\omega_{C},\alpha_{C},\beta_{C}]/(\omega_{C}^{5},\alpha_{C}^{3},\beta_{C}^{3})\,.$
As verified in [Nur16, §3.4.3], the three homology classes in (5.12) all correspond to 111-sweepouts under $[\Psi|_{C}]$, and thus the element
$\lambda_{C}\mathrel{\mathop{\ordinarycolon}}=([\Psi|_{C}])^{*}(\bar{\lambda})\in H^{1}(C;\mathbb{Z}_{2})$
is equal to $\omega_{C}+\alpha_{C}+\beta_{C}$. Hence, by the Künneth formula,
$(\lambda_{C})^{7}=\omega_{C}^{4}\cup\alpha_{C}^{2}\cup\beta_{C}+\omega_{C}^{4}\cup\alpha_{C}\cup\beta_{C}^{2}\neq 0$
in $\mathbb{Z}_{2}$-coefficients. So $[\Psi|_{C}]$ is a 7-sweepout, as $[\Psi]$ is continuous in the $\mathcal{F}$-norm by Proposition 2.8.
5.5. Proof of Theorem 3.13
Recall the three elements $\lambda,\alpha,\beta$ in $H^{1}(Y;\mathbb{Z}_{2})$ defined by:
  • Let $\bar{\lambda}$ be the non-trivial element of $H^{1}(\mathcal{Z}_{2}(S^{3};\mathbb{Z}_{2});\mathbb{Z}_{2})$; define $\lambda=[\Psi]^{*}(\bar{\lambda})$.
  • Let $a={\mathbb{RP}}^{1}\times{\mathbb{RP}}^{2}\subset{\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$, and $A\subset Y$ be the ${\mathbb{RP}}^{5}$-subbundle over $a$; define $\alpha$ to be the Poincaré dual $PD(A)$ of $A$.
  • Let $b={\mathbb{RP}}^{2}\times{\mathbb{RP}}^{1}\subset{\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$, and $B\subset Y$ be the ${\mathbb{RP}}^{5}$-subbundle over $b$; define $\beta$ to be the Poincaré dual $PD(B)$ of $B$.
Our goal is to prove $\lambda^{5}\cup\alpha^{2}\cup\beta^{2}\neq 0$.
Let us first write down the generators of $H^{*}(Y;\mathbb{Z}_{2})$. Let $C\subset Y$ be the 8-dimensional subset obtained in the previous subsection such that $[\Psi|_{C}]$ is a 777-sweepout, and $i_{C}\mathrel{\mathop{\ordinarycolon}}C\to Y$ be the inclusion map. Recall that the element
$\lambda_{C}\mathrel{\mathop{\ordinarycolon}}=i_{C}^{*}(\lambda)=([\Psi]\circ i_{C})^{*}(\bar{\lambda})\in H^{1}(C;\mathbb{Z}_{2})$
satisfies
(5.13) $(\lambda_{C})^{7}\neq 0\;\;{\textrm{in}}\;H^{*}(C;\mathbb{Z}_{2})\,.$
Now, set $\gamma\mathrel{\mathop{\ordinarycolon}}=PD(C)\in H^{1}(Y;\mathbb{Z}_{2})$.
Recall that $Y$ is an ${\mathbb{RP}}^{5}$-bundle over $Z={\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$:
${\mathbb{RP}}^{5}\to Y=\mathrm{dmn}(\Psi)\to Z\,.$
Let us denote the projection map by $\pi\mathrel{\mathop{\ordinarycolon}}Y\to Z$.
Claim 5.4.
The cohomology ring $H^{*}(Y;\mathbb{Z}_{2})$ is generated by $\alpha,\beta,\gamma.$
Proof.
First, by definition, $\alpha,\beta$ are the pullbacks of the two generators of $H^{*}(Z;\mathbb{Z}_{2})$ under the projection map $\pi\mathrel{\mathop{\ordinarycolon}}Y\to Z$. Moreover, note that for each $z\in Z$, the intersection of $C$ with every fiber $Y_{z}\cong{\mathbb{RP}}^{5}$ of $Y$ is an ${\mathbb{RP}}^{4}$, whose Poincaré dual in $H^{1}(Y_{z};\mathbb{Z}_{2})$ generates
$H^{*}(Y_{z};\mathbb{Z}_{2})\cong H^{*}({\mathbb{RP}}^{5};\mathbb{Z}_{2})\,.$
Thus, if we pullback $\gamma$ to $H^{1}(Y_{z};\mathbb{Z}_{2})$ via the inclusion map $Y_{z}\hookrightarrow Y$, we obtain the generator of $H^{*}(Y_{z};\mathbb{Z}_{2})$. Then the claim follows immediately from the Leray–Hirsch Theorem (see [Hat02, Theorem 4D.1]). ∎
[…]
Claim 5.5.
$\alpha^{2}\cup\beta^{2}=PD(Y_{(p,p)})$, and $\alpha^{r}\cup\beta^{s}=0$ whenever $r\geq 3$ or $s\geq 3$.
Proof.
By definition, $\alpha$ is the Poincaré dual of $A_{i}\mathrel{\mathop{\ordinarycolon}}=(\pi_{1}\circ\pi)^{-1}(k_{i})$ while $\beta$ is the Poincaré dual of $B_{i}\mathrel{\mathop{\ordinarycolon}}=(\pi_{2}\circ\pi)^{-1}(k_{i})$, for $i=1,2$. Therefore, $\alpha^{2}\cup\beta^{2}$ is the Poincaré dual of $A_{1}\cap A_{2}\cap B_{1}\cap B_{2}$, which is the fiber $Y_{(p,p)}$.
If $r\geq 3$ or $s\geq 3$, we just consider three perturbed copies of $k_{1}\subset{\mathbb{RP}}^{2}$ that do not share a common point, and their preimages under $\pi_{1}\circ\pi$ or $\pi_{2}\circ\pi$, similarly as above. Then second claim would follow. ∎
Claim 5.6.
In $H^{1}(Y;\mathbb{Z}_{2})$, $\lambda$ is not a linear combination of just $\alpha$ and $\beta$.
Proof.
Suppose by contradiction that $\lambda=c_{1}\alpha+c_{2}\beta$ for some $c_{1},c_{2}\in\mathbb{Z}_{2}$. Then
$\lambda_{C}=i_{C}^{*}(\lambda)=i_{C}^{*}(c_{1}\alpha+c_{2}\beta)\,.$
Recall that $(\lambda_{C})^{7}\neq 0$ by (5.13). But by Claim 5.5, $(c_{1}\alpha+c_{2}\beta)^{7}=0$, which leads to a contradiction. ∎
From Claim 5.4 and 5.6, we have $\lambda=\gamma+c_{1}\alpha+c_{2}\beta$ for some $c_{1},c_{2}\in\mathbb{Z}_{2}$. Theorem 3.13 will follow from the following claim.
Claim 5.7.
$\gamma^{5}\cup\alpha^{2}\cup\beta^{2}\neq 0$.
Proof.
Since $C\to Z$ is a trivial $\mathbb{RP}^{4}$-bundle, we can perturb it to obtain $C_{1},C_{2},C_{3},C_{4},C_{5}$ such that
[…]
  • All $\hat{C}_{i}$’s transversely intersect at a single point.
Thus, $Y_{(p,p)}\cap\bigcap_{i=1}^{5}C_{i}$ is just a single point in $Y$. Together with Claim 5.5 and the definition $\gamma=PD(C)$, we know $\gamma^{5}\cup\alpha^{2}\cup\beta^{2}\neq 0$. ∎
Thus, by the above claim and Claim 5.5, we have
$\lambda^{5}\cup\alpha^{2}\cup\beta^{2}=(\gamma+(c_{1}\alpha+c_{2}\beta))^{5}\cup\alpha^{2}\cup\beta^{2}=\gamma^{5}\cup\alpha^{2}\cup\beta^{2}\neq 0$
in $H^{*}(Y;\mathbb{Z}_{2})$. This completes the proof of Theorem 3.13.
6. Perturbing the metric
Note that in a 333-dimensional Riemannian manifold $(S^{3},g)$ with positive Ricci curvature, there is no stable minimal surface, and moreover, by [CS85], there exists $L>0$, such that the area of every minimal surface with genus at most 111 is no greater than $L/2$. Therefore, Proposition 3.3 follows from the following general result.
Proposition 6.1 (Generalization of Proposition 3.3).
Let $(S^{3},g_{0})$ be a Riemannian 333-sphere with only finitely many embedded minimal tori and no degenerate stable embedded minimal sphere. For any $\varepsilon>0$ and any $L$ greater than the largest area among all minimal tori in $(S^{3},g_{0})$, there exists a metric $g^{\prime}$ that is $\varepsilon$-close to $g_{0}$ in $C^{\infty}$ such that among all embedded $g^{\prime}$-minimal surfaces with area less than $L$:
  1. (1)
    The number of embedded $g^{\prime}$-minimal tori is the same as that of $g_{0}$-minimal tori.
  2. (2)
    There are only finitely many embedded minimal spheres $\{S_{1},\cdots,S_{q}\}$, and each of them is non-degenerate.
  3. (3)
    The $g^{\prime}$-areas
    $\operatorname{area}_{g^{\prime}}(S_{1}),\dots,\operatorname{area}_{g^{\prime}}(S_{q})$
    are $\mathbb{Z}$-linearly independent. Moreover, their $\mathbb{Z}$-linear combination can never achieve the area of any $g^{\prime}$-minimal tori with $g^{\prime}$-area less than $L$.
Let us recall a compactness result for minimal surfaces essentially following from [CS85, Page 390, Claim(*)] (cf. [And85, Whi87][CM11, Proposition 7.14]).
Lemma 6.2.
In $S^{3}$, given a non-negative integer $\mathfrak{g}_{0}$, a sequence of Riemannian metrics $\{g_{i}\}^{\infty}_{i=1}$ and a sequence of connected embedded closed surfaces $\{\Sigma_{i}\}^{\infty}_{i=1}$, suppose that there exists another Riemannian metrics $g_{\infty}$ such that $g_{i}\to g_{\infty}$ in $C^{3}$ and each $\Sigma_{i}$ is a $g_{i}$-minimal surface satisfying
$\sup_{i}\operatorname{area}_{g_{i}}(\Sigma_{i})<+\infty\,,\quad\sup_{i}\mathfrak{g}(\Sigma_{i})\leq\mathfrak{g}_{0}\,.$
Then there exists a subsequence of $\{\Sigma_{i}\}$, still denoted by $\{\Sigma_{i}\}$, a connected $g_{\infty}$-minimal surface $\Sigma_{\infty}$ and a finite number of points $Z=\{x_{1},\cdots,x_{l}\}$ such that in the varifold topology,
$|\Sigma_{i}|\to k|\Sigma_{\infty}|,\quad k\mathfrak{g}(\Sigma_{\infty})\leq\mathfrak{g}_{0}\,,$
with some $k\in\mathbb{N}^{+}$, and for any $r>0$, $\Sigma_{i}\to\Sigma_{\infty}$ in $S^{3}\setminus B_{r}(Z)$ (multi-)graphically in the $C^{3}$ topology.
Furthermore, if $k=1$, then $Z=\emptyset$, and thus, for sufficiently large $i$, $\mathfrak{g}(\Sigma_{i})=\mathfrak{g}(\Sigma_{\infty})$; if $k>1$, then $\Sigma_{\infty}$ is degenerate stable.
Remark 6.3.
Note that $S^{3}$ does not admit non-orientable closed surfaces, so all $\Sigma_{i}$ and $\Sigma_{\infty}$ are oriented.
Proof.
The first part, except for the genus bound of $\Sigma_{\infty}$, follows from [Whi87, Theorem 3], since a $g_{i}$-minimal surface can be viewed as an embedded $\Phi_{i}$-stationary surface in $(S^{3},g_{\infty})$ for some even elliptic integrand $\Phi_{i}$, and $\Phi_{i}$ converges to the constant function 111. By the area monotonicity formula for $g_{i}$-minimal surfaces, $\Sigma_{i}$ also converges to $\Sigma_{\infty}$ in the Hausdorff distance sense.
Then there exists $r_{0}>0$ such that $\mathfrak{g}(\Sigma_{\infty}\setminus B_{r_{0}}(Z))=\mathfrak{g}(\Sigma_{\infty})$. Since $\Sigma_{i}\to\Sigma_{\infty}$ in $S^{3}\setminus B_{r_{0}}(Z)$ (multi-)graphically in the $C^{3}$ topology, for sufficiently large $i$, $\Sigma_{i}\setminus B_{r_{0}}(Z)$ has exactly $k$ connected components $\{\Gamma^{j}_{i}\}^{k}_{j=1}$, and
$\sum^{k}_{j=1}\mathfrak{g}(\Gamma^{j}_{i})\leq\mathfrak{g}(\Sigma_{j})\leq\mathfrak{g}_{0}\,.$
Hence, for sufficiently large $i$ we have
$\mathfrak{g}(\Sigma_{\infty})=\mathfrak{g}(\Sigma_{\infty}\setminus B_{r_{0}}(Z))=\min_{j}\mathfrak{g}(\Gamma^{j}_{i})\leq\frac{\mathfrak{g}_{0}}{k}\,.$
When $k=1$, $Z=\emptyset$ follows immediately from Allard’s regularity [All72]. When $k>1$, it follows from [Sha17, Claim 5 and Claim 6] that there exists a positive Jacobi fields on $\Sigma_{\infty}$, and thus, $\Sigma_{\infty}$ is degenerate stable. ∎
We prove the following useful lemma.
Lemma 6.4.
Suppose $(S^{3},g_{0})$ contains only finitely many embedded minimal tori, given by $\mathcal{T}\mathrel{\mathop{\ordinarycolon}}=\{T_{1},T_{2},\dots,T_{l}\}$. Then for any $L>0$, there exists an open subset $U\supset\bigcup_{1\leq j\leq l}T_{j}$ in $S^{3}$ and an open neighborhood $\mathcal{O}$ of $g_{0}$ in $C^{\infty}$ such that for every $g\in\mathcal{O}$,
  1. (1)
    Every embedded $g$-minimal 222-sphere $S$ with $g$-area less than $L$ in $S^{3}$ is not entirely contained in $U$;
  2. (2)
    If we further assume that $(S^{3},g_{0})$ does not admit any degenerate stable embedded minimal sphere with area less than $L$, and $g|_{U}=g_{0}|_{U}$. Then every embedded $g$-minimal torus $T$ with $g$-area less than $L$ in $(S^{3},g)$ is inside $\mathcal{T}$.
Proof.
[…]
By Lemma 6.2, up to a subsequence, $S_{j}$ converges to some minimal sphere $S_{\infty}$ with multiplicity in $\bigcup_{1\leq j\leq l}T_{j}$ as varifolds. However, there is no minimal sphere lying in $\bigcup_{1\leq j\leq l}T_{j}$, a contradiction.
Hence, we finish the proof of (1).
Now that we fix $L,U$ as above and argue that for a smaller neighborhood $\mathcal{O}^{\prime}\subset\mathcal{O}$ of $g_{0}$, (b) is true. To see that, suppose again for contradiction that there exists a sequence of metrics $\{g_{j}\}_{J}$ $C^{3}$-converging to $g_{0}$ with $g_{j}|_{U}=g_{0}|_{U}$, as well as a sequence of $g_{j}$-minimal torus $\{T_{j}\notin\mathcal{T}\}$ in $S^{3}$ with $g_{j}$-area less than $L$. By Lemma 6.2, $|T_{j}|$ varifold subconverges to $k|\Sigma|$ for some smooth minimal surface $\Sigma\subset(S^{3},g_{0})$.
Moreover, if $\Sigma$ is also a torus, then $\Sigma\in\mathcal{T}$ and $T_{j}\subset U$ for $j>>1$, but since $g_{j}|_{U}=g_{0}|_{U}$, such $T_{j}$ is also $g_{0}|_{U}$ minimal, which contradicts to the definition of $\mathcal{T}$. Therefore, $\Sigma$ is a sphere and the multiplicity $k\geq 2$. Then again by Lemma 6.2, $\Sigma$ is a degenerate stable minimal sphere in $(S^{3},g_{0})$, which contradicts to the assumption in (2).
This completes the proof of (2). ∎
Now we adapt White’s bumpy metric argument to our setting. In the following, let $3\leq q\leq\infty$ be an integer (or infinity), $L>\sup_{j}\{\operatorname{area}_{g_{0}}(T_{j})\}$ be fixed, $U,\mathcal{O}$ be determined from previous Lemma. Let
$\Gamma^{q}$ :=$\mathrel{\mathop{\ordinarycolon}}=\{C^{q}\text{-metric }g\in\mathcal{O}\text{ conformal to }g_{0}\mathrel{\mathop{\ordinarycolon}}g|_{U}=g_{0}|_{U}\}$
$\mathcal{M}^{q}$ :=$\mathrel{\mathop{\ordinarycolon}}=\{(g,S)\mathrel{\mathop{\ordinarycolon}}g\in\Gamma^{q}\,;\;S\text{ is a minimal sphere in }(S^{3},g),\operatorname{area}_{g}(S)<L\}\,;$
Here we use $C^{q}$-topology on Riemannian metrics and $C^{2,\alpha}$-topology on submanifolds, $\alpha\in(0,1)$.
Lemma 6.5.
For every integer $3\leq q<\infty$, $\mathcal{M}^{q}$ is a separable $C^{q-2}$ Banach manifold modeled on $C^{q}(S^{3})$ and
$\Pi\mathrel{\mathop{\ordinarycolon}}\mathcal{M}^{q}\to\Gamma^{q}\,,\;\;\;(g,S)\mapsto g$
is a $C^{q-2}$ Fredholm map with index 0.
Moreover, $g\in\Gamma^{q}$ is a regular value of $\Pi$ if and only if for every $(g,S)\in\mathcal{M}^{q}$, $S$ is non-degenerate.
Proof.
For every $(\hat{g},\hat{S})\in\mathcal{M}^{q}$ with unit normal field $\hat{\nu}$, we apply [Whi91, Theorem 1.2] with
$\Gamma\mathrel{\mathop{\ordinarycolon}}=\Gamma^{q}\,,\;\;\;\;\;X\mathrel{\mathop{\ordinarycolon}}=C^{2,\alpha}(\hat{S})\,,\;\;\;\;\;Y\mathrel{\mathop{\ordinarycolon}}=C^{\alpha}(\hat{S})\,,\;\;\;\;\;\mathcal{H}\mathrel{\mathop{\ordinarycolon}}=L^{2}(\hat{S})\,,$
$A\mathrel{\mathop{\ordinarycolon}}\Gamma\times X\to\mathbb{R}\,,\;\;(g,w)\mapsto\operatorname{area}_{g}(\mathrm{graph}_{\hat{S},\hat{g}}(w))\,,$
$H\mathrel{\mathop{\ordinarycolon}}\Gamma\times X\to Y\,,\;\text{ given by }\langle H(g,w),v\rangle_{L^{2}}=\left.\frac{d}{dt}\right|_{t=0}A(g,w+tv)\,.$
where every norm is measured under metric $\hat{g}$. Note that $\Gamma^{q}$ can be identified as an open subset of
$C^{q}_{U}(S^{3})\mathrel{\mathop{\ordinarycolon}}=\{f\in C^{q}(S^{3})\mathrel{\mathop{\ordinarycolon}}f|_{U}=0\}\,.$
endowed with $C^{q}$-topology. This is also a separable Banach space.
Notice that near $(\hat{g},\hat{S})$, $\mathcal{M}^{q}=\{(g,S)\in\Gamma\times X\mathrel{\mathop{\ordinarycolon}}H(g,S)=0\}$. And as is verified in [Whi91, Theorem 1.1],
$D_{2}H(\hat{g},0)=-(\Delta_{\hat{S}}+|A_{\hat{S}}|^{2}+\operatorname{Ric}_{\hat{g}}(\hat{\nu},\hat{\nu}))=\mathrel{\mathop{\ordinarycolon}}L_{\hat{S},\hat{g}}\mathrel{\mathop{\ordinarycolon}}X\to Y$
is a Fredholm map with index 0. Hence to prove the first part of the Lemma, it suffices to verify the condition (C) in [Whi91, Theorem 1.2], namely, for every $0\neq\kappa\in\mathrm{ker}D_{2}H(\hat{g},0)$, there exists a one parameter family $\{\gamma(s)\in\Gamma^{q}\}_{s\in(-1,1)}$ such that $\gamma(0)=\hat{g}$ and
(6.1) $\left(\frac{\partial^{2}}{\partial s\partial t}\right)_{s=t=0}A(\gamma(s),t\kappa)\neq 0\,.$
To see this, let $\gamma(s)=(1+sf)\hat{g}$ where $f\in C^{q}_{U}(S^{3})$. According to the calculation in [Whi91, Theorem 2.1],
[…]
Recall that by Lemma 6.4, $\hat{S}\setminus\bar{U}\neq\emptyset$; And by the unique continuation property, $\kappa|_{\hat{S}\setminus U}$ is not identically 0. Therefore, (6.1) is true by taking $f\in C_{c}^{q}(S^{3}\setminus\bar{U})$ such that $\hat{\nu}(f)$ is an $L^{2}$-approximation of $\kappa\cdot\chi_{S^{3}\setminus U}$.
To study the regular value of $\Pi$, first notice that for every $(g^{\prime},S^{\prime})\in\mathcal{M}^{q}$,
$\operatorname{Tan}_{(g^{\prime},S^{\prime})}\mathcal{M}^{q}$ =$=\{(f,w)\mathrel{\mathop{\ordinarycolon}}f\in C^{q}_{U}(S^{3}),\;L_{S^{\prime},g^{\prime}}=\hat{\nu}(f)\text{ on }S^{\prime}\}\,;$
$d\Pi|_{(g^{\prime},S^{\prime})}$ :$\mathrel{\mathop{\ordinarycolon}}\operatorname{Tan}_{(g^{\prime},S^{\prime})}\mathcal{M}^{q}\to C^{q}_{U}(S^{3})\,,\;\;(f,w)\mapsto f\,.$
Thus, if $g^{\prime}\in\Gamma^{q}$ is a regular value of $\Pi$, then for every $(g^{\prime},S^{\prime})\in\mathcal{M}^{q}$ with unit normal field $\nu^{\prime}$, every $\kappa\in\mathrm{ker}L_{S^{\prime},g^{\prime}}$ and every $\phi\in C^{q}_{c}(S^{\prime}\setminus\bar{U})$, we take $f\in C^{q}_{U}(S^{3})$ such that $\nu^{\prime}(f)=\phi$, then there exists $w_{\phi}\in C^{2,\alpha}(S^{\prime})$ such that $L_{S^{\prime},g^{\prime}}w_{\phi}=\nu^{\prime}(f)=\phi$. Then
$0=\int_{S^{\prime}}w_{\phi}\cdot L_{S^{\prime},g^{\prime}}\kappa=\int_{S^{\prime}}\kappa\cdot L_{S^{\prime},g^{\prime}}w_{\phi}=\int_{S^{\prime}}\kappa\phi\,.$
In other words, $\kappa\perp C^{q}_{c}(S^{\prime}\setminus\bar{U})$. Therefore, $\kappa|_{S^{\prime}\setminus\bar{U}}=0$. Again by Lemma 6.4 and the unique continuation property, $\kappa\equiv 0$ on $S^{\prime}$. Hence, $\mathrm{ker}L_{S^{\prime},g^{\prime}}=0$, that is, $S^{\prime}$ is non-degenerate. ∎
To obtain the third assertion of Proposition 6.1, we need the following lemma in linear algebra.
Lemma 6.6.
Let $k,l\geq 1$. Given $\mathbf{a}\in\mathbb{R}^{k},\mathbf{b}\in\mathbb{R}^{l}$, consider the linear function
$f_{\mathbf{a}}\mathrel{\mathop{\ordinarycolon}}\mathbb{Z}^{k}\to\mathbb{R}\,,\quad\mathbf{u}\mapsto\mathbf{u}\cdot\mathbf{a}\,;\qquad h_{\mathbf{b}}\mathrel{\mathop{\ordinarycolon}}\mathbb{Z}^{l}\to\mathbb{R}\,,\quad\mathbf{v}\mapsto\mathbf{v}\cdot\mathbf{b}\,.$
Then for every $\mathbf{b}\in\mathbb{R}^{l}$,
:=$\mathcal{R}\mathrel{\mathop{\ordinarycolon}}=\{\mathbf{a}\in\mathbb{R}^{k}\mathrel{\mathop{\ordinarycolon}}\mathrm{ker}(f_{\mathbf{a}})=\{0\}\,,\;f_{\mathbf{a}}(\mathbb{Z}^{k})\cap h_{\mathbf{b}}(\mathbb{Z}^{l})=\{0\}\}$
is a countable intersection of open dense subset of $\mathbb{R}^{k}$ (and thus dense in $\mathbb{R}^{k}$).
Proof.
[…]
$\mathcal{R}_{N}\mathrel{\mathop{\ordinarycolon}}=\left\{\mathbf{a}\in\mathbb{R}^{k}\mathrel{\mathop{\ordinarycolon}}\mathrm{ker}(f_{\mathbf{a}})\cap\Lambda_{k,N}\subset\{0\}\,,\;f_{\mathbf{a}}(\mathbb{Z}^{k}\cap\Lambda_{k,N})\cap h_{\mathbf{b}}(\mathbb{Z}^{l}\cap\Lambda_{l,N})\subset\{0\}\right\}$
over $N\in\mathbb{N}$, where each $\mathcal{R}_{N}$ is open and dense in $\mathbb{R}^{k}$. ∎
Proof of Proposition 6.1.
For every fixed integer $q\geq 3$, $\varepsilon>0$ and $L$ greater than the largest area among all minimal tori in $(S^{3},g_{0})$, let $U\subset S^{3}$ and $\mathcal{O}\ni g_{0}$ be determined by Lemma 6.4. By the Sard-Smale Theorem, the regular values of $\Pi\mathrel{\mathop{\ordinarycolon}}\mathcal{M}^{q}\to\Gamma^{q}$ forms a residual (and hence dense) subset $\Gamma^{q}_{bumpy}\subset\Gamma^{q}$. By Lemma 6.26.46.5 and Sharp’s compactness theorem [Sha17], the first two assertion of Proposition 6.1 hold for every $g\in\Gamma^{q}_{bumpy}$. Then by the argument in [Whi17](See also the proof of [ACS18, Theorem 9]), the first two assertion of Proposition 6.1 also hold for every $g\in\Gamma^{q}_{bumpy}\cap\Gamma^{\infty}$, which is dense in $\Gamma^{\infty}$.
Moreover, for each $g^{\prime}_{0}\in\Gamma^{q}_{bumpy}\cap\Gamma^{\infty}$, there are only finitely many minimal spheres with area at most $L$. Therefore, by Lemma 6.2, and Sharp’s compactness theorem again, there exists a neighborhood $\mathcal{U}$ of $g^{\prime}_{0}$ in $\Gamma^{\infty}$, such that $\mathcal{U}\subset\Gamma^{q}_{bumpy}\cap\Gamma^{\infty}$. In other words, $\Gamma^{q}_{bumpy}\cap\Gamma^{\infty}$ is both open and dense in $\Gamma^{\infty}$.
To obtain the third assertion, let $g_{0}^{\prime}\in\Gamma^{q}_{bumpy}\cap\Gamma^{\infty}$ be $\varepsilon/2$-$C^{\infty}$ close to $g_{0}$ and let
$\{S_{1},S_{2},\dots,S_{k}\}$
[…]
$B_{r}(p_{j})\cap U\cup\bigcup_{i\neq j}S_{i}=\emptyset\,.$
by deforming in $B_{r}(p_{j})$, one can construct an $\mathbb{R}^{k}$-parametrized family of metrics $\{g(\mathbf{a})^{\prime}\}_{\mathbf{a}\in\mathbb{R}^{k}}\subset\Gamma^{q}_{bumpy}\cap\Gamma^{\infty}$ such that $g(0)^{\prime}=g_{0}^{\prime}$, all the embedded $g(\mathbf{a})^{\prime}$-minimal spheres $\{S_{j}(\mathbf{a})\}_{j=1}^{k}$ are continuous deformation of $\{S_{j}\}_{j=1}^{k}$ and that
$\mathbf{a}\mapsto\left(\operatorname{area}_{g(\mathbf{a})^{\prime}}(S_{1}(\mathbf{a})),\operatorname{area}_{g(\mathbf{a})^{\prime}}(S_{2}(\mathbf{a})),\dots,\operatorname{area}_{g(\mathbf{a})^{\prime}}(S_{k}(\mathbf{a}))\right)$
is a local diffeomorphism near 0 onto its image. By Lemma 6.6 with
$\mathbf{b}=(\operatorname{area}_{g_{0}}(T_{1}),\operatorname{area}_{g_{0}}(T_{2}),\dots,\operatorname{area}_{g_{0}}(T_{l}))\,,$
in a small neighborhood of 0, those $\mathbf{a}\in\mathbb{R}^{k}$ such that $g(\mathbf{a})^{\prime}$ satisfies the third assertion of Proposition 6.1 is residual and hence dense. We choose such a $g^{\prime}=g(\mathbf{a})^{\prime}$ within $\varepsilon/2$-neighborhood of $g_{0}^{\prime}$ in $\Gamma^{\infty}$, which satisfies all the assertion in Proposition 6.1. ∎
7. Min-max results II: Convergence as currents
Theorem 3.5 is analogous to Proposition 4.10 and Theorem 4.11 in [MN21]. However, in our definition of homotopy class (Definition 2.10, two Simon-Smith families must be connected through an isotopy, which precludes the use of the well-known Almgren interpolation scheme. Therefore, the proofs in [MN21] cannot be followed verbatim to justify Theorem 3.5. Instead, we will use results from [Ket19] and [DP10] to construct appropriate isotopies.
Let $\Phi^{\prime}\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}^{*}(M)$ be a Simon-Smith family of genus $\leq\mathfrak{g}_{0}$ in the homotopy class $\Lambda$ from Theorem 3.5. Suppose that $X$ is a cubical subcomplex of some $I(m,k)$ and $K(m)=3^{m3^{m}}$. In addition, by Remark 2.11, we may assume that $N_{P}\equiv N\in\mathbb{N}$ within the homotopy class $\Lambda$. For $L=\mathbf{L}(\Lambda)$, since $\mathcal{W}_{L,\leq\mathfrak{g}_{0}}$ consists of finitely many varifolds associated with connected multiplicity-one minimal surfaces, we can enumerate these surfaces as
$\Gamma_{1},\Gamma_{2},\cdots,\Gamma_{q}\,.$
We fix $r_{0}>0$ such that for each $\Gamma_{k}$, each point $p\in\Gamma_{k}$, and any $r\in(0,2r_{0})$,
$B_{r}(p)\cap\Gamma_{k}$
is a topological disk, and moreover, $\partial B_{r}(p)$ is a mean-convex sphere with transversal intersection with $\Gamma_{k}$:
$\partial B_{r}(p)\pitchfork\Gamma_{k}\,.$
Lemma 7.1 ([Pit81, 2.1 (f)]).
Suppose that $T,T_{1},T_{2},\cdots\in\mathcal{Z}_{2}(M;\mathbb{Z}_{2})$, $\lim_{i}T_{i}=T$ in the flat topology and $\lim_{i}|T_{i}|=V\in\mathcal{V}_{2}(M)$ in the weak topology. Then
$\|T\|\leq\|V\|\,.$
Corollary 7.2.
For any $r>0$, there exists $r_{1}>0$ and $\delta_{1}>0$ such that for every $V\in\mathcal{W}_{L,\leq\mathfrak{g}_{0}}$ associated with some $\Gamma_{k}$, the set
$\mathcal{D}_{k}\mathrel{\mathop{\ordinarycolon}}=\{T\in\mathcal{Z}_{2}(M;\mathbf{F};\mathbb{Z}_{2})\mathrel{\mathop{\ordinarycolon}}\mathbf{F}(|T|,V)<r_{1}\}$
has exactly two connected components $\mathcal{D}^{0}_{k}\cup\mathcal{D}^{1}_{k}$ characterized by
$\mathcal{D}^{0}_{k}$ =$=\{T\in\mathcal{D}_{k}\mathrel{\mathop{\ordinarycolon}}\mathcal{F}(T,0)<\delta_{1}\}\,,$
$\mathcal{D}^{1}_{k}$ =$=\{T\in\mathcal{D}_{k}\mathrel{\mathop{\ordinarycolon}}\mathcal{F}(T,0)>\delta_{1}\}=\{T\in\mathcal{D}_{k}\mathrel{\mathop{\ordinarycolon}}\mathbf{F}(T,[\Gamma_{k}])<r\}\,.$
Proof.
By Lemma 7.1, every sequence $\{T_{i}\}_{i}\subset\mathcal{Z}_{2}(M;\mathbb{Z}_{2})$ satisfying $\lim_{i}T_{i}=T$ and $\lim_{i}|T_{i}|=V$ has
$\|T\|\leq\|V\|\,.$
Since $V$ is associated with a connected multiplicity-one minimal surface $\Gamma_{k}$ and $\partial T=0$, by the constancy theorem, we have either $T=[\Gamma_{k}]$ or $T=0$. Thus, by choosing suitable $r_{1}$ and $\delta_{1}$, the conclusion follows immediately. ∎
Fix $r$ from Theorem 3.5 and then choose $r_{1}$ and $\delta_{1}$ from the previous Corollary. The crucial proposition we need is the following:
Proposition 7.3.
For every $\alpha\in(0,2r_{0})$, there exists $\varepsilon\in(0,r_{1})$ such that for any punctate surface $\Sigma\in\mathcal{S}(M)$ and any $\delta>0$, if for some $\Gamma_{k}$ and some $p\in\Gamma_{k}$,
  1. (i)
    $\mathbf{F}(|\Sigma|,|\Gamma_{k}|)<\varepsilon$,
  2. (ii)
    $\Sigma\cap B_{\alpha}(p)$ is smooth,
  3. (iii)
    $\mathfrak{g}(\Sigma\cap B_{\alpha}(p))=0$,
  4. (iv)
    $[\Sigma]\in\mathfrak{a}(B_{\alpha}(p);\varepsilon,\delta)$,
then $\mathcal{F}([\Sigma],0)>\delta_{1}$.
Proof.
Suppose for the sake of contradiction that for $\alpha>0$, $\Gamma_{k}$ and $p\in\Gamma_{k}$, there exists a sequence of punctate surfaces $\{\Sigma_{i}\}\subset\mathcal{S}(M)$ and a sequence of positive number $\{\delta^{(i)}\}$ such that for all large $i$,
  1. (i)
    $\mathbf{F}(|\Sigma_{i}|,|\Gamma_{k}|)<\frac{1}{i}$,
  2. (ii)
    $\Sigma_{i}\cap B_{\alpha}(p)$ is smooth,
  3. (iii)
    $\mathfrak{g}(\Sigma_{i}\cap B_{\alpha}(p))=0$,
  4. (iv)
    $[\Sigma_{i}]\in\mathfrak{a}(B_{\alpha}(p);\frac{1}{i},\delta^{(i)})$,
but $\mathcal{F}([\Sigma_{i}],0)\leq\delta_{1}$. Hence, by Lemma 7.1 and Corollary 7.2, we have $\mathcal{F}([\Sigma_{i}],0)<\delta_{1}$ and
$\lim_{i\to\infty}\mathcal{F}([\Sigma_{i}],0)=0\,.$
We can choose $\bar{r}\in(\alpha/2,\alpha)\subset(0,2r_{0})$ such that for each $i$, $\partial B_{\bar{r}}(p)\pitchfork\Sigma_{i}$. Note that by definition, we also have $|\Sigma_{i}|\in\mathfrak{a}(B_{\bar{r}}(p);\frac{1}{i},\delta^{(i)})$.
Following [Ket19, Section 3.2], let $\mathfrak{Is}_{i}(\Sigma_{i},B_{\bar{r}}(p))$ denote the set of all isotopies $\phi$ of $M$ supported in $B_{\bar{r}}(p)$ such that for all $t\in[0,1]$,
$\mathcal{H}^{2}(\phi(t,\Sigma_{i}))\leq\mathcal{H}^{2}(\Sigma_{i})+\delta^{(i)}\,.$
[…]
$\lim_{l\to\infty}\mathcal{H}^{2}(\phi^{l}(1,\Sigma_{i}))=\inf_{\phi\in\mathfrak{Is}_{i}(\Sigma_{i},B_{\bar{r}}(p))}\mathbf{M}(\phi(1,\Sigma_{i}))\,,$
and let $V_{i}$ be the varifold limit of $\phi^{l}(1,\Sigma_{i})$.
It follows from [Ket19, Lemma 3.12] that as $i\to\infty$,
$V_{i}\to|\Gamma_{k}|$
in the varifold topology, and the convergence is smooth in compact subsets of $B_{\bar{r}}(p)$. By [DP10, Proposition 3.2], we also know that $V_{i}\cap B_{\bar{r}}(p)$ is associated with a smooth minimal surface $\Delta_{i}$ such that $\partial\Delta_{i}=\partial(\Sigma_{i}\cap B_{\bar{r}}(p))$.
By [Ket19, Proposition 4.7], we have
$\mathfrak{g}(\Delta_{i})\leq\mathfrak{g}(\Sigma_{i}\cap B_{\bar{r}}(p))=0\,,$
so $\Delta_{i}$ is a minimal disk. In particular, $\Delta_{i}$ separates $B_{\bar{r}}(p)$, and $V_{i}$ is associated with a cycle $T_{i}\in\mathcal{Z}_{2}(M;\mathbb{Z}_{2})$.
On one hand, note that
$T_{i}\llcorner(M\setminus B_{\bar{r}}(p))=[\Sigma_{i}]\llcorner(M\setminus B_{\bar{r}}(p))\,.$
In particular,
$T_{i}\to 0$
outside $B_{2r_{0}}(p)$.
On the other hand, in $B_{\alpha/2}(p)$, for sufficiently large $i$, $\Delta_{i}$ is a normal graph over $\Sigma_{i}$, $B_{\alpha/2}(p)\setminus\Delta_{i}$ has two connected components $E^{1}_{i}$ and $E^{2}_{i}$ and $B_{\alpha/2}(p)\setminus\Gamma_{k}$ has two connected components $E^{1}$ and $E^{2}$ satisfying
$\min\{\operatorname{vol}(E^{1}_{i}),\operatorname{vol}(E^{2}_{i})\}\geq\min\{\operatorname{vol}(E^{1}),\operatorname{vol}(E^{2})\}/2>0\,.$
Therefore, we have
$\liminf_{i}\mathcal{F}(T_{i},0)\geq\min\{\operatorname{vol}(E^{1}),\operatorname{vol}(E^{2})\}/2>0\,.$
Since $|T_{i}|=V_{i}\to|\Gamma_{k}|$, by Lemma 7.1, we have $\lim T_{i}=[\Gamma_{k}]$. This contradicts that $T_{i}\to 0$ outside $B_{2r_{0}}(p)$. ∎
Proof of Theorem 3.5.
As above, for $r>0$, we choose $r_{1}$ and $\delta_{1}$ from Corollary 7.2 and also choose $\varepsilon$ from Proposition 7.3 with the parameters $r_{1}$ and $\delta_{1}$.
Let $\alpha=r_{0}/(16\cdot 8^{\bar{N}})$ and choose $\varepsilon$ from Proposition 7.3. We set $r_{3}\mathrel{\mathop{\ordinarycolon}}=\min(r,r_{1},\varepsilon)$.
For each $\Gamma_{k}$, choose $p_{k}\in\Gamma_{k}$ and for each $i=1,\cdots,\bar{N}$, we set $r_{i}=\frac{7r_{0}}{8^{i}}$ and $s_{i}=\frac{5r_{0}}{8^{i}}$, and we obtain a $\bar{N}$-admissible collection of annuli, $\{A^{k}_{i}\mathrel{\mathop{\ordinarycolon}}=A(p_{k};s_{i},r_{i})\}$. We also choose $\{p^{1}_{k},\cdots,p^{\bar{N}}_{k}\}\subset\Gamma_{k}$ such that for each $i=1,\cdots,\bar{N}$,
$\operatorname{dist}(p_{k},p^{i}_{k})=\frac{3r_{0}}{4\cdot 8^{i-1}}\,.$
and thus, $B(p^{i}_{k},\alpha)\subset A^{k}_{i}$.
By Theorem 2.16, there exists a pulled-tight sequence $\{\Phi_{i}\}$ in $\Lambda$ and some $\eta_{3}\in(0,r_{3})$ such that for each $i$ and each $x\in X$,
$\mathcal{H}^{2}(\Phi_{i}(x))\geq L-\eta_{3}\implies|\Phi_{i}(x)|\in\mathbf{B}^{\mathbf{F}}_{r_{3}}(\mathcal{W}_{L,\leq\mathfrak{g}_{0}})\,.$
It follows from Corollary 7.2 that the compact set
$Y_{i}\mathrel{\mathop{\ordinarycolon}}=\{x\in X\mathrel{\mathop{\ordinarycolon}}\mathcal{H}^{2}(\Phi_{i}(x))\geq L-\eta_{3}\}$
can be decomposed as two disjoint compact sets $Y^{0}_{i}$ and $Y^{1}_{i}$:
$Y^{0}_{i}$ =$=\{x\in Y_{i}\mathrel{\mathop{\ordinarycolon}}\mathcal{F}([\Phi_{i}(x)],0)<\delta_{1}\}$
$Y^{1}_{i}$ =$=\{x\in Y_{i}\mathrel{\mathop{\ordinarycolon}}\mathcal{F}([\Phi_{i}(x)],0)>\delta_{1}\}=\{x\in Y_{i}\mathrel{\mathop{\ordinarycolon}}\mathbf{F}([\Phi_{i}(x)],[\mathcal{W}_{L,\leq\mathfrak{g}_{0}}])<r\}\,.$
Now for each $x\in Y^{0}_{i}$, we can choose $k$ such that $\mathbf{F}(|\Phi_{i}(x)|,|\Gamma_{k}|)<r_{3}\leq\varepsilon$. Since $\mathfrak{g}(\Phi_{i}(x))\leq\mathfrak{g_{0}}$ and we can choose a set of at most $N$ points $P_{\Phi_{i}}(x)$ such that $\Phi_{i}(x)\setminus P_{\Phi_{i}}(x)$ is a smooth surface, there are at least $K(m)$ points $\{p^{j_{l}}_{k}\}^{K(m)}_{j_{l}}$ from $\{p^{j}_{k}\}^{\bar{N}}_{j=1}$ where for each $l$,
  1. (1)
    $\Phi_{i}(x)\cap B_{\alpha}(p^{j_{l}}_{k})$ is smooth,
  2. (2)
    $\mathfrak{g}(\Phi_{i}(x)\cap B_{\alpha}(p^{j_{l}}_{k}))=0$,
hence, $\Sigma$ admits an $(\varepsilon,\frac{\eta_{3}}{1000K(m)})$ deformation in $B_{\alpha}(p^{j_{l}}_{k})$.
By Proposition 4.3, we can obtain $\Phi^{*}_{i}$ for each $\Phi_{i}$. Moreover, for sufficiently large $i$, since $\eta_{3}<\varepsilon$, we have the following properties.
For every $x\in Y^{0}_{i}$,
$\mathcal{H}^{2}(\Phi^{*}_{i}(x))$ $<\mathcal{H}^{2}(\Phi_{i}(x))-\varepsilon+(3^{m}-1)\frac{\eta_{3}}{1000K(m)}$
$<\mathcal{H}^{2}(\Phi_{i}(x))-\eta_{3}/2$
$<L-\eta_{3}/2\,.$
For every $x\in X\setminus Y_{i}$,
$\mathcal{H}^{2}(\Phi^{*}_{i}(x))$ $<\mathcal{H}^{2}(\Phi_{i}(x))+3^{m}\frac{\varepsilon}{1000K(m)}$
$<L-\eta_{3}+3^{m}\frac{\eta_{3}}{1000K(m)}$
$<L-\eta_{3}/2\,.$
Since $Y^{0}_{i}$ and $Y^{1}_{i}$ are disjoint compact set, we also have $\Phi^{*}_{i}(x)=\Phi_{i}(x)$ holds for every $x\in Y^{1}_{i}$.
Therefore, for every $x\in X$ with $\mathcal{H}^{2}(\Phi^{*}_{i}(x))\geq L-\eta_{3}/2$, then $x\in Y^{1}_{i}$ and thus,
$[\Phi_{i}(x)]\in\mathbf{B}^{\mathbf{F}}_{r_{3}}([\mathcal{W}_{L,\leq\mathfrak{g}_{0}}])\subset\mathbf{B}^{\mathbf{F}}_{r}([\mathcal{W}_{L,\leq\mathfrak{g}_{0}}])\,,$
which completes the proof with $\eta=\eta_{3}/2$ and $\Phi=\Phi^{*}_{i}$. ∎
8. Min-max results III: Interpolation
The goal of this section is to prove Theorem 3.6. Let $(M,g)$ be a 333-dimensional closed Riemannian manifold.
The key is the following deformation proposition. Intuitively, for a Simon-Smith family of genus 1$\leq$ 1≤ 1 very close to spheres, we show that we can pinch all the small necks and shrink all small connected components in a continuous way to obtain a Simon-Smith family of genus 0.
Proposition 8.1.
[…]
If $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ is a Simon-Smith family of genus 1$\leq$ 1≤ 1, $N_{P}(\Phi)<q$ and $S\mathrel{\mathop{\ordinarycolon}}X\to\mathscr{S}$ is a map such that for every $x\in X$,
$\mathbf{F}([\Phi(x)],[S(x)])\leq\delta$
then there exists a Simon-Smith family $H\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}(M)$ of genus 1$\leq$ 1≤ 1 such that:
  1. (1)
    For every $x\in X$, $H(0,x)=\Phi(x)$, $\mathfrak{g}(H(1,x))=0$.
  2. (2)
    For every $x\in X$ and $0\leq t\leq t^{\prime}\leq 1$,
    $\mathcal{H}^{2}(H(t,x))\leq\mathcal{H}^{2}(\Phi(x))+\varepsilon\,,\quad\mathfrak{g}(H(t^{\prime},x))\leq\mathfrak{g}(H(t,x))\,.$
  3. (3)
    For each $x\in X$, the Simon-Smith family $\{H(x,t)\}_{t\in[0,1]}$ is a pinch-off process.
We will prove the proposition using the following technical lemmas.
8.1. Nontrivial loop in a small ball
In this subsection, we show that if a (singular) torus is sufficiently close to a sphere, we can find a loop contained within a small ball. In the next subsection, we will explain how to pinch the torus along the boundary of the small ball and then shrink all connected components inside to a point. The small size of the ball helps control the area increase, as required in Proposition 8.1 (2). This existence result is closely related to the well-known systolic inequality.
Theorem 8.2 (Loewner torus inequality [Pu52, Section 3], Gromov’s homological systolic inequality [Gro96, Section 2.C]).
Given a smooth orientable closed surface $\Gamma$ of positive genus, there exists a universal constant $C>0$ such that
$\operatorname{sys}H_{1}(\Gamma;\mathbb{Z}_{2})\mathrel{\mathop{\ordinarycolon}}=\inf\left\{\mathcal{H}^{1}(c)\mathrel{\mathop{\ordinarycolon}}[c]\neq 0\in H_{1}(\Gamma;\mathbb{Z}_{2})\right\}\leq C\sqrt{\mathcal{H}^{2}(\Gamma)}\,.$
In particular, there exists a simple closed geodesic $c\in\Gamma$ such that $[c]\neq 0\in H_{1}(\Gamma;\mathbb{Z}_{2})$ and $\mathcal{H}^{1}(c)\leq C\sqrt{\mathcal{H}^{2}(\Gamma)}$.
Let us first prove the existence of a nontrivial small loop for embedded tori sufficiently close to an embedded sphere.
Lemma 8.3 (Existence of small nontrivial loop: Smooth surfaces).
In a closed Riemannian 333-manifold $(M,g)$, let $S$ be an embedded sphere. For any $\varepsilon\in(0,1)$, there exists $\delta=\delta_{0}(S,\varepsilon)\in(0,\varepsilon)$ with the following property.
In $(M,g)$, for any embedded torus $T$ with $\mathbf{F}([S],[T])\leq\delta$, there exists $p\in M$ and a simple loop $\gamma\subset\Sigma\cap B(p,\varepsilon)$ such that the closed surface obtained by neck-pinch along $\gamma$ of $\Sigma$ has genus 0.
Proof.
Without loss of generality, we may assume that $S$ is a unit round 222-sphere, and it has a tubular neighborhood $N_{\zeta}(S)\cong(S\times[-\zeta,\zeta],g_{S}+dt^{2})$ with $\zeta<\varepsilon/10$. We denote by
$\pi\mathrel{\mathop{\ordinarycolon}}N_{\zeta}(S)\to S$
the corresponding closest-point projection.
The general case follows by changing the metric $g$ and taking a smaller $\delta$.
Step 0. Set up
Fix a triangulation $\mathcal{T}$ of $S$ such that:
  1. (0.i)
    For any $\Delta\in\mathcal{T}^{(2)}$, $\mathcal{H}^{2}(\Delta)\in\left(0,\frac{\varepsilon^{2}}{100}\right)$.
  2. (0.ii)
    For any $e\in\mathcal{T}^{(1)}$, $\mathcal{H}^{1}(e)\in\left(0,\frac{\varepsilon^{2}}{100}\right)$.
Let $m$ denote the total number of cells in $\mathcal{T}$. Set $\tau\mathrel{\mathop{\ordinarycolon}}=\frac{1}{1000m^{4}}$.
For convenience, in the following, for any $v\in\mathcal{T}^{(0)}$, $e\in\mathcal{T}^{(1)}$, $\Delta\in\mathcal{T}^{(2)}$ and $\varphi\in\operatorname{Isom}(S)=SO(3)$, we define:
  • the vertical line $\mathrm{I}_{v,\varphi}\mathrel{\mathop{\ordinarycolon}}=\pi^{-1}(\varphi(v))$;
  • the rectangle $\mathrm{R}_{e,\varphi}\mathrel{\mathop{\ordinarycolon}}=\pi^{-1}(\varphi(e))$;
  • the cylinder $\mathrm{Cyl}_{\Delta,\varphi}\mathrel{\mathop{\ordinarycolon}}=\pi^{-1}(\partial\varphi(\Delta))$;
  • the prism $\mathrm{Pri}_{\Delta,\varphi}\mathrel{\mathop{\ordinarycolon}}=\pi^{-1}(\varphi(\Delta))$.
Step 1. Existence of a good triangulation
We show that there exists $\delta^{\prime}\in(0,1)$ with the following property:
For any closed embedded surface $T\subset N_{\zeta}(S)$ and $\mathbf{F}([S],[T])<\delta^{\prime}$, there exist $\varphi\in\operatorname{Isom}(S)=SO(3)$ such that:
  1. (1)
    For any $v\in\mathcal{T}^{(0)}$, $\mathrm{I}_{v,\varphi}\pitchfork T$ is a single point in $T$.
  2. (2)
    For any $\Delta\in\mathcal{T}^{(2)}$, $\mathrm{Cyl}_{\Delta,\varphi}\pitchfork T=\bigcup^{n_{\Delta}}_{k=1}c_{\Delta,k}$ is a finite union of disjoint simple loops in $T$ satisfying
    $\frac{\mathcal{H}^{1}(\mathrm{Cyl}_{\Delta,\varphi}\cap T)}{\mathcal{H}^{1}(\partial\Delta)}\in(1-\tau,1+\tau)\,,$
    and
    $\frac{\mathcal{H}^{2}(\mathrm{Pri}_{\Delta,\varphi})\cap T)}{\mathcal{H}^{2}(\Delta)}\in(1-\tau,1+\tau)\,,$
Moreover, for any $\Delta\in\mathcal{T}^{(2)}$, we can arrange $\{c_{\Delta,k}\}^{n_{\Delta}}_{k=1}$ in such an order that:
  1. (3)
    $c_{\Delta,1}$ satisfies $[c_{\Delta,1}]\neq 0\in H_{1}(\mathrm{Cyl}_{\Delta,\varphi};\mathbb{Z}_{2})$ and
    $\frac{\mathcal{H}^{1}(c_{\Delta,1})}{\mathcal{H}^{1}(\partial\Delta)}\in(1-\tau,1+\tau)\,;$
    we call it the long loop for $\Delta$.
  2. (4)
    Every $c_{\Delta,k}$ with $k\in\{2,\dots,n_{\Delta}\}$ satisfies $[c_{\Delta,k}]=0\in H_{1}(\mathrm{Cyl}_{\Delta,\varphi};\mathbb{Z}_{2})$, $c_{\Delta,k}\cap\bigcup_{v\in\mathcal{T}^{(0)}}\mathrm{I}_{v,\varphi})=\emptyset$, and
    $\frac{\sum^{n_{\Delta}}_{k=2}\mathcal{H}^{1}(c_{\Delta,k})}{\mathcal{H}^{1}(\partial\Delta)}\in(0,\tau)\,;$
    we call them the short loops for $\Delta$.
Let $\mu$ be the Haar measure on $\operatorname{Isom}(S)=SO(3)$.
By the isoperimetric inequality, we choose $\delta$ small enough, depending on $S$ and $\zeta$, such that $\mathbf{F}([S],[T])\leq\delta$ and $T\subset N_{\zeta}(S)$ imply $T$ and $S$ are homologous to each other in $N_{\zeta}(S)$ with $\mathbb{Z}_{2}$ coefficients.
This implies that for any $v\in\mathcal{T}^{(0)}$, $e\in\mathcal{T}^{(1)}$, $\Delta\in\mathcal{T}^{(2)}$, and $\varphi\in\operatorname{Isom}(S)$, if $\mathrm{I}_{v,\varphi}$ intersects $T$ transversely, then
$[\mathrm{I}_{v,\varphi}\cap T]=[\varphi(v)]\neq 0\in H_{0}(\mathrm{I}_{v,\varphi};\mathbb{Z}_{2})\,;$
if $\mathrm{R}_{e,\varphi}$ intersects $T$ transversely, then
$[\mathrm{R}_{e,\varphi}\cap T]=[\varphi(e)]\neq 0\in H_{1}(\mathrm{R}_{e,\varphi};\mathbb{Z}_{2})\,;$
if $\mathrm{Cyl}_{\Delta,\varphi}$ intersects $T$ transversely, then
(8.1) $[\mathrm{Cyl}_{\Delta,\varphi}\cap T]=[\varphi(\partial\Delta)]\neq 0\in H_{1}(\mathrm{Cyl}_{\Delta,\varphi};\mathbb{Z}_{2})\,,$
and
$[\mathrm{Pri}_{\Delta,\varphi}\cap T]=[\varphi(\Delta)]\neq 0\in H_{2}(\mathrm{Pri}_{\Delta,\varphi};\mathbb{Z}_{2})\,,$
By Sard’s theorem, the transversality holds for a full measure subset $E$ of $SO(3)$. Note that these give lower bounds on $\mathcal{H}^{0}(\mathrm{I}_{v,\varphi}\cap T)$, $\mathcal{H}^{1}(\mathrm{R}_{e,\varphi}\cap T)$ and $\mathcal{H}^{2}(\mathrm{Cyl}_{\Delta,\varphi}\cap T)$ for all $\varphi\in E$.
Therefore, by the coarea formula and possibly taking $\delta$ smaller, there exists a subset $W_{T}\subset E$ with $\mu(W_{T})\geq 1/2$ such that (1) and (2) hold. Meanwhile, (3) and (4) follow from (8.1).
Step 2. Special case
By Step 1, for every torus $T\subset N_{\zeta}(S)$ with $\mathbf{F}([T],[S])<\delta^{\prime}$, we can choose a $\varphi\in W_{T}$. Then for every $\Delta\in\mathcal{T}^{(2)}$, $v\in\Delta^{(0)}$,
$\mathrm{I}_{v,\varphi}\cap T=\mathrm{I}_{v,\varphi}\cap c_{\Delta,1}\,.$
Therefore, every short loop $c_{\Delta,k}$ with $k\geq 2$ lies in the interior of $\mathrm{R}_{e,\varphi}$ for some $e\in\Delta^{(1)}$. In particular, every short loop for $\Delta\in\mathcal{T}^{(2)}$ coincides with exactly one short loop in another $\Delta^{\prime}\in\mathcal{T}^{(2)}$.
First, we cut $T$ along every loop $\bigcup_{\Delta\in\mathcal{T}^{(2)}}\{c_{\Delta,k}\}^{n_{\Delta}}_{k=2}$. Then for any $\Delta\in\mathcal{T}^{(2)}$,
$T_{\Delta}\mathrel{\mathop{\ordinarycolon}}=T\cap\mathrm{Pri}_{\Delta,\varphi}$
is a surface with boundary; the boundary consists of smooth short loops $\{c_{\Delta,k}\}^{n_{\Delta}}_{k=1}$ and a piecewise smooth long loop $c_{\Delta,1}$. We glue disks $\{D_{\Delta,k}\}^{n_{\Delta}}_{k=2}$ of arbitrary small areas to its boundary and obtain a orientable surface $T^{\prime}_{\Delta}$ with boundary. If $\mathfrak{g}(T^{\prime}_{\Delta})=1$, then we find a nontrivial simple loop $\gamma$ of $T$ contained in $\mathrm{Pri}_{\Delta,\varphi}\subset B(p,\varepsilon/2)$ for some $p\in M$, so we are done. Otherwise, each $T^{\prime}_{\Delta}$ is a union of multiple spheres and a disk whose boundary is $c_{\Delta,1}$. So by Step 1 (1) and (3), the whole surface $T^{\prime}$ is a genus 0 surface. This implies that at least one short loop $\gamma$ is a non-trivial simple loop in $T$, and by Step 2 (4), it is contained in a small ball $B(p,\varepsilon/2)$.
Step 3. General case
By choosing $\delta$ sufficiently small, for a general torus $T$ with $\mathbf{F}([T],[S])<\delta$, by the coarea formula, we can find $d\in(\zeta/2,\zeta)$ such that we can cut $T\pitchfork\partial N_{d}(S)=\bigcup^{n_{T}}_{j=1}c_{j}$ and attach disks $\{D_{j}\}^{n_{T}}_{j=1}$ to obtain two disjoint orientable closed surfaces $T^{1}$ and $T^{2}$ satisfying:
  1. (3.i)
    $T^{1}\subset N_{\zeta}(S)$.
  2. (3.ii)
    $\mathbf{F}([T^{1}],[S])<\delta^{\prime}$.
  3. (3.iii)
    $\mathcal{H}^{2}(T^{2})<\varepsilon^{2}/100C$, where $C$ is the universal constant from Theorem 8.2.
  4. (3.iv)
    $\sum^{n_{T}}_{j=1}\mathcal{H}^{1}(c_{j})<\varepsilon/100$.
If $\mathfrak{g}(T^{1})=1$, we can apply Step 2 to $T^{1}$ to obtain a nontrivial simple loop $\gamma^{1}\in T^{1}$ and a small ball $B(p,\varepsilon/2)\supset\gamma^{1}$.
If $\mathfrak{g}(T^{2})=1$, then by Theorem 8.2, we can find a nontrivial simple loop $\gamma^{2}\in T^{2}$ of length $\leq\varepsilon/10$. Therefore, there exists a geodesic ball $B(p,\varepsilon/2)\supset\gamma^{2}$.
In either case, $\gamma^{1}$ or $\gamma^{2}$ might intersect with some $D_{j}$’s. Since $D_{j}$’s are contractible, we can push $\gamma^{1}$ or $\gamma^{2}$ to the boundary, and obtain a nontrivial simple loop $\gamma\subset T$. In addition, $\gamma\subset B(p,\varepsilon)$ by the length estimates of $\mathcal{H}^{1}(c_{j})$.
Finally, if $\mathfrak{g}(T^{1})=\mathfrak{g}(T^{2})=0$, then one of $c_{j}$ is a nontrivial simple loop in $T$ and the conclusion follows. ∎
To prove the existence of a nontrivial loop for punctate surfaces, we first need to approximate them by smooth surfaces as follows.
Lemma 8.4.
In an orientable closed Riemannian 333-manifold $(M,g)$, suppose that $S$ is an embedded smooth orientable closed surface. For any $q\in\mathbb{N}^{+}$ and $\varepsilon\in\left(0,\operatorname{injrad}(M,\mathbf{g})/2\right)$, there exists a constant $\delta=\delta_{1}(S,q,\varepsilon)\in(0,\varepsilon)$ with the following property.
For any punctate surface $\Sigma\in\mathcal{S}(M)$ with a punctate set $P$ such that
#$\#P\leq q\,,\quad\mathbf{F}([S],[\Sigma])<\delta\,,$
[…]
  1. (1)
    The closed balls $\{\overline{B(p_{i},r)}\}^{\widetilde{q}}_{i=1}$ are pairwise disjoint geodesic balls.
  2. (2)
    $\Sigma\setminus\bigcup^{\widetilde{q}}_{i=1}B(p_{i},r)$ is an embedded smooth surface with boundary and genus at most $\mathfrak{g}(\Sigma)$.
  3. (3)
    By replacing $\Sigma\cap\bigcup^{\widetilde{q}}_{i=1}B(p_{i},r)$ by finitely many disks $\{D_{j}\}$, we obtain an embedded smooth orientable closed surface $T$ satisfying $\mathbf{F}([T],[S])<\varepsilon$ and $\sum^{n}_{j=1}\mathcal{H}^{1}(\partial D_{j})<\varepsilon$.
Proof.
Since $S$ is a smooth closed surface, there exists a positive constant $C_{S}>1$, such that for any $p\in M$ and $r\in(0,\operatorname{injrad}(M,g)/2)$,
$\mathcal{H}^{2}(\overline{B(p,r)}\cap S)\leq C_{S}r^{2}\,,\quad\mathcal{H}^{3}(\overline{B(p,r)})\leq C_{S}r^{3}\,.$
Let $\eta$ be a sufficiently small positive constant determined later.
By the definition of $\mathbf{F}$-metric, there exists $\delta\in(0,\varepsilon/2)$ such that for any $\Sigma\in\mathcal{S}(M)$ with $\mathbf{F}([S],[\Sigma])\leq\delta$ and any $p\in M$,
$\mathcal{H}^{2}(\overline{B(p,2\eta)}\cap\Sigma)\leq 8C_{S}\eta^{2}\,.$
Suppose that the punctate set $P$ has #$\#P\leq q$.
By Sard’s theorem and the coarea formula, we can choose a constant $C_{q}>0$, only depending on $q$, and a radius $\widetilde{r}\in\left(\eta_{3}/3^{q},2\eta_{3}/3^{q}\right)$ such that, for any $p\in P$ and any $l\in\{1,3,3^{2},\dots,3^{q}\}$, $\partial B(p,l\widetilde{r})\pitchfork S$ is a finite union of disjoint simple loops and
$\mathcal{H}^{1}(\partial B(p,l\widetilde{r})\cap\Sigma)\leq C_{q}\mathcal{H}^{2}(\overline{B(p,2\eta)}\cap\Sigma)/\eta<8C_{q}C_{S}\eta\,.$
Then one can find a subset $P^{\prime}\subset P$ and a radius $r\in\{\widetilde{r},3\widetilde{r},3^{2}\widetilde{r},\dots,3^{q}\widetilde{r}\}\subset(0,\eta)$ such that $\Sigma\setminus\bigcup_{p^{\prime}\in P^{\prime}}B(p^{\prime},r)$ is a smooth surface with boundary satisfying
$\sum_{p^{\prime}\in P^{\prime}}\mathcal{H}^{1}(\partial B(p^{\prime},r)\cap\Sigma)<8qC_{q}C_{S}\eta\,,$
$P\subset\bigcup_{p^{\prime}\in P^{\prime}}B(p^{\prime},r)$ and $\{\overline{B(p^{\prime},r)}\}_{p^{\prime}\in P^{\prime}}$ are disjoint from each other. Clearly, we also have $\mathfrak{g}(\Gamma)\leq\mathfrak{g}(S)$. These confirms (1) and (2).
By choosing $\eta$ sufficiently small,  (3) follows from the isoperimetric inequality. ∎
Lemma 8.5 (Existence of small nontrivial loops: Punctate surfaces).
[…]
Suppose that $\Sigma\in\mathcal{S}(M)$ has $\mathfrak{g}(\Sigma)=1$, its punctate set $P$ has #$\#P\leq q$, and there exists a $S\in\mathscr{S}$ such that $\mathbf{F}([\Sigma],[S])\leq\delta$. Then there exists $p\in S^{3}$ and an embedded loop $\gamma\subset\Sigma\cap B(p,\varepsilon)$ such that the surface obtained by neckpinch along $\gamma$ of $\Sigma$ has genus 0.
Proof.
For every $S\in\mathscr{S}$, let $\delta_{0}(S)\mathrel{\mathop{\ordinarycolon}}=\delta_{0}(S,\varepsilon/2)$ be as defined in Lemma 8.3 and $\delta_{1}(S)\mathrel{\mathop{\ordinarycolon}}=\delta_{1}\left(S_{i},q,\delta_{0}(S)/2\right)$ as defined in Lemma 8.4. The compactness of $\mathscr{S}$ implies that there exists finite subset $\{S_{i}\}$ such that for every $S^{\prime}\in\mathscr{S}$, there exists $S_{i_{0}}$ with
$\mathbf{F}([S^{\prime}],[S_{i_{0}}])\leq\delta_{1}(S_{i_{0}})/2\,.$
We define
$\delta\mathrel{\mathop{\ordinarycolon}}=\min_{i}\delta_{1}\left(S_{i}\right).$
For every $\Sigma\in\mathcal{S}(M)$ with $\mathfrak{g}(\Sigma)=1$, if its punctate set $P$ has #$\#P\leq q$, and there exists a $S\in\mathscr{S}$ such that $\mathbf{F}([\Sigma],[S])\leq\delta$, then there exists $S_{i_{0}}$ such that
$\mathbf{F}([\Sigma],[S_{i_{0}}])\leq\delta_{1}(S_{i_{0}})\,.$
We can apply Lemma 8.4 to approximate $\Sigma$ by an embedded closed surface $T$ where
$T\setminus\bigcup B(p_{i},r)=\Sigma\setminus\bigcup B(p_{i},r)$
for a disjoint union of balls and $r\in(0,\varepsilon/2)$ and $\mathbf{F}([T],[S_{i_{0}}])<\delta_{0}(S_{i_{0}})$.
Since $\mathfrak{g}(\Sigma)=1$, $\mathfrak{g}(T)=0$ or 111:
  • If $\mathfrak{g}(T)=0$, then there exists a nontrivial simple loop of $\Sigma$ lying inside some $B(p_{i},r)$.
  • Otherwise, by Lemma 8.3, there exists a nontrivial simple loop $\gamma\subset T$ contained in a ball $B(p,\varepsilon/2)$. As before, $\gamma$ might not lie inside $\Sigma$ but since each disk in $T\cap\bigcup B(p_{i},r)$, we can push $\gamma$ to their boundary. The resulting simple loop $\gamma^{\prime}\subset\Sigma$ and is inside the ball $B(p,\varepsilon)$.
These complete the proof. ∎
8.2. Local deformations
In the previous subsection, we located nontrivial loops for singular tori within small balls, provided that they are sufficiently close to a sphere. One key reason we do not simply pinch along the loops, which resemble mean curvature flow, is the need to construct the pinching in a continuous manner along a Simon-Smith family. The singular set presents a technical obstruction to this straightforward approach. Therefore, we employ a two-step modification process, Pinching and Shrinking, as detailed in the following lemmas.
Lemma 8.6 (Local Deformation A: Pinching).
[…]
Let $\Lambda>0$, $0<r<\mathrm{injrad}(M,g)/4$, $p\in M$. Suppose that $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}^{*}(M)$ is a Simon-Smith family and $x_{0}\in X$ so that $\Phi(x_{0})\in\mathcal{S}(M)$ satisfies
$\mathcal{H}^{2}(\Phi(x_{0})\cap A(p;r,2r))\leq\Lambda r^{2}\,.$
Then there exist $s\in(r,2r)$, $\zeta\in(0,\min\{\Lambda r,s-r,2r-s\}/5)$, an integer $K\geq 0$, a neighborhood $O_{x_{0}}$ of $x_{0}$ in $X$, and a Simon-Smith family
$H\mathrel{\mathop{\ordinarycolon}}[0,1]\times O_{x_{0}}\to\mathcal{S}(M)\,,$
with the following properties.
  1. (1)
    For every $y\in O_{x_{0}}$, $\Sigma_{y}\mathrel{\mathop{\ordinarycolon}}=\Phi(y)\cap A(p;s-5\zeta,s+5\zeta)$ is a surface, varying smoothly in $y$.
  2. (2)
    For every $s^{\prime}\in[s-4\zeta,s+4\zeta]$ and every $y\in O_{x_{0}}$, we have $\Phi(y)$ intersects $\partial B_{s^{\prime}}(p)$ transversally, and
    $\Phi(y)\cap\partial B_{s^{\prime}}(p)=\mathrel{\mathop{\ordinarycolon}}\bigsqcup_{i=1}^{K}\gamma_{i}(s^{\prime},y)\,,\quad\sum_{i=1}^{K}\mathcal{H}^{1}(\gamma_{i}(s^{\prime},y))\leq 2\Lambda r\,.$
If $K=0$, then $H(t,y)\equiv\Phi(y)$ for every $y\in O_{x_{0}}$ and $t\in[0,1]$; and if $K\geq 1$, then
  1. (3)
    For every $y\in O_{x_{0}}$, $H(0,y)=\Phi(y)$.
  2. (4)
    There exists $t_{0}=0<t_{1}<t_{2}<\dots<t_{K}<1=t_{K+1}$ such that for every $0\leq i\leq K$, $t\in(t_{i},t_{i+1})$ and $y\in O_{x_{0}}$,
    $\Sigma_{t,y}\mathrel{\mathop{\ordinarycolon}}=H(t,y)\cap A(p,s-5\zeta,s+5\zeta)$
[…]
    $H(1,y)\cap A(p,s-\zeta,s+\zeta)=\emptyset\,.$
  3. (5)
    For every $i\in\{1,2,\cdots,K\}$ and $y\in O_{x_{0}}$, near $t_{i}$, $H(t,y)$ is a surgery process via pinching the cylinder $\bigcup_{s^{\prime}\in(s-2\zeta,s+2\zeta)}\gamma_{i}(s^{\prime},y)$.
  4. (6)
    For every $y\in O_{x_{0}}$ and $t\in[0,1]$, we have
    $H(t,y)\setminus A(p;s-3\zeta,s+3\zeta)$ $=\Phi(y)\setminus A(p;s-3\zeta,s+3\zeta)\,,$
    $\mathcal{H}^{2}(H(t,y))$ $\leq\mathcal{H}^{2}(\Phi(y))+C_{g}\Lambda^{2}r^{2}\,.$
Remark 8.7.
For every $y\in O_{x_{0}}$, by (5), we see that $\{H(t,y)\}_{t\in[0,1]}$ is a pinch-off process.
Proof.
Since $\mathcal{H}^{2}(\Phi(x_{0}))\leq\Lambda r^{2}$ and $\Phi(x_{0})\setminus P$ is a smooth surface for some finite set $P$, by the slicing theorem and Sard’s lemma, there exists $s\in(r,2r)$ such that $\Phi(x_{0})\pitchfork\partial B_{r}(p)$ is a finite union of loops $\{\gamma_{i}(s,x_{0})\}^{K}_{i=1}$ with
$\sum_{i=1}^{K}\mathcal{H}^{1}(\gamma_{i}(s,x_{0}))\leq\Lambda r\,.$
By Definition 2.4 (4), we can choose $\zeta$ and $O_{x_{0}}$ such that (1) and (2) hold. Moreover, one can find a continuous family of $\{\varphi_{y}\}_{y\in O_{x_{0}}}\subset\operatorname{Diff}^{\infty}(M)$ such that
$\varphi_{y}(\Phi(x_{0}))\cap A(p;s-5\zeta,s+5\zeta)=\Sigma_{y}\,.$
Therefore, by possibly choosing a smaller $O_{x_{0}}$, it suffices to find a continuous map $\{H(t,x_{0})\}_{t\in[0,1]}$ for $\Sigma$ supported in $A(p;s-4\zeta,s+4\zeta)$ satisfying (3) and
  1. (4)’
    There exists $t_{0}=0<t_{1}<t_{2}<\dots<t_{K}<1=t_{K+1}$ such that for every $0\leq i\leq K$, $t\in(t_{i},t_{i+1})$ and $x_{0}\in O_{x_{0}}$,
    $\Sigma_{x_{0},t}\mathrel{\mathop{\ordinarycolon}}=H(t,x_{0})\cap A(p,s-5\zeta,s+5\zeta)$
[…]
    $H(1,x_{0})\cap A(p,s-1.1\zeta,s+1.1\zeta)=\emptyset\,.$
  2. (5)’
    For every $i\in\{1,2,\cdots,K\}$, near $t_{i}$, $H(t,x_{0})$ is a surgery process via pinching the cylinder $\bigcup_{s^{\prime}\in(s-1.9\zeta,s+1.9\zeta)}\gamma_{i}(s^{\prime},x_{0})$;
  3. (6)’
    For every $t\in[0,1]$, we have
    $H(t,x_{0})\setminus A(p;s-3.1\zeta,s+3.1\zeta)$ $=\Phi({x_{0}})\setminus A(p;s-3.1\zeta,s+3.1\zeta)\,;$
    $\mathcal{H}^{2}(H(t,x_{0}))$ $\leq\mathcal{H}^{2}(\Phi({x_{0}}))+C_{g}\Lambda^{2}r^{2}/1.1\,.$
Indeed, the general $H\mathrel{\mathop{\ordinarycolon}}[0,1]\times O_{x}\to\mathcal{S}(M)$ can be defined by
$H(t,y)\mathrel{\mathop{\ordinarycolon}}=\left(\varphi_{y}(H(t,x_{0}))\cap A(p;s-5\zeta,s+5\zeta)\right)\cup\left(\Phi(y)\setminus A(p;s-5\zeta,s+5\zeta)\right)\,,$
and one can verify that $H$ satisfies (3) - (6) following the properties of $H(\cdot,x_{0})$.
Note that in $(M,g)$, for every $r\in(0,\mathrm{injrad}(M,g)/4)$ and $p\in S$, the annulus $A(p;r,2r)$ is diffeomorphic to $\mathbb{A}(\mathbf{0};r,2r)\subset\mathbb{R}^{3}$ with a uniform bi-Lipschitz constant $C^{(1)}_{g}$ depending only on $g$. Without loss of generality, we may assume that $A(p;r,2r)$ is $\mathbb{A}(\mathbf{0};r,2r)$.
By the isoperimetric inequality in standard spheres, there exists a dimensional constant $C^{(2)}>0$ such that for any loop $\gamma$ in $\gamma$ in a sphere $\partial\mathbb{B}_{s}$ for any $s>0$, $\gamma$ bounds a disk $D$ in $\partial\mathbb{B}_{s}$ with
$\mathcal{H}^{2}(D)\leq C^{(2)}\mathcal{H}^{1}(\gamma)^{2}\,.$
Therefore, together with Jordan curve theorem, given a finite set of disjoint loops $\{\gamma_{i}\}$ in $\partial\mathbb{B}_{s}$, each of them bounds a disk $D_{i}$ in $\partial\mathbb{B}_{s}$ such that
$\sum_{i}\mathcal{H}^{2}(D_{i})\leq C^{(2)}\sum_{i}\mathcal{H}^{1}(\gamma_{i})^{2}\leq C^{(2)}(\sum_{i}\mathcal{H}^{1}(\gamma_{i}))^{2}\,.$
In addition, for every pair $i$ and $j$, we have a trichotomy: $D_{i}\cap D_{j}=\emptyset$, $D_{i}\subset D_{j}$ or $D_{j}\subset D_{i}$.
Hence, for $\Phi(x_{0})$ with $\Phi(x_{0})\cap\partial\mathbb{B}_{s}(p)=\bigsqcup^{K}_{i=1}\gamma_{i}$, we can perform the neckpinch surgery starting from the innermost $D_{i}$ to the outermost $D_{i}$ in a the neighborhood $\mathbb{A}(\mathbf{0};s-1.9\zeta,s+1.9\zeta)$ through removing cylinders in $\mathbb{A}(\mathbf{0};s-1.1\zeta,s+1.1\zeta)$ and gluing disks. Since the total area of all the gluing disks is bounded by
$3\sum_{i}\mathcal{H}^{2}(D_{i})+6(\sum_{i}\mathcal{H}^{1}(\gamma_{i}))\zeta\leq C^{(3)}\Lambda^{2}r^{2}\,,$
for some dimensional constant $C^{(3)}$.
In a general annulus $A(p;r,2r)$, we only need to replace $C^{(3)}$ by a constant $C(C^{(1)}_{g},C^{(2)},C^{(3)})$, and thus, the statements (3), (4)’, (5)’ and (6)’ above follow immediately. ∎
Lemma 8.8 (Local Deformation B: Shrinking).
In a 333-dimensional Riemannian manifold $(M,g)$, let $p\in M$ and $0<r^{-}<r^{+}<\mathrm{injrad}(M,g)/4$. Then there exists a smooth one parameter family of maps $\{\mathcal{R}_{t}\mathrel{\mathop{\ordinarycolon}}M\to M\}_{t\in[0,1]}$ such that
  1. (1)
    $\mathcal{R}_{t}=\mathrm{id}$ when $t\in[0,1/4]$; $\mathcal{R}_{t}(B(p,r^{-}))=\{p\}$ when $t\in[3/4,1]$.
  2. (2)
    $\mathcal{R}_{t}$ is a diffeomorphism when $t\in[0,3/4)$.
  3. (3)
    $\mathcal{R}_{t}|_{M\setminus B(p,r^{+})}=\mathrm{id}$ and $\mathcal{R}_{t}(B(p,r^{\pm}))\subset B(p,r^{\pm})$ for every $t\in[0,1]$.
  4. (4)
    $\mathcal{R}_{t}|_{B(p,r^{-})}$ is 111-Lipschitz for every $t\in[0,1]$.
Proof.
Let $\psi\in C^{\infty}(\mathbb{R})$ be a non-decreasing function such that $\psi=0$ on $\mathbb{R}_{\leq 1/4}$, $\psi=1$ on $\mathbb{R}_{\geq 3/4}$, $\psi\in(0,1)$ on (1/4,3/4)(1/4,3/4)( 1 / 4 , 3 / 4 ). Let $\eta\in C^{\infty}_{c}(\mathbb{R}_{<r_{+}})$ be a decreasing function such that $\eta=1$ on $\mathbb{R}_{\leq r_{-}}$.
Working under normal coordinates of $(M,g)$ at $p$, it is easy to check that the following $\mathcal{R}_{t}$ satisfies (1)-(4) in the Lemma:
$\mathcal{R}_{t}(q)\mathrel{\mathop{\ordinarycolon}}=\begin{cases}q\,,&\text{ if }q\notin B(p,r^{+})\,,\\ \left(1-\psi(t)\eta(|q|)\right)q\,,&\text{ if }q\in B(p,r^{+})\,.\end{cases}$
Lemma 8.9.
In $(M,g)$, for $N\in\mathbb{N}^{+}$, let $\{A(p_{j};r_{j}^{-},r_{j}^{+})\}_{j=1}^{N}$ be a collection of pairwise disjoint annuli in $M$, where $r_{j}^{+}<\mathrm{injrad}(M,g)/4$. For each $j\in\{1,2,\cdots,N\}$, let $\mathcal{R}^{j}_{t}\mathrel{\mathop{\ordinarycolon}}M\to M$ be a map from Lemma 8.8 with $p_{j},r_{j}^{\pm}$ in place of $p,r^{\pm}$. For simplicity, we denote
$A_{j}\mathrel{\mathop{\ordinarycolon}}=A(p_{j};r^{-}_{j},r^{+}_{j})\,,\quad B^{\pm}_{j}\mathrel{\mathop{\ordinarycolon}}=B(p_{j};r^{\pm}_{j})\,.$
Then the following hold.
  1. (1)
    There exists a permutation $\sigma\in\mathfrak{S}_{N}$ such that for every $1\leq i<j\leq N$, either $B_{\sigma(i)}^{+}\cap B_{\sigma(j)}^{+}=\emptyset$, or $B_{\sigma(i)}^{+}\subset B_{\sigma(j)}^{-}$. In this case, we call such a $\sigma$ admissible.
  2. (2)
    For every admissible $\sigma$, the map
    $\mathcal{R}^{\sigma(N)}_{t}\circ\mathcal{R}^{\sigma(N-1)}_{t}\circ\cdots\circ\mathcal{R}^{\sigma(2)}_{t}\circ\mathcal{R}^{\sigma(1)}_{t}\mathrel{\mathop{\ordinarycolon}}M\to M$
    is 111-Lipschitz on every connect component of $M\setminus\bigcup_{j=1}^{N}A_{j}$; and it is independent of the choice of admissible $\sigma$.
Proof.
Since $\{A_{j}\}$ is pairwise disjoint, for every $i\neq j$, we have exactly one of the three possibilities:
  • $B_{\sigma(i)}^{+}\cap B_{\sigma(j)}^{+}=\emptyset$;
  • $B_{\sigma(i)}^{+}\subset B_{\sigma(j)}^{-}$;
  • $B_{\sigma(j)}^{+}\subset B_{\sigma(i)}^{-}$.
Hence, to achieve (1), we can simply choose a permutation $\sigma$ such that for every $1\leq i<j\leq N$, $r^{+}_{\sigma(i)}\leq r^{+}_{\sigma(j)}$.
For (2), the 111-Lipschitz property follows from Lemma 8.8 (3) and (4). The independence on the choice of admissible permutations follows from the fact that if $B^{+}_{i}\cap B^{+}_{j}=\emptyset$, then $\operatorname{spt}\mathcal{R}^{i}_{t}\cap\operatorname{spt}\mathcal{R}^{j}_{t}=\emptyset$. ∎
8.3. Combinatorial arguments
To extend the local deformations from the previous subsection to a global deformation along a Simon-Smith family, we rely on the following two useful combinatorial arguments, inspired by [Pit81, Lemma 4.8, Proposition 4.9] and [DR18a, Lemma 5.7].
Lemma 8.10 (Combinatorial Argument I).
For any positive integers $m\in\mathbb{N}^{+}$ and $q\in\mathbb{N}^{+}$, there exists $N=N(m,q)\in\mathbb{N}^{+}$ with the following property.
Given any $k\in\mathbb{N}^{+}$ and a cubical complex $I(m,k)$, suppose that $\{\mathcal{F}_{\sigma}\}_{\sigma\in I(m,k)}$ is a family of collections of open sets in $M$ assigned to each cell $\sigma$ of $I(m,k)$, where each collection $\mathcal{F}_{\sigma}$ has the form
$\mathcal{F}_{\sigma}=(\mathcal{O}_{\sigma,1},\mathcal{O}_{\sigma,2},\cdots,\mathcal{O}_{\sigma,q})$
and each $\mathcal{O}_{\sigma,i}=\{U^{1}_{\sigma,i},U^{2}_{\sigma,i},\cdots,U^{N}_{\sigma,i}\}$ satisfies
(8.2) $\operatorname{dist}(U^{r}_{\sigma,i},U^{s}_{\sigma,i})\geq 2\min\{\mathrm{diam}(U^{r}_{\sigma,i}),\mathrm{diam}(U^{s}_{\sigma,i})\}\,,$
for all $\sigma\in I(m,k)$, $i\in\{1,2,\cdots,q\}$ and $r,s\in\{1,2,\cdots,N\}$ with $r\neq s$. Then we can extract a family of open sets
$\{(U_{\sigma,1},U_{\sigma,2},\cdots,U_{\sigma,q})\}_{\sigma\in I(m,k)}$
where $U_{\sigma,i}\in\mathcal{O}_{\sigma,i}$ such that
$\operatorname{dist}(U_{\sigma,i}\cap U_{\tau,j})>0\,,$
whenever $\sigma,\tau\in I(m,k)$, $(\sigma,i)\neq(\tau,j)$ and $\sigma,\tau$ are faces of a common cell $\gamma$ of $I(m,k)$.
Moreover, this is also true if the number of collections in each $\mathcal{F}_{\sigma}$ is no greater than $q$.
Proof.
The proof is similar to that of [DR18a, Lemma 5.7].
Let $\{\mathcal{F}_{\sigma}\}_{\sigma\in I(m,k)}$ be a family as in the lemma with some $N\in\mathbb{N}^{+}$ to be determined later.
Note that by our assumption (8.2), for each $U\in\mathcal{O}_{\sigma,i}$ and $V^{1},\cdots,V^{l}\in\mathcal{O}_{\tau,j}$ with $\sigma\neq\tau$, if $\mathrm{diam}(U)\leq\mathrm{diam}(V^{j})$ for all $j\in\{1,2,\cdots,l\}$, then there exists at most one $V^{j}$ with
$\operatorname{dist}(U,V^{j})=0\,.$
Here is the algorithm to generate the family of open sets:
  1. (1)
    We take an open set $U^{r}_{\sigma,i}\in\mathcal{O}_{\sigma,i}$ with the smallest diameter among all the remaining open sets $U^{\cdot}_{\cdot,\cdot}$ and define $U_{\sigma,i}\mathrel{\mathop{\ordinarycolon}}=U^{r}_{\sigma,i}$, and then remove all the open sets in $\mathcal{O}_{\sigma,i}$.
  2. (2)
    For every $\tau\in I(m,k)$ and every $j\in\{1,2,\cdots,q\}$, where $(\sigma,i)\neq(\tau,j)$ and $\sigma,\tau$ are faces of some cell $\gamma\in I(m,k)$, and we remove all $V\in\mathcal{O}_{\tau,j}$ with $\operatorname{dist}(U_{\sigma,i},V)=0$.
  3. (3)
    Repeat Step (1) and Step (2), until there is no open set left.
In Step (1), we remove all the open sets in $\mathcal{O}_{\sigma,i}$ after defining $U_{\sigma,i}$, so $U_{\sigma,i}$ can be only defined at most once.
In Step (2), by the argument above, before $U_{\tau,j}$ is defined, the size of $\mathcal{O}_{\tau,j}$ will decrease at most by one if some $U_{\sigma,i}$ is defined where $(\sigma,i)\neq(\tau,j)$ and $\sigma,\tau$ are faces of some cell $\gamma\in I(m,k)$. Since there are at most $5^{m}$ many such $\sigma$, the size of $\mathcal{O}_{\tau,j}$ will decrease at most by $5^{m}\cdot q$ in total before $U_{\sigma,i}$ is defined.
Therefore, it suffices to choose
$N\mathrel{\mathop{\ordinarycolon}}=5^{m}\cdot q+1\,,$
and then for every $\sigma\in I(m,k)$ and $i\in\{1,\cdots,q\}$, the open set $U_{\sigma,i}$ with required properties is defined before our algorithm halts.
For the general case where $|\mathcal{F}_{\sigma}|=q^{\prime}<q$, we can replace each $\mathcal{F}_{\sigma}$ by
$(\mathcal{O}_{\sigma,1},\mathcal{O}_{\sigma,2},\cdots,\mathcal{O}_{\sigma,q^{\prime}},\mathcal{O}_{\sigma,q^{\prime}},\cdots,\mathcal{O}_{\sigma,q^{\prime}})\,.$
The same arguments apply immediately. ∎
Lemma 8.11 (Combinatorial argument II).
Given any integers $q,N\in\mathbb{N}^{+}$ and any $R_{0}\in(0,\mathrm{injrad}(M,g)/4)$, if $P$ is a finite set of at most $q$ points in $M$, then there exists a radius $R\in(5^{-2Nq^{2}}R_{0},5^{-2N}R_{0})$ such that for every $p\in P$,
$A(p;R,5^{2N}R)\cap P=\emptyset\,.$
Proof.
For $1\leq l\leq q^{2}$, we define
$s_{l}=5^{-2Nl}R_{0}\,,\quad r_{l}=5^{-2N(l-1)}R_{0}\,.$
For every $p\in P$, since #$\#P\leq q$, there are at most $q-1$ many $l$ such that
$A(p;s_{l},r_{l})\cap P\neq\emptyset\,.$
Since $q(q-1)<q^{2}$, by the pigeon hole principle, there is a $l_{0}$ such that $R\mathrel{\mathop{\ordinarycolon}}=5^{-2Nl_{0}}R_{0}$ satisfies the requirement of the lemma. ∎
8.4. Proof of Proposition 8.1
The proof will proceed in three steps.
Step 1. Set up
Let $C_{g}$ be as in Lemma 8.6 and $N\mathrel{\mathop{\ordinarycolon}}=N(m,q)$ as in Lemma 8.10. Without loss of generality, we assume that
$\varepsilon<5^{-N-1}\min\left\{\mathrm{injrad}(M,g),\frac{1}{10000\cdot 3^{m}qC_{g}}\right\}\,,$
and we set $\varepsilon_{1}\mathrel{\mathop{\ordinarycolon}}=5^{-2N-10}\varepsilon$, and $\varepsilon_{2}\mathrel{\mathop{\ordinarycolon}}=5^{-2Nq^{2}}\varepsilon_{1}$.
Since $\mathscr{S}$ is a compact set, we can choose $\delta_{0}=\delta_{0}(\mathscr{S})>0$ such that for every $\Sigma\in\mathcal{S}^{*}(M)$, if $\mathbf{F}([\Sigma],[\mathscr{S}])<\delta_{0}$, then for every $r\in(\varepsilon_{2},\varepsilon_{1})$,
(8.3) $\sup_{p\in M}\left\{\mathcal{H}^{2}\left(\Sigma\cap A(p,r,2r)\right)\right\}\leq 100r^{2}\,.$
We select $\delta_{1}\mathrel{\mathop{\ordinarycolon}}=\delta_{1}(\mathscr{S},q,\varepsilon_{2})$ from Lemma 8.5, and set
$\delta(\varepsilon,q,m,\mathscr{S})\mathrel{\mathop{\ordinarycolon}}=\min\{\delta_{0},\delta_{1}\}\,.$
For a Simon-Smith family $\Phi$ from the proposition, by Lemma 8.5, we see that for every $x\in X$ with $\mathfrak{g}(\Phi(x))=1$, there exists $p_{x}\in S^{3}$ such that $B_{\varepsilon_{2}}(p_{x})$ contains a non-trivial loop in $\Phi(x)$.
For each $x\in X$, we set
$\widetilde{P}(x)\mathrel{\mathop{\ordinarycolon}}=\begin{cases}P_{\Phi}(x)\cup\{p_{x}\}&\text{if }\mathfrak{g}(\Phi(x))=1\\ P_{\Phi}(x)&\text{if }\mathfrak{g}(\Phi(x))=0\,.\end{cases}$
By Lemma 8.11, since #$\#\widetilde{P}(x)\leq q$, for each $x\in X$, take $r_{x}\in(\varepsilon_{2},5^{-2N}\varepsilon_{1})$ such that for every pair $p,p^{\prime}\in\widetilde{P}(x)$, we have
$p^{\prime}\notin A(p;r_{x},5^{2N}r_{x})\,.$
For each $l\in\{1,2,\cdots,N\}$, we set
$r_{x,l}\mathrel{\mathop{\ordinarycolon}}=5^{2(l-1)}r_{x}\,.$
By (8.3), for every $1\leq l\leq N$, $x\in X$ and $p\in\widetilde{P}(x)$ we have
(8.4) $\mathcal{H}^{2}\left(\Phi(x)\cap A(p;r_{x,l},2r_{x,l})\right)\leq 100r_{x,l}^{2}\,.$
and thus, applying Lemma 8.6 to $\Phi(x)$ in $A(p;r_{x,l},2r_{x,l})$ with $\Lambda=100$, we obtain the following data:
  1. (1)
    $s_{x,l,p}\in(r_{x,l},2r_{x,l})$,
  2. (2)
    $\zeta_{x,l,p}\in(0,\min\{100r_{x,l},s_{x,l}-r_{x,l},2r_{x,l}-s_{x,l}\}/5)$,
  3. (3)
    a neighborhood $O_{x,l,p}\subset X$ of $x$,
  4. (4)
    and a Simon-Smith family
    $H_{x,l,p}\mathrel{\mathop{\ordinarycolon}}[0,1]\times O_{x,l,p}\to\mathcal{S}(M)\,.$
For each $x\in X$, we set a neighborhood of $x$
$O_{x}\subset\bigcap_{l\in\{1,2,\cdots,N\},p\in\widetilde{P}(x)}O_{x,l,p}\,,$
such that $y\mapsto\Phi(y)\setminus\bigcup_{p\in\widetilde{P}(x)}B(p;r_{x,1,p},2r_{x,1,p})$ is continuous in the smooth topology for $y\in O_{x}$.
Note that by the definition of $\widetilde{P}(x)$, $\mathfrak{g}(\Phi(x)\setminus\bigcup_{p\in\widetilde{P}(x)}B(p;r_{x,1,p}))=0$ and for every $y\in O_{x}$, we also have $\mathfrak{g}(\Phi(y)\setminus\bigcup_{p\in\widetilde{P}(x)}B(p;r_{x,1,p}))=0$. Therefore, if $\mathfrak{g}(\Phi(y))=1$, then there exists at least one nontrivial loop contained in $\Phi(y)\cap\bigcup_{p\in\widetilde{P}(x)}B(p;r_{x,1,p})$.
Step 2. Refinement
Since $X$ is compact, we can take a finite cover from $\{O_{x}\}_{x\in X}$. Then we refine $X$ so that for every cell $\sigma$ of $X$, we can find an $x_{\sigma}\in X$ satisfying that every cell $\tau$ of $X$ is a subset of $O_{x_{\sigma}}$ provided that $\sigma$ and $\tau$ are faces of some cell $\gamma$ of $X$. Note that for each $\sigma$, the number of such $\tau$ is no more than $5^{m}$.
By applying Lemma 8.10 with
$\mathcal{F}_{\sigma}=(\{A(p;r_{x_{\sigma},i},2r_{x_{\sigma},i})\}^{N}_{i=1})_{p\in\widetilde{P}(x_{\sigma})}\,,$
we obtain, associated with each $\sigma$, a collection of annuli denoted by
$(A_{\sigma,p})_{p\in\widetilde{P}(x_{\sigma})}$
such that whenever $\sigma$ and $\tau$ are faces of some cell $\gamma$ of $X$ and for any $p\in\widetilde{P}(x_{\sigma})$ and $p^{\prime}\in\widetilde{P}(x_{\tau})$, we have
$A_{\sigma,p}\cap A_{\tau,p^{\prime}}=\emptyset\,,$
unless $(\sigma,p)=(\tau,p^{\prime})$.
For each cell $\sigma$ and $p\in\widetilde{P}(x_{\sigma})$, for convenience, we denote the data associated with each $A_{\sigma,p}$, constructed from Lemma 8.6 at the end of Step 1, as follows:
  1. (1)
    $s_{\sigma,p}\in(r_{\sigma,p},2r_{\sigma,p})$,
  2. (2)
    $\zeta_{\sigma,p}\in(0,\min\{s_{\sigma,p}-r_{\sigma,p},2r_{\sigma,p}-s_{\sigma,p}\}/5)$,
  3. (3)
    a neighborhood $O_{\sigma}\mathrel{\mathop{\ordinarycolon}}=O_{x_{\sigma}}$,
  4. (4)
    and a Simon-Smith family
    $H_{\sigma,p}\mathrel{\mathop{\ordinarycolon}}[0,1]\times O_{\sigma}\to\mathcal{S}(M)\,.$
We also denote
$r_{\sigma,p}^{\pm}\mathrel{\mathop{\ordinarycolon}}=s_{\sigma,p}\pm\zeta_{\sigma,p}\,,\quad B_{\sigma,p}^{\pm}\mathrel{\mathop{\ordinarycolon}}=B_{r_{\sigma,p}^{\pm}}(p_{\sigma,p})\,,\quad\hat{A}_{\sigma,p}\mathrel{\mathop{\ordinarycolon}}=\overline{B_{\sigma,p}^{+}\setminus B_{\sigma,p}^{-}}\subset A_{\sigma,p}\,.$
Step 3. Construction of $H$
After the refinement in the previous step, suppose that $X$ is a cubical subcomplex of $I(m,k^{\prime})$. For each cell $\sigma$ of $X$ and for each $x\in X$, we define
$\mathbf{d}_{\sigma}(x)\mathrel{\mathop{\ordinarycolon}}=\min\left\{\frac{2\|x-c_{\sigma}\|_{\ell_{\infty}}}{3^{-k^{\prime}}},1\right\}\,.$
to be the normalized $l^{\infty}$ distant function to $\sigma$. Here, $c_{\sigma}$ is the center of $\sigma$.
For each $t\in[0,1/2]$ and each $x\in X$, we define $H(x,t)$ using pinching surgeries as described in Lemma 8.6:
For every cell $\sigma$ of $X$ and every $p\in\widetilde{P}(x_{\sigma})$,
$H(t,x)\cap A_{\sigma,p}=H_{\sigma,p}\left(\min\{8t(1-\mathbf{d}_{\sigma}(x)),1\},x\right)\cap A_{\sigma,p}$
and that $H(t,x)=\Phi(x)$ outside $\bigcup_{\sigma\in X,p\in\widetilde{P}(x_{\sigma})}A_{\sigma,p}$. The map has the following properties for every $x\in X$.
  • The set
    $Z_{x}\mathrel{\mathop{\ordinarycolon}}=\{\sigma\mathrel{\mathop{\ordinarycolon}}\mathbf{d}_{\sigma}(x)<1\}$
    has the property that every cell $\sigma$ in $Z_{x}$ is a face of the cell $\gamma$, which is the smallest cell containing $x$. By Step 2, the corresponding annuli $\{A_{\sigma,p}\}_{\sigma\in Z_{x},p\in\widetilde{P}(x_{\sigma})}$ are pairwise disjoint, and the number is no greater than $3^{m}\cdot q$. In particular, the map $H(x,t)$ is well-defined.
  • By Lemma 8.6 (4) and (5), $t\mapsto\mathfrak{g}(H(x,t))$ is non-increasing and $t\mapsto H(x,t)$ is a pinch-off process.
  • By (8.4) and Lemma 8.6 (6), we have for every $t\in[0,1/2]$,
    $\mathcal{H}^{2}(H(t,x))$ $\leq\mathcal{H}^{2}(\Phi(x))+(3^{m}\cdot q)\cdot C_{g}100^{2}\varepsilon^{2}_{1}$
    $\leq\mathcal{H}^{2}(\Phi(x))+\varepsilon\,.$
  • If $\mathbf{d}_{\sigma}(x)\leq 3/4$, then for each $p\in\widetilde{P}(x_{\sigma})$, in $A_{\sigma,p}$, $H(1/2,x)=H_{\sigma,p}(1,x)$, thus by Lemma 8.6 (4), we know that
    $H(1/2,x)\cap\hat{A}_{\sigma,p}=\emptyset\,.$
  • If $\mathfrak{g}(\Phi(x))=1$ and $x\in\sigma$, then by the discussion at the end of Step 1, we know that for some $p\in\widetilde{P}(x_{\sigma})$, $B_{s_{\sigma,p}}(p)$ contains a nontrivial loop of $\Phi(x)$. Thus, the components of $H(1/2,x)$ outside $B_{\sigma}^{+}$ always has genus 0.
Finally, for each $t\in[1/2,1]$ and each $x\in X$, we define $H(x,t)$ using shrinking process as described in Lemma 8.8:
For each cell $\sigma$ of $X$ and $p\in\widetilde{P}(x_{\sigma})$, we let $\mathcal{R}_{(\sigma,p),t}\mathrel{\mathop{\ordinarycolon}}S^{3}\to S^{3}$ be the one-parameter family of shrinking deformation from Lemma 8.8 with $r^{-}=r^{-}_{\sigma,p}$ and $r^{+}=r^{+}_{\sigma,p}$.
Then for each $x\in X$, we label $\{(\sigma,p)\mid\sigma\in Z_{x},\ p\in P^{\prime}(x_{\sigma})\}$ as
$\{(\sigma_{i},p_{i})\}_{1\leq i\leq N^{\prime}}$
such that for every $1\leq i<j\leq N^{\prime}$, either $B_{\sigma_{i},p_{i}}^{+}\cap B_{\sigma_{j},p_{j}}^{+}=\emptyset$, or $B_{\sigma_{i},p_{i}}^{+}\subset B_{\sigma_{j},p_{j}}^{-}$. The existence of such labeling follows from Lemma 8.9 (1). Then, we denote for simplicity,
^$\hat{\mathcal{R}}^{(i)}_{t,x}\mathrel{\mathop{\ordinarycolon}}=\mathcal{R}_{(\sigma_{i},p_{i}),(2t-1)(1-\mathbf{d}_{\sigma}(x))}\,.$
Note that
  • By Lemma 8.8 (1), if $\mathbf{d}_{\sigma_{i}}(x)\geq 3/4$ or $t\in[1/2,5/8]$, then ^$\hat{\mathcal{R}}^{i}_{t,x}=\mathrm{id}$; when $\mathbf{d}_{\sigma_{i}}(x)<3/4$, by the construction above, $H(1/2,x)\cap\hat{A}_{\sigma_{i},p_{i}}=\emptyset$.
  • if $\mathbf{d}_{\sigma_{i}}(x)\leq 1/4$, then ^$\hat{\mathcal{R}}^{(i)}_{1,x}(B_{\sigma_{i},p_{i}}^{-})\subset\{p_{i}\}$.
Now for $t\in[1/2,1]$ and $x\in X$, we define
$H(t,x)\mathrel{\mathop{\ordinarycolon}}=\hat{\mathcal{R}}^{(N^{\prime})}_{t,x}\circ\cdots\circ\hat{\mathcal{R}}^{(2)}_{t,x}\circ\hat{\mathcal{R}}^{(1)}_{t,x}\left(H(x,1/2)\right)\,.$
Intuitively, $\{H(t,x)\}_{t\in[1/2,1]}$ is obtained by shrinking some connected components of $H(1/2,x)$ to points. Therefore, for every $x\in X$, $t\mapsto\mathfrak{g}(H(t,x))$ is non-increasing, and $\{H(t,x\}\}_{t\in[1/2,1]}$ is a pinch-off process. This confirms statement (3) of the proposition.
By Lemma 8.8 (3), (4) and Lemma 8.9 (2), we see that $\mathcal{H}^{2}(H(t,x))$ is non-increasing for $t\in[1/2,1]$. In particular, for every $t\in[1/2,1]$,
$\mathcal{H}^{2}(H(t,x))\leq\mathcal{H}^{2}(\Phi(x))+\varepsilon\,.$
This confirms statement (2) of the proposition.
Furthermore, if $\mathfrak{g}(H(1/2,x))=1$ and $x\in\sigma$, then by the construction above, $H(1,x)$ consists of some connected components of $H(1/2,x)$ outside $\bigcup_{p\in\widetilde{P}(x_{\sigma})}B_{s_{\sigma,p}}(p)$, along with a finite set of points. From the final property of $H(1/2,x)$, it follows that $\mathfrak{g}(H(1,x))=0$. This verifies statement (1) of the proposition and completes the proof.
8.5. Proof of Theorem 3.6
Now we use Proposition 8.1 and the following local min-max theorem to prove Theorem 3.6.
Theorem 8.12 (Local min-max theorem [MN21, Theorem 6.1]).
[…]
  1. (1)
    $F_{0}=\operatorname{Id}$, $F_{v}=F^{-1}_{v}$ for all $v\in\overline{\mathbb{B}}^{k}$.
  2. (2)
    $\|F_{v}-\operatorname{Id}\|_{C^{1}}<\beta$ for all $v\in\overline{\mathbb{B}}^{k}$.
  3. (3)
    The function
    $A^{\Sigma}\mathrel{\mathop{\ordinarycolon}}\overline{\mathbb{B}}^{k}\to[0,\infty],v\mapsto\mathcal{H}^{2}((F_{v})_{\#}\Sigma)\,,$
    is strictly concave.
  4. (4)
    For every $T\in\mathcal{Z}_{2}(M;\mathbb{Z}_{2})$ with $\mathcal{F}(T,[\Sigma])<\varepsilon_{0}$, we have
    $\max_{v\in\overline{\mathbb{B}}^{k}}\mathbf{M}((F_{v})_{\#}T)\geq\mathcal{H}^{2}(\Sigma)$
    with equality only if $[\Sigma]=(F_{v})_{\#}T$ for some $v\in\overline{\mathbb{B}}^{k}$.
We proceed the proof in three steps.
Step 1. Set up
Since $\mathcal{W}_{L,\leq\mathfrak{g}_{0}}(M,g)$ consists of a varifold associated with a non-degenerate multiplicity-one minimal sphere $S$, let $\{F_{v}\}_{v\in\overline{\mathbb{B}}^{k}}$ be the smooth family associated with $S$ given by Theorem 8.12 such that
  • (i)
    For every $T_{1},T_{2}\in\mathcal{Z}_{n}(M;\mathbb{Z}_{2})$ and $v\in\overline{\mathbb{B}}^{k}$,
    $\mathbf{F}((F_{v})_{\#}(T_{1}),(F_{v})_{\#}(T_{2}))\leq 2\mathbf{F}((T_{1}),(T_{2}))\,.$
We can choose smaller $\varepsilon_{0}>0$ such that
  • (ii)
    For every $T\in\mathcal{Z}_{2}(M;\mathbb{Z}_{2})$ with $\mathbf{F}(T,[S])<\varepsilon_{0}$, the function
    $A^{T}\mathrel{\mathop{\ordinarycolon}}\overline{\mathbb{B}}^{k}\to[0,\infty],v\mapsto\mathbf{M}((F_{v})_{\#}T)\,,$
    is strictly concave with a unique maximum in $\mathbb{B}^{k}_{1/2}$.
  • (iii)
    There exists $\varepsilon_{1}>0$ such that for every $T\in\mathcal{Z}_{2}(M;\mathbb{Z}_{2})$ with $\mathbf{F}(T,[S])<\varepsilon_{0}$,
    $\varepsilon_{1}<\min_{v\in\partial\mathbb{B}^{k},T}(\mathbf{M}(S)-\mathbf{M}((F_{v})_{\#}S))\,.$
[…]
$\varepsilon_{1}\mathrel{\mathop{\ordinarycolon}}=\min_{v\in\partial\mathbb{B}^{k}}(\mathbf{M}(S)-\mathbf{M}((F_{v})_{\#}S))>0\,.$
Let $\delta(\varepsilon_{1}/10,N_{P}(\Phi)+1,m+1,\mathscr{S})$ be chosen as in Proposition 8.1.
Step 2. Initial Simon-Smith family
It follows from Theorem 3.5 with $r=\min(\varepsilon_{0},\delta(\varepsilon_{1}/10,\mathcal{S},m)/10)$ that there exists $\eta>0$ and $\Phi_{1}\in\Lambda(\Phi)$ such that
$\mathbf{M}(\Phi_{1}(x))\geq L-4\eta\implies[\Phi_{1}(x)]\in\mathbf{B}^{\mathbf{F}}_{r}([S])\,.$
In particular, there exists a Simon-Smith family
$H_{1}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}^{*}(M)$
with $H_{1}(0,\cdot)=\Phi$ and $H_{1}(1,\cdot)=\Phi_{1}$, which is a pinch-off process, because for every $x\in X$, $H_{1}(\cdot,x)$ is induced by a one-parameter group of diffeomorphisms.
We can refine the cubical subcomplex $X$ so that, each cell $\sigma$ of $X$, exactly one of the following conditions holds:
  • There exists a point $x_{0}\in\sigma$ such that $\mathbf{M}(\Phi_{1}(x_{0}))\geq L-2\eta$. In this case, for every $x\in\sigma$, $\mathbf{M}(\Phi_{1}(x))\geq L-4\eta$.
  • For every $x\in\sigma$, $\mathbf{M}(\Phi_{1}(x))<L-2\eta$.
Let $X_{0}$ be the smallest cubical subcomplex of $X$ containing all cells that satisfy the first condition, and let $X_{1}$ be the smallest cubical subcomplex containing all cells that satisfy the second condition. Let
$Z\mathrel{\mathop{\ordinarycolon}}=X_{0}\cap X_{1}\,,$
which is also compact.
Clearly for every $z\in Z$,
$L-4\eta\leq\mathcal{H}^{2}(\Phi_{1}(z))\leq L-2\eta\,.$
and thus, $\mathbf{F}([\Phi_{1}(z)],[S])<r\leq\varepsilon_{0}$. For every $z\in Z$, let $A^{z}\mathrel{\mathop{\ordinarycolon}}\overline{\mathbb{B}}^{k}\to[0,\infty)$ be the function
$A^{z}(v)\mathrel{\mathop{\ordinarycolon}}=\mathbf{M}((F_{v})_{\#}([\Phi_{1}(z)]))\,,$
By (ii) of Step 1, $A^{z}$ is strictly concave and has a unique maximum $m(z)\in\mathbb{B}^{k}_{1/2}$. By Proposition 2.8, $[\Phi_{1}]$ is continuous in the $\mathbf{F}$-metric, so the function
$m\mathrel{\mathop{\ordinarycolon}}Z\to\mathbb{B}^{k}_{1/2}$
is continuous.
It follows from Theorem 8.12 (4) and $\mathbf{M}([\Phi_{1}(z)])=\mathcal{H}^{2}(\Phi_{1}(z))<L$, that $m(z)\neq 0$ for every $z\in Z$. Hence, there exists $\alpha>0$ such that
$\alpha\leq|m(z)|<1/2\,.$
Consider the one-parameter flow $\{\phi^{z}(\cdot,t)\}_{t\geq 0}\subset\operatorname{Diff}(\overline{\mathbb{B}}^{k})$ generated by
$v\mapsto-(1-|v|^{2})\nabla A^{z}(v)\,.$
For every $v\in\overline{\mathbb{B}}^{k}\setminus m(y)$, $t\mapsto A^{z}(\phi^{z}(v,t))$ is decreasing, and the limit $\lim_{t\to\infty}\varphi^{z}(v,t)\in\partial\mathbb{B}^{k}$. In particular, by (iii) of Step 1, we have
$A^{z}(0)-\lim_{t\to\infty}A^{z}(\phi^{z}(v,t))\geq\varepsilon_{1}\,.$
By the compactness of $Z$, we can choose $T_{0}>0$ such that for every $z\in Z$, $t\mapsto A^{z}(\phi^{z}(v,t))$ is decreasing along $[0,T_{0}]$ and
$A^{z}(0)-A^{z}(\phi^{z}(v,T_{0}))\geq\varepsilon_{1}/2\,.$
Now, let us consider a cubical complex $X^{\prime}$ in $I^{m+1}$, whose underlying space is
$X_{0}\times\{0\}\cup Z\times[0,1]\cup X_{1}\times\{1\}\,.$
From the construction, we also know that the map
$f\mathrel{\mathop{\ordinarycolon}}X^{\prime}\to X,\quad(x,y)\mapsto x$
is a homotopy equivalence. Note that $f$ is a surjective cubical map, after refining both $X$ and $X^{\prime}$.
We can define a Simon-Smith family $\Phi^{\prime}_{2}\mathrel{\mathop{\ordinarycolon}}X^{\prime}\to\mathcal{S}^{*}(M)$ by
$\Phi^{\prime}_{2}(x,0)$ $=\Phi_{1}(x)\,,\quad\forall x\in X_{0}\,,$
$\Phi^{\prime}_{2}(x,1)$ $=\Phi_{1}(x)\,,\quad\forall x\in X_{1}\,,$
$\Phi^{\prime}_{2}(x,y)$ $=F_{\phi^{z}(0,3yT_{0})}(\Phi^{\prime}(x))\,,\quad\forall x\in Z,y\in[0,1/3]\,,$
$\Phi^{\prime}_{2}(x,y)$ $=F_{\phi^{z}(0,T_{0})}(\Phi^{\prime}(x))\,,\quad\forall x\in Z,y\in[1/3,2/3]\,,$
$\Phi^{\prime}_{2}(x,y)$ $=F_{\phi^{z}(0,3(1-y)T_{0})}(\Phi^{\prime}(x))\,,\quad\forall x\in Z,y\in[2/3,1]\,.$
Note that we can define two Simon-Smith family $\Phi^{\prime}\mathrel{\mathop{\ordinarycolon}}X^{\prime}\to\mathcal{S}^{*}(M)$ and $\Phi^{\prime}_{1}\mathrel{\mathop{\ordinarycolon}}X^{\prime}\to\mathcal{S}^{*}(M)$ by extending $\Phi$ and $\Phi_{1}$ to the domain $X^{\prime}$:
$\Phi^{\prime}(x,y)=\Phi(x)\,,\quad\forall(x,y)\in X^{\prime}\,,$
and
$\Phi^{\prime}_{1}(x,y)=\Phi_{1}(x)\,,\quad\forall(x,y)\in X^{\prime}\,,$
Similarly, we also have two deformations $H^{\prime}_{1}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X^{\prime}\to\mathcal{S}^{*}(M)$ from $\Phi^{\prime}$ to $\Phi^{\prime}_{1}$ and $H^{\prime}_{2}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X^{\prime}\to\mathcal{S}^{*}(M)$ from $\Phi^{\prime}_{1}$ to $\Phi^{\prime}_{2}$:
$H^{\prime}_{1}(t,x,y)=H_{1}(t,x)\,,\quad\forall t\in[0,1],(x,y)\in X^{\prime}\,,$
and
$H^{\prime}_{2}(t,x,0)$ $=\Phi_{1}(x)\,,\quad\forall t\in[0,1],x\in X_{0}\,,$
$H^{\prime}_{2}(t,x,1)$ $=\Phi_{1}(x)\,,\quad\forall t\in[0,1],x\in X_{1}\,,$
$H^{\prime}_{2}(t,x,y)$ $=F_{\phi^{z}(0,3tyT_{0})}(\Phi^{\prime}(x))\,,\quad\forall t\in[0,1],x\in Z,y\in[0,1/3]$
$H^{\prime}_{2}(t,x,y)$ $=F_{\phi^{z}(0,tT_{0})}(\Phi^{\prime}(x))\,,\quad\forall t\in[0,1],x\in Z,y\in[1/3,2/3]\,,$
$H^{\prime}_{2}(t,x,y)$ $=F_{\phi^{z}(0,3t(1-y)T_{0})}(\Phi^{\prime}(x))\,,\quad\forall t\in[0,1],x\in Z,y\in[2/3,1]\,.$
For any $i\in{1,2}$ and any $(x,y)\in X^{\prime}$, $\{H^{\prime}_{i}(t,x,y)\}_{t\in[0,1]}$ is induced by an isotopy, and thus, a pinch-off process. Furthermore, we have
(8.5) $N_{P}(\Phi)=N_{P}(\Phi_{1})=N_{P}(\Phi^{\prime})=N_{P}(\Phi^{\prime}_{1})=N_{P}(\Phi^{\prime}_{2})\,.$
Furthermore, $\Phi^{\prime}_{2}$ has the following properties: For every $(x,y)\in X^{\prime}$ with $y\geq 1/3$,
(8.6) $\mathcal{H}^{2}(\Phi^{\prime}_{2}(x,y))\leq L-2\eta\,.$
For every $(x,y)\in X^{\prime}$ with $y\leq 2/3$, by (i) of Step 1,
(8.7) $\mathbf{F}([\Phi^{\prime}_{2}(x,y)],[\mathscr{S}])\leq 2r<\delta(\varepsilon_{1}/10,m+1,N_{P}(\Phi^{\prime}_{2})+1,\mathscr{S})\,.$
For every $(x,y)\in X^{\prime}$ with $y\in[1/3,2/3]$,
(8.8) $\mathcal{H}^{2}(\Phi^{\prime}_{2}(x,y))\leq L-2\eta-\varepsilon_{1}/2\,.$
Step 3. Interpolation
Let $Y^{\prime}\mathrel{\mathop{\ordinarycolon}}=\{(x,y)\in X\mathrel{\mathop{\ordinarycolon}}y\leq 1/3\}$. By (8.7), for every $x^{\prime}\in Y^{\prime}$, $\Phi^{\prime}_{2}(x^{\prime})\in\mathcal{S}(M)$, and we can apply Proposition 8.1 to $\Phi^{\prime}_{2}|_{Y^{\prime}}$ and obtain a deformation
$H^{\prime}_{3,Y^{\prime}}\mathrel{\mathop{\ordinarycolon}}[0,1]\times Y^{\prime}\to\mathcal{S}(M)\,.$
Consequently, we can define a Simon-Smith family $H^{\prime}_{3}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X^{\prime}\to\mathcal{S}^{*}(M)$ as
$H^{\prime}_{3}(t,x,y)$ $=H^{\prime}_{3}(t,x,y)\,,\quad\forall t\in[0,1],(x,y)\in X^{\prime},y\in[0,1/3]\,,$
$H^{\prime}_{3}(t,x,y)$ $=H^{\prime}_{3}((2-3y)t,x,1/3)\,,\quad\forall t\in[0,1],x\in Z,y\in[1/3,2/3]\,,$
$H^{\prime}_{3}(t,x,y)$ $=\Phi^{\prime}_{2}(x,y)\,,\quad\forall t\in[0,1],(x,y)\in X^{\prime},y\in[2/3,1]\,.$
We denote $H^{\prime}_{3}(1,\cdot,\cdot)\mathrel{\mathop{\ordinarycolon}}X^{\prime}\to\mathcal{S}^{*}(M)$ by $\Phi^{\prime}_{3}$.
By Proposition 8.1 (1), for every $(x,y)\in X^{\prime}$ with $y\in[0,1/3]$,
$\mathfrak{g}(\Phi^{\prime}_{3}(x,y))=\mathfrak{g}(H^{\prime}_{3}(1,x,y))=0\,.$
By Proposition 8.1 (2) and (8.8), for every $(x,y)\in X^{\prime}$ with $y\in[1/3,2/3]$,
$\mathcal{H}^{2}(\Phi^{\prime}_{3}(x,y))=\mathcal{H}^{2}(H^{\prime}_{3}((2-3y)t,x,1/3))\leq\mathcal{H}^{2}(\Phi^{\prime}_{2}(x,1/3))+\varepsilon_{1}/10\leq L-2\eta\,.$
By (8.6), for every $(x,y)\in X^{\prime}$ with $y\in[2/3,1]$,
$\mathcal{H}^{2}(\Phi^{\prime}_{3}(x,y))=\mathcal{H}^{2}(\Phi^{\prime}_{2}(x,y))\leq L-2\eta\,.$
Therefore, for every $(x,y)\in X^{\prime}$,
(8.9) $\mathcal{H}^{2}(\Phi^{\prime}_{3}(x,y))\geq L-\eta\implies y\in[0,1/3]\implies\mathfrak{g}(\Phi^{\prime}_{3}(x,y))=0\,.$
Moreover, by Proposition 8.1 (3), for every $(x,y)\in X^{\prime}$, $\{H^{\prime}_{3}(t,x,y)\}_{t\in[0,1]}$ is a pinch-off process.
Finally, we can set $\Phi^{\prime}=\Phi^{\prime}_{3}$, which satisfies (1) of the theorem. Let $W\mathrel{\mathop{\ordinarycolon}}=M_{f}$, and consider the deformation map
$H^{\prime}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X^{\prime}\to\mathcal{S}^{*}(M)\,,$
by
$H^{\prime}(t,x)=\begin{cases}H_{1}(3t,x)&t\in[0,1/3]\,,\\ H_{2}(3t-1,x)&t\in[1/3,2/3]\,,\\ H_{3}(3t-2,x)&t\in[2/3,1]\,.\end{cases}$
Since $H^{\prime}(0,x)=\Phi^{\prime}(x)=\Phi(f(x))$, $H^{\prime}$ induces a Simon-Smith family $H\mathrel{\mathop{\ordinarycolon}}W\to\mathcal{S}^{*}(M)$ of genus 1$\leq$ 1≤ 1 satisfying (2)(a) and (2)(b) of the theorem.
This completes the proof.
9. Pinch-off process and mean curvature flow
In this section, we show that a pinch-off process (Definition 3.7) would possess certain topological properties similar to those of mean curvature flow. These properties will be used in §10.
[…]
$W(t)\mathrel{\mathop{\ordinarycolon}}=\{t\}\times(N\backslash M(t))\subset[0,\infty)\times N\,,$
and
$W[t_{1},t_{2}]\mathrel{\mathop{\ordinarycolon}}=\bigcup_{t\in[t_{1},t_{2}]}W(t)\subset[0,\infty)\times N\,.$
In his work [Whi95], B. White ingeniously applied the avoidance principle to show:
  • For any $T>0$, any loop in $W[0,T]$ is homotopic to some loop in $W[0]$.
  • If $N$ is a complete Riemannian manifold with Ricci curvature bounded from below, then for any $T>0$, the homomorphism
    $H_{n-1}(W(T);\mathbb{Z})\to H_{n-1}(W[0,T];\mathbb{Z})$
    induced by the inclusion map $W(T)\hookrightarrow W[0,T]$ is an monomorphism.
  • The rank of $H_{1}(W(t);\mathbb{Z})$ is non-increasing in $t$.
In the following, we prove a similar theorem regarding pinch-off processes. Given a compact Riemannian 333-manifold $M$, and a pinch-off process $\Phi\mathrel{\mathop{\ordinarycolon}}[0,T]\to\mathcal{S}(M)$, consider the complements of the time-slices,
$W(t)\mathrel{\mathop{\ordinarycolon}}=\{t\}\times(M\backslash S(t))\subset[0,T]\times M\,,$
and
$W[t_{1},t_{2}]\mathrel{\mathop{\ordinarycolon}}=\bigcup_{t\in[t_{1},t_{2}]}W(t)\subset[0,T]\times M\,.$
Proposition 9.1.
If $\Phi\mathrel{\mathop{\ordinarycolon}}[0,T]\to\mathcal{S}(M)$ is a pinch-off process, we have:
  1. (1)
    Any loop in $W(T)$ is homotopic in $W[0,T]$ to a loop in $W(0)$.
  2. (2)
    If a loop is homologically non-trivial in $W(T)$, then it is homotopic in $W[0,T]$ to a homologically non-trivial loop in $W(0)$.
  3. (3)
    By Definition 2.1, for each $t\in[0,T]$, $W(t)$ can be written as the disjoint union of two open sets $\mathrm{in}(\Phi(t))$ and $\mathrm{out}(\Phi(t))$, such that both of their reduced boundaries are $\Phi(t)\setminus\Phi(t)_{\text{iso}}$, and they both vary continuously in $t$ (as Caccioppoli sets). Then,
    $\operatorname{rank}(H_{1}(\mathrm{in}(\Phi(t));\mathbb{Z}))\text{ and }\operatorname{rank}(H_{1}(\mathrm{out}(\Phi(t));\mathbb{Z}))$
    are both non-increasing in $t$.
In the remainder of this section, we will prove the above proposition. We will follow the notation in Definition 3.7. For simplicity, we assume that the singular times $t_{1},\cdots,t_{n}$ are distinct. It would be clear that the following strategy can still be applied when some of the $t_{i}$ coincide, by treating the neighborhood of each $(t_{i},p_{i})$ individually.
Item (1)
Let $\gamma$ be a loop in $W(T)$. In order to homotope $\gamma$ back to some loop in $W(0)$, let us construct a homotopy $\{\gamma_{t}\}_{t\in[0,T]}$ with $\gamma_{T}=\gamma$, and $\gamma_{t}\subset W(t)$ for each $t$, backwardly from time $t=T$ to $t=0$. When we decrease $t$ from $T$ to 0, if $t$ is not equal to any of the $t_{i}$, then near $t$, $\Phi(\cdot)$ is moving by isotopy, so clearly the homotopy $\{\gamma_{t}\}$ can be constructed. So, let us assume $t$ is approaching some $t_{i}$ from above. Suppose $(t_{i},p_{i})$ corresponds to “shrinking some components into $p_{i}$”. In this case, $\Phi(t)$ varies smoothly on some $[t_{i},t_{i}+\epsilon]$, so we can define $\{\gamma_{t}\}$ up to $t=t_{i}$. As $t$ further decreases from $t_{i}$, the reversed shrinking process just produces some connected components, so clearly $\{\gamma_{t}\}$ can exist past $t_{i}$. Finally, the case where $(t_{i},p_{i})$ corresponds to a neck-pinch surgery can be argued similarly (we can by perturbation assume $\gamma_{t_{i}}$ avoids $p_{i}$, and continue to decrease $t$). Thus, the desired homotopy $\{\gamma_{t}\}_{t\in[0,T]}$ can be constructed, and so (1) is true.
Item (2)
To prove (2), considering the homotopy $\{\gamma_{t}\subset W(t)\}_{t\in[0,T]}$ of loops obtained previously, it suffices to show that if $\gamma_{0}$ is homologically trivial in $W(0)$ then $\gamma_{T}$ is homologically trivial in $W(T)$. Suppose $\gamma_{0}$ bounds a 2-chain in $W(0)$. We are going to show that $\gamma_{t}$ bounds some 2-chain $C_{t}$ for each $t$. Note $C_{t}$ need not vary continuously in $t$.
When we increase $t$ from 0 to $T$, if $t$ is not equal to any of the $t_{i}$, then as before, near $t$, $\Phi(\cdot)$ is moving by isotopy, so it is possible to construct the desired homotopy of 2-chains. So let us assume $(t_{i},p_{i})$ is a singularity. Now, we fix a continuous choice of inside region and outside region for $\Phi(t)$, for all $t$. Then there are four cases regarding $(t_{i},p_{i})$:
  1. (a)
    $(t_{i},p_{i})$ is a neck-pinch point, such the solid cylinder region lies in the inside region.
  2. (b)
    $(t_{i},p_{i})$ corresponds to shrinking some component of the inside region into the point $p_{i}$, at time $t_{i}$.
  3. (c)
    $(t_{i},p_{i})$ is a neck-pinch point, such the solid cylinder region lies in the outside region.
  4. (d)
    $(t_{i},p_{i})$ corresponds to shrinking some component of the outside region into the point $p_{i}$, at time $t_{i}$.
Without loss of generality let us assume every $\gamma_{t}$ lies in the inside region. For the general case, we just consider the inside and outside components of $\gamma_{t}$ separately, and argue similarly.
In each of these four cases, let $[s_{1},t_{i}]\times U$ be a spacetime neighborhood within which the pinch-off process takes the form as described above. We can assume that we already have the family $\{C_{t}\}$ defined for $t$ up to $s_{1}$. Our goal is to extend it pass the time $t_{i}$.
Case (b)
By choosing the spacetime neighborhood small enough, we can assume that $[s_{1},t_{i}]\times U$ does not intersect $\gamma_{t}$ for every $t\in[s_{1},T]$. Hence, by the definition of case (b), $C_{s_{1}}\cap U$ is a union of 222-cycles. Removing these 2-cycles from $C_{s_{1}}$, and we can extend the family onto the time interval $[s_{1},t_{i}]$, and after that.
Case (a)
From our construction of the family $\{\gamma_{t}\}_{t\in[0,T]}$, we may, by taking the surgery region $U$ to be sufficiently small, assume that $\gamma_{s_{1}}$ avoids $U$, even though $C_{s_{1}}$ could intersect $U$. Now, the boundary of the cylinder $\Phi(s_{1})\cap U$ consists of two loops. Pick one such loop, and denote by $D$ the disc that it bounds in the sphere $\partial U$. When we intersect the two chain $C_{s_{1}}$ with $D$, we can assume that the cross section consists of finitely many loops $c_{1},\cdots,c_{m}$, with each $c_{j}$ bounding a disc $d_{j}$ within $D$. By gluing two copies of each $d_{j}$ to $C_{s_{1}}$, and slightly opening up a gap between the two $d_{j}$, one can obtain a 2-chain bounded by $\gamma_{s_{1}}$ that avoids $D$. By deforming this 2-chain, we can further assume that it avoids the whole ball $U$. Now, let $C_{s_{1}}$ denote this new 2-chain instead, replacing the old one. Then, we can extend the family $\{C_{t}\}$ pass the surgery time $t_{i}$.
Case (c)
It is clear from the definition of Case (c) that we can construct the desired family $\{C_{t}\}$ pass the time $t_{i}$. Intuitively, this is because the neck-pinch is “outward”, so that the inside region is gaining ground.
Case (d)
Note $C_{s_{1}}$ may intersect $U$. We perform surgery for $C_{s_{1}}$ along every loop component of $C_{s_{1}}\cap\partial U$, so that it avoids the sphere $\partial U$. After discarding everything inside $U$, we still call the new 2-chain $C_{s_{1}}$, replacing the old one. Now we can construct the desired family $\{C_{t}\}$ pass the time $t_{i}$.
Item (3)
Let us just do the case for $\mathrm{in}(\Phi(t))$, as the case for $\mathrm{out}(\Phi(t))$ is similar. Suppose we have loops $\gamma^{1}_{T},\cdots,\gamma^{k}_{T}$ in $\mathrm{in}(\Phi(T))$ such that the elements they induce in $H_{1}(\mathrm{in}(\Phi(T));\mathbb{Z})$ are linearly independent. From the proof of (1), we can construct for each $j=1,\cdots,k$ a homotopy $\{\gamma^{j}_{t}\subset\mathrm{in}(\Phi(t))\}_{t\in[0,T]}$ of loops. To prove (3), it suffices to show that $\gamma^{1}_{0},\cdots,\gamma^{k}_{0}$ are also linear independent in $H_{1}(\mathrm{in}(\Phi(0));\mathbb{Z})$. Suppose by contradiction that $\gamma^{1}_{0}+\cdots+\gamma^{k}_{0}$ bounds some 2-chain $C_{0}\subset\mathrm{in}(\Phi(0))$. Following the proof of (2), we can similarly construct for each $t$ a 2-chain $C_{t}\subset\mathrm{in}(\Phi(t))$ bounded by $\gamma^{1}_{t}+\cdots+\gamma^{k}_{t}$. Putting $t=T$, contradiction arises. Thus (3) is true too.
10. Topology of genus one cap
The goal of this section is to prove Theorem 3.16.
As explained at the end of §3, it suffices to show that for each genus 111 cap $C\subset\mathrm{dmn}(\Xi)$, the map
$i_{*}\mathrel{\mathop{\ordinarycolon}}H_{1}(C;\mathbb{Z}_{2})\to H_{1}(\mathrm{dmn}(\Xi);\mathbb{Z}_{2})$
induced by the inclusion $i\mathrel{\mathop{\ordinarycolon}}C\hookrightarrow\mathrm{dmn}(\Xi)$ is trivial.
In particular, for any fixed loop $c\subset C$, we aim to show that
$[i(c)]=0\in H_{1}(\mathrm{dmn}(\Xi);\mathbb{Z}_{2})$
For simplicity, we may just view $c$ as a subset of $\mathrm{dmn}(\Xi)$, we will show
(10.1) $[c]=0\in H_{1}(\mathrm{dmn}(\Xi);\mathbb{Z}_{2})\,.$
Note that, by the definition of a genus 111 cap, the image
(10.2) $[\Xi](c)\subset\mathbf{B}^{\mathbf{F}}_{d_{0}}([T])$
for some embedded minimal torus $T$.
The proof of (10.1) will consist of three main steps. First, we need to understand better the topology of the members of $\Xi|_{c}$. In the second step, using family $\widetilde{\Xi}$ obtained in Proposition 3.11, we, in a certain sense, homotope the family $\Xi|_{c}$ back to some subfamily $\Phi|_{c_{0}}$ of $\Phi$, for some loop $c_{0}\subset Y$, while keeping track of the topology of the members of this homotopy. We will show that one can assume $\Phi|_{c_{0}}$ to consist entirely Clifford tori (if $\Psi$ is viewed in $\mathbb{S}^{3}$), and to prove (10.1) it suffices to show that $c_{0}$ is homologically trivial in $Y$. In the third step, we prove that $c_{0}$ is homologically trivial by studying the family $\Psi$.
10.1. The family $\Xi|_{c}$
In Proposition 3.11, we obtained a mapping cone $\widetilde{W}$ containing $Y$ and $\mathrm{dmn}(\Xi)$, a map $\widetilde{F}\mathrel{\mathop{\ordinarycolon}}[0,1]\times\widetilde{W}\to\widetilde{W}$, and a Simon-Smith family $\widetilde{\Xi}\mathrel{\mathop{\ordinarycolon}}\widetilde{W}\to\mathcal{S}^{*}(S^{3})$ of genus 1$\leq$ 1≤ 1. Then we can define a map
$G_{1}\mathrel{\mathop{\ordinarycolon}}[0,1]\times S^{1}\to\widetilde{W}$
such that $G_{1}(1,\cdot)$ parametrizes a loop $c\subset\widetilde{W}$, and for every $t\in[0,1]$ and every $\theta\in S^{1}$,
$G_{1}(t,\theta)\mathrel{\mathop{\ordinarycolon}}=\widetilde{F}(t,G_{1}(1,\theta))\,.$
By Proposition 3.11, for each $\theta\in S^{1}$, the family $\{\widetilde{\Xi}\circ G_{1}(t,\theta)\}_{t\in[0,1]}$ is a pinch-off process.
To prove Theorem 3.16, we need more information about $\Xi(x)$ for $x\in C$ beyond the fact that $C$ is a genus 111 cap. More precisely, we need that the two regions, inside and outside, enclosed by each surface resemble solid tori.
Proposition 10.1.
In a Riemannian 333-sphere $(S^{3},g)$, let $\Sigma$ be a smooth, embedded torus. There exists a constant $\tau(\Sigma)>0$ with the following property.
Let $S\in\mathcal{S}(S^{3})$. By definition 2.1, let $\Omega,\Omega^{\prime}\subset S^{3}$ be the two open subsets of $S^{3}\setminus S$, whose reduced boundaries are both $S\setminus S_{\text{iso}}$. Assume that:
  1. (i)
    $H_{1}(\Omega;\mathbb{Z})$ and $H_{1}(\Omega^{\prime};\mathbb{Z})$ are both either 0 or $\mathbb{Z}$.
  2. (ii)
    $\mathbf{F}([S],[\Sigma])<\tau(\Sigma)$.
Then,
  1. (1)
    $H_{1}(\Omega;\mathbb{Z})\cong H_{1}(\Omega^{\prime};\mathbb{Z})\cong\mathbb{Z}.$
  2. (2)
    Fixing an inside direction for $\Sigma$ pointing towards $\mathrm{in}(\Sigma)$, there is a unique choice of an inside direction for $S$ such that
    $\operatorname{vol}(\mathrm{in}(S)\triangle\mathrm{in}(\Sigma)),\operatorname{vol}(\mathrm{out}(S)\triangle\mathrm{out}(\Sigma))<\tau(\Sigma)\,.$
    Here, $\{\mathrm{in}(S),\mathrm{out}(S)\}=\{\Omega,\Omega^{\prime}\}$ and $\{\mathrm{in}(\Sigma),\mathrm{out}(\Sigma)\}$ is the connected components of $S^{3}\setminus\Sigma$.
  3. (3)
    Fixing a generator $a_{0}$ for $H_{1}(\mathrm{in}(\Sigma);\mathbb{Z})$, there is a unique generator $a_{1}$ for $H_{1}(\mathrm{in}(S);\mathbb{Z})$ such that there exists a loop $\gamma_{\mathrm{in}}\subset\mathrm{in}(S)\cap\mathrm{in}(\Sigma)$ that induces both $a_{0}$ and $a_{1}$. And an analogous statement holds with $\mathrm{out}(\cdot)$ in place of $\mathrm{in}(\cdot)$.
Remark 10.2.
The condition “$H_{1}(\Omega;\mathbb{Z})$ and $H_{1}(\Omega^{\prime};\mathbb{Z})$ are both either 0 or $\mathbb{Z}$” is different from requiring “$\mathfrak{g}(S)=0$ or 1”. Indeed, consider this counterexample: Take three 2-spheres, identify all of their north poles as a point, and similarly, identify all of their three south poles as another point.
Proof.
First, choose a constant $d_{1}>0$ small enough such that $\Sigma$ has a tubular neighborhood of width $2d_{1}$. Fix an inside direction for $\Sigma$. For each $t\in(0,d_{1})$, the boundary of the $t$-neighborhood of $\Sigma$ consists of two smooth surfaces, $S_{t}$ and $S_{-t}$: We assume that $S_{t}$ lies outside $\Sigma$, while $S_{-t}$ inside. We denote $S_{0}\mathrel{\mathop{\ordinarycolon}}=\Sigma$. Now, for sufficiently small $\tau(\Sigma)$, from the fact that
$\mathcal{F}([S],[\Sigma])\leq\mathbf{F}([S],[\Sigma])<\tau(\Sigma)\,,$
item (2) follows.
Fix a small $\delta=\delta(\Sigma)>0$ to be determined by $\Sigma$ only. By $\mathbf{F}(|S|,|\Sigma|)<\tau(\Sigma)$ and by assuming $\tau(\Sigma)$ sufficiently small (depending on $\delta$ and $\Sigma$), we can assert that: By the coarea formula and Sard’s theorem, there exists a $t_{0}\in(-d_{1},-d_{1}/2)$ such that:
  • The intersection $S\cap S_{t_{0}}$ is transverse and is a finite union of smooth loops.
  • These loops are contained in finitely many balls $\{B_{r_{i}}(q_{i})\}_{i}$ with
    $\sum_{i}r_{i}<\delta\,.$
Since $\Sigma$ is a torus, we can fix generating elements
$a_{0}\in H_{1}(\mathrm{in}(\Sigma);\mathbb{Z})\cong\mathbb{Z}\,,\quad b_{0}\in H_{1}(\mathrm{out}(\Sigma);\mathbb{Z})\cong\mathbb{Z}\,,$
such that they have linking number 1 in $S^{3}$. Then, by the bullet points above and $\mathcal{F}([S],[\Sigma])<\tau(\Sigma)$, we can choose a sufficiently small $\delta=\delta(\Sigma)$, and thus $\tau(\Sigma)$, such that there exists a loop $\gamma_{\mathrm{in}}\subset S_{t_{0}}\cap\mathrm{in}(S)$ which avoids all loops in $S_{t_{0}}\cap S$ and satisfies $[\gamma_{\mathrm{in}}]=a_{0}\in H_{1}(\mathrm{in}(\Sigma);\mathbb{Z})$. Arguing similarly, we can choose a loop $\gamma_{\mathrm{out}}\subset\mathrm{out}(\Sigma)\cap\mathrm{out}(S)$ such that $[\gamma_{\mathrm{out}}]=b_{0}\in H_{1}(\mathrm{out}(\Sigma);\mathbb{Z})$. Noting the linking number $\operatorname{link}(\gamma_{\mathrm{in}},\gamma_{\mathrm{out}})$ is 111, we obtain item (1) of the proposition. Now, set
$a_{1}\mathrel{\mathop{\ordinarycolon}}=[\gamma_{\mathrm{in}}]\in H_{1}(\mathrm{in}(S);\mathbb{Z})\,,\quad b_{1}\mathrel{\mathop{\ordinarycolon}}=[\gamma_{\mathrm{out}}]\in H_{1}(\mathrm{out}(S);\mathbb{Z})\,.$
From $\operatorname{link}(\gamma_{\mathrm{in}},\gamma_{\mathrm{out}})=1$, item (3) also follows easily. ∎
Recall that by assumption, $(S^{3},g^{\prime})$ has only finitely many embedded minimal tori. For each embedded minimal torus $\Sigma$, the above proposition gives a constant $\tau(\Sigma)>0$.
Without loss of generality, we now assume that the constant $d_{0}>0$, chosen in the proof of Theorem A (in the paragraph right before §3.2.1) is smaller than $\tau(\Sigma)$ for every embedded minimal torus $\Sigma$. Then for each $x\in c\;(\subset\mathrm{dmn}(\Xi))$, we have the following:
  • $\mathbf{F}([\Xi(x)],[T])<\tau(T)$. In particular, $\Xi(x)\in\mathcal{S}(M)$.
  • Since the family $\widetilde{\Xi}\circ\widetilde{F}(\cdot,x)$ is a pinch-off process (by Proposition 3.11), by considering the shape of the members in original family $\Psi$ (see §5.3.2), we know by Proposition 9.1, that $\Xi(x)$ satisfies the assumptions on $S$ in the first bullet point of Proposition 10.1.
As a result, we can apply Proposition 10.1 to $\widetilde{\Xi}\circ G_{1}(1,\theta)$ for each $\theta\in S^{1}$ (as $G_{1}(1,\cdot)$ is a parametrization for the loop $c$). Thus, fixing an inside direction for the torus $T$, we have:
  1. (1)
    For each $\theta\in S^{1}$,
    (10.3) $H_{1}(\mathrm{in}(\widetilde{\Xi}\circ G_{1}(1,\theta));\mathbb{Z})\cong H_{1}(\mathrm{out}(\widetilde{\Xi}\circ G_{1}(1,\theta));\mathbb{Z})\cong\mathbb{Z}\,.$
  2. (2)
    For each $\theta$, there is a unique way to choose an inside direction for $\widetilde{\Xi}\circ G_{1}(1,\theta)$ such that
    $\operatorname{vol}(\mathrm{in}(T)\triangle\mathrm{in}(\widetilde{\Xi}\circ G_{1}(1,\theta))),\operatorname{vol}(\mathrm{out}(T)\triangle\mathrm{out}(\widetilde{\Xi}\circ G_{1}(1,\theta)))<\tau(T)\,,$
    and this choice is continuous in $\theta$.
  3. (3)
    Let us fix a generator $a_{0}$ for $H_{1}(\mathrm{in}(T);\mathbb{Z})$. For each $\theta$, there is a unique generator $a(\theta)$ of $H_{1}(\mathrm{in}(\widetilde{\Xi}\circ G_{1}(1,\theta));\mathbb{Z})$ such that there exists a loop $\gamma_{a}(\theta)\subset\mathrm{in}(\widetilde{\Xi}\circ G_{1}(1,\theta))\cap\mathrm{in}(T)$ that induces both $a_{0}$ and $a(\theta)$.
  4. (4)
    Let us fix a generator $b_{0}$ for $H_{1}(\mathrm{out}(T);\mathbb{Z})$. For each $\theta$, there is a unique generator $b(\theta)$ of $H_{1}(\mathrm{out}(\widetilde{\Xi}\circ G_{1}(1,\theta));\mathbb{Z})$ such that there exists a loop $\gamma_{b}(\theta)\subset\mathrm{out}(\widetilde{\Xi}\circ G_{1}(1,\theta))\cap\mathrm{out}(T)$ that induces both $b_{0}$ and $b(\theta)$.
10.2. Relating $\Xi|_{c}$ to $\Psi$
Then, using the item (2) above above, we can continuously and uniquely choose an inside direction for $\widetilde{\Xi}\circ G_{1}(t,\theta)$ for each $(t,\theta)\in[0,1]\times S^{1}$. Moreover, from the fact that $\widetilde{\Xi}\circ G_{1}(\cdot,\theta)$ is a pinch-off process for each $\theta$, we know by Proposition 9.1 that
$\operatorname{rank}(H_{1}(\mathrm{in}(\widetilde{\Xi}\circ G_{1}(t,\theta));\mathbb{Z}))\,,\quad\operatorname{rank}(H_{1}(\mathrm{out}(\widetilde{\Xi}\circ G_{1}(t,\theta));\mathbb{Z}))\,,$
are both non-increasing in $t$. Thus, together with (10.3) and the description of the family $\Psi$ given in §5.3.2, we obtain:
  • For each $(t,\theta)\in[0,1]\times S^{1}$,
    $H_{1}(\mathrm{in}(\widetilde{\Xi}\circ G_{1}(t,\theta));\mathbb{Z})\cong H_{1}(\mathrm{out}(\widetilde{\Xi}\circ G_{1}(t,\theta));\mathbb{Z})\cong\mathbb{Z}\,.$
    (Recall that the first homology, in $\mathbb{Z}$-coefficients, of an open subset of $S^{3}$ has no torsion.)
  • The family $\widetilde{\Xi}\circ G_{1}(0,\cdot)$, which can be viewed as the same as $\Psi|_{\widetilde{F}(0,c)}$, consists entirely of smooth tori, where $\widetilde{F}(0,c)$ is a loop in $Y$.
Now, we prove a lemma stating that all smooth tori in $\Psi$ can be deformation retracted to Clifford tori.
Lemma 10.3.
If we view $\Psi$ as a Simon-Smith family in the unit 333-sphere $\mathbb{S}^{3}$, and let $Z_{0},Z_{1}\subset Y$ denote the sets of parameters corresponding to Clifford tori and smooth tori, respectively, then $Z_{1}$ can be deformation retracted onto $Z_{0}$.
Proof.
Proposition 5.2 tells us that for each $v\in{\mathbb{B}^{4}}$ and $z\in\widetilde{\mathcal{C}}$, we have an interval $I_{v,z}\subset(-\pi,\pi)$ such that for each $t\in I_{v,z}$,
$\Sigma(v,t,z)\mathrel{\mathop{\ordinarycolon}}=\partial\{x\in S^{3}\mathrel{\mathop{\ordinarycolon}}d(v,z)(x)<t\}$
[…]
$Z_{2}\subset\{[(v,t,z)]\in Y\mathrel{\mathop{\ordinarycolon}}v\in{\mathbb{B}^{4}},t=0\}.$
Thus, $Z_{2}$ can be deformation retracted onto the set
$Z_{0}=\{[(v,t,z)]\in Y\mathrel{\mathop{\ordinarycolon}}v=0,t=0\}\,,$
which parametrizes the set of Clifford tori under $\Psi$ in $\mathbb{S}^{3}$. ∎
Thus, by the above lemma, there exists a homotopy
$G_{2}\mathrel{\mathop{\ordinarycolon}}[0,1]\times S^{1}\to Y\subset\widetilde{W}$
such that $G_{2}(0,\cdot)$ lies in $Z_{0}$ and $G_{2}(1,\cdot)=G_{1}(0,\cdot)$. Let $c_{0}$ denote the loop in $Z_{0}$ parametrized by $G_{2}(0,\cdot)$. By concatenating the two homotopies $G_{1}$ and $G_{2}$, we obtain a homotopy
$G_{3}\mathrel{\mathop{\ordinarycolon}}[0,1]\times S^{1}\to\widetilde{W}$
such that:
  • $G_{3}(0,\cdot)$ parametrizes $c_{0}\subset Z_{0}\subset\widetilde{W}$.
  • $G_{3}(1,\cdot)$ parametrizes $c\subset\mathrm{dmn}(\Xi)\subset\widetilde{W}$.
  • For each $\theta\in S^{1}$, the family $\widetilde{\Xi}\circ G_{3}(\cdot,\theta)$ is a pinch-off process.
  • For each $(t,\theta)\in[0,1]\times S^{1}$,
    $H_{1}(\mathrm{in}(\widetilde{\Xi}\circ G_{3}(t,\theta));\mathbb{Z})\cong H_{1}(\mathrm{out}(\widetilde{\Xi}\circ G_{3}(t,\theta));\mathbb{Z})\cong\mathbb{Z}\,.$
[…]
Recall that $Z_{0}\cong{\mathbb{RP}}^{2}\times{\mathbb{RP}}^{2}$, so $H_{1}(Z_{0};\mathbb{Z}_{2})=\mathbb{Z}^{2}\times\mathbb{Z}^{2}$ has 3 non-trivial elements: (1,0),(0,1),(1,1)(1,0),(0,1),(1,1)( 1 , 0 ) , ( 0 , 1 ) , ( 1 , 1 ). Note, on one hand, (1,0)(1,0)( 1 , 0 ) and (0,1)(0,1)( 0 , 1 ) (but not (1,1)(1,1)( 1 , 1 )) both correspond to 1-sweepouts under $\Psi$ as we discussed in §5.4. On the other hand, $\Xi|_{c}$ is not a 1-sweepout as it is close to a single embedded minimal torus in the flat topology by (10.2), and this implies $\Psi|_{c_{0}}$, which is homotopic to $\Xi|_{c}$ in the flat topology via $\widetilde{\Xi}\circ G_{3}$, is not a 1-sweepout either. Thus, $[c_{0}]$ cannot be (1,0)(1,0)( 1 , 0 ) or (0,1)(0,1)( 0 , 1 ).
It suffices rule out the possibility that $[c_{0}]=(1,1)$. Let us extend the inside directions we chose for the family $\widetilde{\Xi}\circ G_{1}$ to the family $\widetilde{\Xi}\circ G_{3}$. Note that this extension is continuous and unique. As a result, we can “naturally identify” the groups
$H_{1}(\mathrm{in}(\widetilde{\Xi}\circ G_{3}(t,\theta));\mathbb{Z})\,,\quad(t,\theta)\in[0,1]\times S^{1}\,,$
in the following sense. Let $k\in\mathbb{N}$. For each $i,j=1,\cdots,k$, define the closed rectangle
$R_{i,j}=\left[\frac{i-1}{k},\frac{i}{k}\right]\times\left[2\pi\frac{j-1}{k},2\pi\frac{j}{k}\right]\subset[0,1]\times S^{1}\,.$
Proposition 10.4.
[…]
  1. (1)
    $\gamma^{\mathrm{in}}_{i,j}$ lies in $\mathrm{in}(\widetilde{\Xi}\circ G_{3}(p))$ and generates $H_{1}(\mathrm{in}(\widetilde{\Xi}\circ G_{3}(p));\mathbb{Z})$.
  2. (2)
    $\gamma^{\mathrm{out}}_{i,j}$ lies in $\mathrm{out}(\widetilde{\Xi}\circ G_{3}(p))$ and generates $H_{1}(\mathrm{out}(\widetilde{\Xi}\circ G_{3}(p));\mathbb{Z})$.
  3. (3)
    $\gamma^{\mathrm{in}}_{i,j}$ and $\gamma^{\mathrm{in}}_{m,n}$ are homologous in $\mathrm{in}(\widetilde{\Xi}\circ G_{3}(p))$.
  4. (4)
    $\gamma^{\mathrm{out}}_{i,j}$ and $\gamma^{\mathrm{out}}_{m,n}$ are homologous in $\mathrm{out}(\widetilde{\Xi}\circ G_{3}(p))$.
Proof.
For each $(t,\theta)\in[0,1]\times S^{1}$, since
$H_{1}(\mathrm{in}(\widetilde{\Xi}\circ G_{3}(t,\theta));\mathbb{Z})\cong H_{1}(\mathrm{out}(\widetilde{\Xi}\circ G_{3}(t,\theta));\mathbb{Z})\cong\mathbb{Z}\,,$
there exist $\gamma^{\mathrm{in}}_{(t,\theta)}\subset\mathrm{in}(\widetilde{\Xi}\circ G_{3}(t,\theta))$ and $\gamma^{\mathrm{out}}_{(t,\theta)}\subset\mathrm{out}(\widetilde{\Xi}\circ G_{3}(t,\theta))$ generating these two homology groups, respectively. Moreover, from the fact that $\widetilde{\Xi}\circ G_{3}$ is a Simon-Smith family (Definition 2.4 (2)), there exists a neighborhood $O_{(t,\theta)}\subset[0,1]\times S^{1}$ of $(t,\theta)$, such that $\gamma^{\mathrm{in}}_{(t,\theta)}\subset\mathrm{in}(\widetilde{\Xi}\circ G_{3}(p))$ and generates
$H_{1}(\mathrm{in}(\widetilde{\Xi}\circ G_{3}(p));\mathbb{Z})\,,$
and $\gamma^{\mathrm{out}}_{(t,\theta)}\subset\mathrm{out}(\widetilde{\Xi}\circ G_{3}(p))$ and generates
$H_{1}(\mathrm{out}(\widetilde{\Xi}\circ G_{3}(p));\mathbb{Z})\,.$
Clearly, $\{O_{(t,\theta)}\}_{(t,\theta)\in[0,1]\times S^{1}}$ is an open covering of $[0,1]\times S^{1}$. Since $[0,1]\times S^{1}$ is compact, we can find a finite covering, denoted by
$\{O_{l}\}^{t}_{l=1}\,,$
together with the corresponding loops $\{\gamma^{\mathrm{in}}_{l}\}$ and $\{\gamma^{\mathrm{out}}_{l}\}$. Moreover, we can find a sufficiently large $k$ such that for each $i,j=1,\cdots,k$, there exists some $l$, such that
$N(R_{i,j})\subset O_{l}$
where $N(R_{i,j})$ is the union of $R_{i,j}$ and all the rectangles adjacent to $R_{i,j}$.
Let us first work on the rectangles $R_{k,1},\cdots,R_{k,k}$.
For each $R_{k,j}$, it contains some point $(1,\theta)$. Using the property (3) and (4) for the family $\widetilde{\Xi}\circ G_{3}(1,\cdot)$, there exists a loop $\gamma_{a}(\theta)$ that lies in
$\mathrm{in}(\widetilde{\Xi}\circ G_{3}(1,\theta))\cap\mathrm{in}(T)\,.$
induces both $a_{0}$ and $a(\theta)$. Suppose that $R_{k,j}\subset O_{l}$, and then we can choose $\gamma^{\mathrm{in}}_{k,j}$ to be either $\gamma^{\mathrm{in}}_{l}$ or $-\gamma^{\mathrm{in}}_{l}$ such that
$[\gamma^{\mathrm{in}}_{k,j}]=a_{0}\in H_{1}(\mathrm{in}(T);\mathbb{Z})\,,$
and thus,
$[\gamma^{\mathrm{in}}_{k,j}]=a(\theta)\in H_{1}(\mathrm{in}(\widetilde{\Xi}\circ G_{3}(1,\theta));\mathbb{Z})\,.$
[…]
$[\gamma^{\mathrm{in}}_{i-1,j}]=[\gamma^{\mathrm{in}}_{i,j}]\in H_{1}(\mathrm{in}(\widetilde{\Xi}\circ G_{3}(i/k,2\pi j/k));\mathbb{Z})\,.$
[…]
$\operatorname{link}(\gamma^{\mathrm{in}}_{m,n},\gamma^{\mathrm{out}}_{k,j})=1,\quad\operatorname{link}(\gamma^{\mathrm{out}}_{m,n},\gamma^{\mathrm{in}}_{k,j})=1\,.$
This confirms statements (3) and (4) of the proposition in this case.
In conclusion, one can find such a large $k$ and the corresponding loops. ∎
10.3. The family $\Phi|_{c_{0}}$
Focusing on the part in $\{0\}\times S^{1}\subset[0,1]\times S^{1}$ in Proposition 10.4, one obtains a map $L_{0}\mathrel{\mathop{\ordinarycolon}}S^{1}\times S^{1}\to S^{3}$ such that for each $\theta\in S^{1}$, the loop $L_{0}(\theta,\cdot)=\gamma^{\mathrm{in}}_{0,j}$ for some $j$, which lies in $\mathrm{in}(\widetilde{\Xi}\circ G_{3}(0,\theta))$ and generates
$H_{1}(\mathrm{in}(\widetilde{\Xi}\circ G_{3}(0,\theta));\mathbb{Z})\cong\mathbb{Z}\,,$
satisfying conclusions (3) and (4) of the previous proposition.
However, by the topological lemma below, this is impossible if $[c_{0}]=(1,1)$.
Lemma 10.5.
Let $d\mathrel{\mathop{\ordinarycolon}}[0,2\pi]\to Z_{0}$ with $d(0)=d(2\pi)$ parametrize a loop representing $(1,1)\in H_{1}(Z_{0};\mathbb{Z}_{2})$. We continuously choose an inward direction for $\Psi\circ d(\theta)$, for each $\theta\in[0,2\pi]$. Let
$L\mathrel{\mathop{\ordinarycolon}}[0,2\pi]\times S^{1}\to S^{3}$
be a continuous map such that each loop $L(\theta,\cdot)$ lies in $\mathrm{in}(\Psi\circ d(\theta))$ and generates
$H_{1}(\mathrm{in}(\Psi\circ d(\theta));\mathbb{Z})\cong\mathbb{Z}\,.$
Then $[L(0,\cdot)]$ and $[L(2\pi,\cdot)]$, which both are elements in $H_{1}(\mathrm{in}(\Psi\circ d(0));\mathbb{Z})$, differ exactly by a sign.
Proof.
[…]
Consider the set
[…]
$a_{0}\textrm{ is one of the two generators of }H_{1}(\Omega;\mathbb{Z})\cong\mathbb{Z}\}.$
Evidently, there is a natural topology that can equipped on this set, such that it is a double cover of the space
[…]
Note that $L$ induces a map $L^{\prime}$ from $[0,2\pi]$ into the space (10.4). To prove the lemma, it suffices to show that, while the solid tori corresponding to $L^{\prime}(0)$ and $L^{\prime}(2\pi)$ are the same, the first homology classes of the solid torus given by $L^{\prime}(0)$ and $L^{\prime}(2\pi)$ differ by a sign.
Following the discussion in §5.1, we know that the space (10.4) is homeomorphic to $G_{2}^{+}(\mathbb{R}^{4})$, the space of oriented 2-planes in $\mathbb{R}^{4}$. But $G_{2}^{+}(\mathbb{R}^{4})\cong S^{2}\times S^{2}$. Thus, to prove the lemma, it suffices to show that the two oriented 2-planes corresponding to $(x,y),(-x,-y)\in S^{2}\times S^{2}$ are the same except differing by an orientation. This was already proved in [Nur16, §3.4.3]. ∎
In conclusion, $[c_{0}]\neq(1,1)$ in $H_{1}(Y;\mathbb{Z}_{2})$, so $[c_{0}]=(0,0)$, as desired. This finishes the proof of Theorem 3.16.